### v2

# Identification of Dynamic Nonlinear Panel Models

Abstract
This paper studies identification for a wide range of nonlinear panel data models, including binary choice, ordered repsonse, and other types of limited dependent variable models. Our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of (potentially contemporary) endogeneity. Our identification strategy relies on a partial stationarity condition, which not only allows for an unknown distribution of errors but also for temporal dependencies in errors. We derive partial identification results under flexible model specifications and provide additional support conditions for point identification. We demonstrate the robust finite-sample performance of our approach using Monte Carlo simulations, and apply the approach to analyze the empirical application of income categories using various ordered choice models.
Keywords: Panel Discrete Choice Models; Stationarity; Dynamic Models; Partial Identification; Endogeneity
1 Introduction
This paper studies semiparametric partial identification of a wide range of panel data models with limited dependent variables, including various discrete choice models, censored dependent variable models, and nonseparable models. In particular, our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of endogenous covariates.
To fix ideas, we start with the following dynamic binary choice model, which is on its own of considerable theoretical and applied interest. Sections 3-6 generalize the approach to other limited dependent variable models. Specifically, consider
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (1)
where $Y_{it}\in\{0,1\}$ denotes a binary outcome variable for individual $i=1,2,...$ and time $t=1,...,T$, while $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$. The objective is to identify the parameter $\theta_{0}:=(\beta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime}$ using a panel of observed variables $(Z_{it},X_{it},Y_{it})_{it}$. We focus on short panels, where the number of time periods $T\geq 2$ is fixed and finite.
The identification of model (1) has been explored in the literature under various assumptions. For example, Chamberlain (1980) examines identification under the logistic distribution of $\epsilon_{it}$ and the independence of $\epsilon_{it}$ with respect to $(\alpha_{i},\{(Z_{it},X_{it})\}_{t=1}^{T})$. Subsequently, Manski (1987) relaxes the distributional assumption and employs the following conditional stationarity of $\epsilon_{it}$ to achieve identification:
$\epsilon_{is}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\ \stackrel{{\scriptstyle d}}{{\sim}}\ \epsilon_{it}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (2)
This condition is also referred to as “group stationarity” or “group homogeneity” and has also been exploited in studies such as Chernozhukov et al. (2013), Shi, Shum, and Song (2018) and Pakes and Porter (2022). Condition (2) does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows dependence between the fixed effect $\alpha_{i}$ and the covariates $(Z_{it},X_{it})$. However, condition (2) does impose substantial restriction on the dependence between $(Z_{it},X_{it})$ and the time-varying error term $\epsilon_{it}$: it effectively requires that all covariates in $(Z_{it},X_{it})$ are exogeneous with respect to the time varying error $\epsilon_{it}$.222 For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=X_{it}^{\prime}\eta$, then the conditional distributions of $\epsilon_{it}$ and $\epsilon_{is}$ cannot be the same as long as $X_{it}^{\prime}\eta\neq X_{is}^{\prime}\eta$, so condition (2) fails in general.
In many economic applications, certain components of the observable covariates, namely $X_{it}$, may exhibit endogeneity. For example, in a dynamic setting where $X_{it}$ includes the lagged outcome variable $Y_{i,t-1}$, then endogeneity of $Y_{i,t-1}$ with respect to $\epsilon_{i,t-1}$ arises immediately. For another example, if $X_{it}$ includes “price” or other variables that may be endogenously chosen by economic agents, then the exogeneity restriction imposed by condition (2) will again fail to hold.
We propose in this paper a general identification approach for various nonlinear panel models in the presence of endogenous regressors, using the following “partial stationarity” condition, which can be viewed as a weaker verision of the stationarity condition (2) above:
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\ \stackrel{{\scriptstyle d}}{{\sim}}\ \epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (3)
Our partial stationary condition (3), as its name suggests, only requires that the errors are stationary across time periods conditional on the realizations of a subvector of the covariates (i.e., the exogenous covariates denoted by $Z$) while allowing the remaining covariates (denoted by $X$) to be endogenous in arbitrary manners. In short, condition (3) imposes exogeneity conditions only on exogenous covariates. Alternatively, we can interpret condition (3) as an assumption of the existence of some covariates being exogenous. [^3] Condition (3) also accommodates the standard stationarity assumption conditional on all covariates.
We describe how to exploit the partial stationarity condition (3) to derive the identified set on the model parameters $\theta_{0}$ through conditional moment inequalities, which take the form of lower and upper bounds for the conditional distribution $\epsilon_{it}+\alpha_{i}\mid Z_{is},Z_{it}$, solely as functions of observed variables and the model parameters $\theta_{0}$. We show that these bounds must have nonzero intersections over time under the partial stationarity assumption, thereby forming identifying restrictions for the parameter $\theta_{0}$.
While extensive work exists on nonlinear panel models under stationarity, previous studies typically focus on individual models. The identification strategies are often context specific and the identification results may have various complicated representations, as seen in studies such as Khan, Ponomareva, and Tamer (2023) and Pakes and Porter (2022). Our approach offers the advantage of providing a simple and unified characterization of the identified set for a broad range of static and dynamic panel models, irrespective of the specific types of variables (discrete/continuous outcome and covariates) and forms of endogeneity (lagged/contemporary endogenous regressors). Furthermore, our strategy does not rely on scalar-additive structures that are often imposed in various parametric and semiparametric models, and can be further extended to accommodate nonseparable panel data models.
We demonstrate the sharpness of the identified set for binary choice models when the support of $X_{it}$ is finite. More precisely, we show that, for any $\theta$ that satisfies all the conditional moment inequalities we derived, we can construct an observationally equivalent joint distribution of the observed and unobserved variables in our model. While our main result is about set identification, we also provide sufficient conditions for the point identification of the coefficients $\beta_{0}$ on exogenous covariates (under scale normalization) as well as the signs of the coefficients on endogenous covariates $\gamma_{0}$.
Our identification strategy based on partial stationarity can be applied more broadly beyond the context of dynamic binary choice models. We demonstrate its applicability to other limited dependent variable models, such as ordered response models, multinomial choice models, and censored outcome models. The results of our approach accommodates both static and dynamic settings across all these models. Furthermore, we illustrate the adaptability of our key strategy to nonseparable semiparametric models with endogeneity.
We characterize the identified set using a collection of conditional moment inequalities, based on which estimation and inference can be conducted using established econometric methods in the literature, such as Chernozhukov, Hong, and Tamer (2007), Andrews and Shi (2013), and Chernozhukov, Lee, and Rosen (2013). Through Monte Carlo simulations, we demonstrate that our identification method yields informative and robust finite-sample confidence intervals for coefficients in both static and dynamic models.
Literature Review
Our paper contributes directly to the line of econometric literature on semiparametric panel discrete choice models. Dating back to Manski (1987), a series of work exploits “full” stationarity conditions for identification, such as Chernozhukov et al. (2013), Khan, Ponomareva, and Tamer (2016), Shi, Shum, and Song (2018), Pakes and Porter (2022), Khan, Ouyang, and Tamer (2021), Khan, Ponomareva, and Tamer (2023), Gao and Li (2020), and Wang (2022). As discussed above, full stationarity conditions given all observable covariates effectively require that all covariates are exogenous with no dynamic effects (i.e., lagged dependent variables). In contrast, we exploit the “partial” stationarity condition, thereby allowing for lagged dependent variables as well as other endogenous covariates.
In the literature on dynamic discrete choice models, our paper is most closely related to Khan, Ponomareva, and Tamer (2023, KPT thereafter), who studies the following dynamic panel binary choice model
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (4)
where the one-period lagged dependent variable $Y_{i,t-1}\in\{0,1\}$ serves as the endogenous covariate, and $Z_{it}$ are exogenous covariates. KPT exploits the following assumption:
$\epsilon_{is}\mid Z_{it},Z_{is},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{it},Z_{is},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
which is exactly a “partial stationarity” condition in the specific context of (4), and derives the sharp identified set for $\theta_{0}$ by explicitly enumerating the realizations of the one-period lagged outcome variable $Y_{i,t-1}$ (across two periods $t,s$). In contrast, our model (1), along with the “partial stationarity” condition, is stated with more general specifications of the endogenous covariates $X_{it}$. The covariates $X_{it}$ can include more than one lagged dependent variables (e.g. $Y_{i,t-1},Y_{i,t-2},$) and other endogenous variables (such as “price” if $Y_{it}$ represents the purchase of a particular product), which may be continuously valued. Consequently, our identification strategy is substantially different from that of KPT, and can be applied more broadly to various other dynamic nonlinear panel models. In the specific model (4), we show that the identifying restrictions we derived are equivalent to those derived in KPT and thus both approaches lead to sharp identification. Relatedly, Mbakop (2023) proposes a computational algorithm to derive conditional moment inequalities in a general class of dynamic discrete choice models (potentially with multiple lags). The focus of Mbakop (2023) is on scenarios where the lagged discrete outcome variables are the only endogenous covariates in the model, and the proposed algorithm relies on the discreteness of these variables. Relative to these works, our paper introduces an analytic approach that directly applies to a more general class of dynamic binary choice models, as well as other types of models with continuous limited dependent variables and any number of endogenous covariates, regardless of whether they are discrete or continuous and whether they take the form of lagged outcome variables or not.
Our identification strategy relies on a type of stationarity condition, while alternative approaches utilize other notions of exogeneity. For example, Aristodemou (2021) exploits a type of “full independence” assumption to identify dynamic binary choice models. The “full independence” assumption essentially requires that the time-varying errors from all time periods and the exogenous variables from all time periods are independent (conditional on initial conditions), but does not make intertemporal restrictions on the errors (such as stationarity”. Hence, such ‘full independence” assumption and the partial stationarity assumption in our paper do not nest each other as special cases. Chesher, Rosen, and Zhang (2023) applies the framework of generalized instrumental variables (Chesher and Rosen, 2017) to the context of various dynamic discrete choice models with fixed effects, and utilizes a similar “full independence” assumption (Aristodemou, 2021) for identification. [^4] Our identification strategy shares some conceptual similarity with the idea of generalized instrumental variable (GIV) in Chesher and Rosen (2017), who proposes a general approach for representing the identified set of structural models with endogeneity. Chesher and Rosen (2017), Chesher and Rosen (2020), and Chesher, Rosen, and Zhang (2023) demonstrate how the GIV framework can be applied to various settings, but focus mostly on the use of exclusion restrictions and/or full independence assumptions. In this paper, we neither impose exclusion restrictions nor independence assumptions but instead explore identification under a partial stationarity condition. More differently, some other papers work with sequential exogeneity in various dynamic panel models and provide (non-)identification results under different model restrictions. For example, Shiu and Hu (2013) imposes a high-level invertibility condition along with a restriction that rules out the dependence of covariates on past dependent variables. More recently, Bonhomme, Dano, and Graham (2023) investigates panel binary choice models with a single binary predetermined covariate under sequential exogeneity, whose evolution may depend on the past history of outcome and covariate values. The sequential exogeneity condition considered in these papers and the partial stationarity condition in ours again do not nest each other as special cases: in particular, our current paper accommodates contemporarily endogenous covariates that violate sequential exogeneity. In summary, the key assumptions, identification strategy, and identification results of these studies are substantially different from and thus not directly comparable to those in our current paper.
Our paper is also complementary to the line of literature that studies dynamic logit models with fixed effects for binary, ordered responses, or multinomial choice models. This literature typically assumes that time-varying errors are conditionally independent across time, independent from all other variables, and follow the logistic distribution. The logit assumption in panel data models has long been studied, such as in Chamberlain (1984) and Chamberlain (2010). In the context of dynamic discrete choice models, Honoré and Kyriazidou (2000) first shows how to conduct differencing of fixed effects under the logit assumption, while recent papers by Honoré and Weidner (2020) and Dano (2023) illustrate how to systematically obtain moment conditions free of fixed effects and time-varing errors. Honoré, Muris, and Weidner (2021) extends the approach in Honoré and Weidner (2020) to dynamic ordered logit models. Meanwhile, Dobronyi, Gu, and Kim (2021) derives sharp identification for dynamic logit models using a different approach based on truncated moments. Alternatively, Honoré and Tamer (2006) proposes a linear programming method to obtain bounds on model parameters and average marginal effects under logit and other parametric error distributions. In addition, Davezies, D’Haultfoeuille, and Laage (2021) provides analytic bounds on average marginal effects in static logit models. Relative to this line of literature, our paper does not require parametric (logistic) or conditional independence assumptions, and provides general semiparametric identification results for various dynamic panel models.
The rest of the paper is organized as follows. Section 2 studies the sharp identification of binary choice models with endogenous covariates. We provide sufficient conditions to achieve point identification and explore two classic models: static and dynamic models. Sections 3-5 apply the identification approach to ordered response models, multinomial choice models, and models with censored dependent variables. Section 6 broadens the scope to general nonseparable models. Section 7 presents simulation results about the finite-sample performances of our approach and Section 8 explores the empirical application of income categories using various ordered response models. We conclude with Section 9.
2 Binary Choice Model
2.1 Model
In this section, we focus on the identification of dynamic binary choice models with endogenous covariates. Specifically, consider the following binary choice model:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (5)
where $Y_{it}\in\{0,1\}$ denotes the binary dependent variable for individual $i\in\{1,2,...\}$ and time $t\leq T$, $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$. The objective is to identify the parameter $\theta_{0}:=(\beta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime}$ using a short panel of observed variables $(Z_{it},X_{it},Y_{it})_{it}$, where the number of time periods $T\geq 2$ is fixed and finite.
The identification of model (5) has been explored in the literature under various assumptions. [^5] See Section 1 for more detailed discussions on related literature. Specifically, Manski (1987) exploits the following conditional stationarity (also referred to as “group homogeneity”) of $\epsilon_{it}$ to achieve identification:
$\epsilon_{is}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (6)
The above condition does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows for dependence of $\epsilon_{it}$ over time. However, condition (2) does impose substantial restriction on the dependence between $(Z_{it},X_{it})$ and the time-varying error $\epsilon_{it}$. For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=Z_{it}^{\prime}\eta_{1}+X_{it}^{\prime}\eta_{2}$, then condition (2) fails in general. Hence, condition (2) can also be interpreted as a form of exogeneity condition on $\epsilon_{it}$ and all covariates $(Z_{it},X_{it})$. Importantly, in a dynamic setting where $X_{it}$ includes the lagged dependent variable $Y_{i,t-1}$, then condition (2) is no longer a justifiable assumption to work with.
Our paper aims to study identification of $\theta_{0}$ robust to endogeneity of $X_{it}$, where $X_{it}$ can be arbitrarily dependent with $(\alpha_{i},\epsilon_{it})$. Before introducing our assumptions, we first provide several applications to illustrate this endogeneity.
Example 1 (Dynamic Models).
Khan, Ponomareva, and Tamer (2023) studies the following dynamic model with one lagged dependent variable:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+Y_{it-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
under the stationarity assumption only given exogenous covariates $Z_{it}$: $\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}$. In this model, the full stationarity condition may naturally fail due to its dynamic nature. For example, suppose $\epsilon_{it}$ is i.i.d. across time and independent of ($(\{Z_{it}\}_{t=1}^{T},\alpha_{i})$. Given $Z_{is}=Z_{it}=z,\alpha_{i}=a$, it is clear that the conditional distribution of $\epsilon_{it}\mid Y_{it-1},Y_{it},z,a$ will be different from $\epsilon_{it+1}\mid Y_{it-1},Y_{it},z,a\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it+1}$, since $Y_{it}$ is informative for $\epsilon_{it}$.
Example 2 (Omitted Variable).
Consider the true model is given as $Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+S_{it}\eta_{0}+\alpha_{i}+v_{it}\geq 0\}$, where $v_{it}$ is independent of $(\alpha_{i},\{Z_{it},X_{it},S_{it}\}_{t=1}^{T})$. However, the variable $S_{it}$ is omitted and we instead estimate the following model:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
where $\epsilon_{it}=v_{it}+S_{it}\eta_{0}$. The covariate $X_{it}$ can be correlated with $\epsilon_{it}$ through the omitted variable $S_{it}$, provided that the omitted variable $S_{it}$ is correlated with $X_{it}$. Then the conditional distribution of $\epsilon_{it}\mid(Z_{is},Z_{it},X_{is},X_{it},\alpha_{i})$ may differ from $\epsilon_{is}\mid(Z_{is},Z_{it},X_{is},X_{it},\alpha_{i})$ since the distribution of $S_{it}\mid X_{it}$ could be different from the distribution $S_{is}\mid X_{is}$ when $X_{it}\neq X_{is}$.
Example 3 (Measurement Error).
Consider the true model is given as $Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{*^{\prime}}\gamma_{0}+\alpha_{i}+v_{it}\geq 0\}$, where $v_{it}$ is independent of $\left(\alpha_{i},\{Z_{it},X_{it}^{*}\}_{t=1}^{T}\right)$. However, the covariate $X_{it}^{*}$ may not be observed and we only observe its measurement $X_{it}$ which may be subject to a measurement error $X_{it}=X_{it}^{*}+e_{it}$. Suppose we estimate the following model using the measurement $X_{it}$ as a regressor:
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
where $\epsilon_{it}=v_{it}-e_{it}^{\prime}\gamma_{0}$. The error term $\epsilon_{it}$ would be correlated with the measurement $X_{it}$ through the measurement error $e_{it}$. Similarly, the conditional stationarity in (2) could fail since $X_{it}$ is informative for $e_{it}$ such that the conditional distribution of $\epsilon_{it}\mid z,X_{is},X_{it},\alpha_{i}$ would vary over time given different covariates $X_{is}\neq X_{it}$ while holding $Z_{is}=Z_{it}=z$.
Motivated by the above examples, this paper aims to establish identification of $\theta_{0}$, while being robust to endogeneity of $X_{it}$ with respect to unobserved factors $(\alpha_{i},\epsilon_{it})$. Below we propose an approach that accommodates various types of endogeneity, regardless of whether the endogenous covariates $X_{it}$ are discrete lagged outcome variables (as in Example 1) or continuously-valued covariates with contemporary endogeneity as in other examples. To achieve this, we introduce the following partial conditional stationarity assumption:
Assumption 1 (Partial Stationarity).
The conditional distribution of $\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}$ is stationary over time:
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
Assumption 1 only assumes the stationarity given the covariate $(Z_{is},Z_{it})$, but does not impose any restrictions on the endogenous covariate $X_{it}$. The standard full stationarity condition (2) is nested in Assumption 1, and it is in general stronger than Assumption 1. The main distinction is that Assumption 1 allows $X_{it}$ to be arbitrarily dependent with both the fixed effect $\alpha_{i}$ and the time-changing error $\epsilon_{it}$. In addition, their correlation can be different across time. Assumption 1 also shares similar features with condition (2), in that it does not impose any distributional assumption on $\epsilon_{it}$ and allows $\epsilon_{it}$ to be correlated across time. The standard i.i.d. assumption of $\epsilon_{it}$ across time or independence of $\epsilon_{it}$ with $(\alpha_{i},Z_{it},X_{it})$ are also nested in Assumption 1 as special cases.
2.2 Identification Strategy
We now explain our key identification strategy based on partial stationarity. While it applies more generally, we first illustrate the key idea in the specific context of the dynamic binary choice model. Sections 3-6 provide generalizations of the idea to various models.
Let $v_{it}:=-(\epsilon_{it}+\alpha_{i})$, $W_{it}:=(Z_{it},X_{it})$, and $W_{ist}:=(W_{is},W_{it})$. The conditional choice probability is given as follows:
$\Pr\left(Y_{it}=1\mid W_{ist}=w_{st}\right)=\Pr\left(v_{it}\leq w_{t}^{\prime}\theta_{0}\mid W_{ist}=w_{st}\right).$
When assuming all covariates are exogenous under condition (2), variation in the choice probability is solely due to variation in the covariate index $w_{t}^{\prime}\theta_{0}$. [^6] Condition 2 implies the stationarity of $v_{it}$: $v_{is}\mid W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid W_{ist}$. Therefore, as shown in Manski (1987), the identified set for $\theta_{0}$ is characterized by the condition: an increase in the value of $w_{t}^{\prime}\theta_{0}$ over time implies an increase in the choice probability over time.
However, Manski (1987)’s method is not applicable to our setting with Assumption 1, given that the endogenous regressor $X_{it}$ can be correlated with $v_{it}$. When the covariate $X_{it}$ changes, the conditional distribution of $v_{it}$ might potentially change as well. Variation in the choice probability contains a mixture of changes in the covariate index and the conditional distribution of $v_{it}$. Hence, it is not feasible to derive identifying restrictions on $\theta_{0}$ merely by examining the intertemporal variation in the choice probability.
Our analysis relies on the partial stationarity condition in Assumption 1, which implies the stationarity of $v_{it}$ given $Z_{ist}:=(Z_{is},Z_{it})$:
$v_{is}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid Z_{ist}.$
The identification strategy proceeds in two steps. We first derive lower and upper bounds for the distribution of $v_{it}\mid Z_{ist}$ at each period $t$. The derived bounds at each period $t$ are functions of observed variables and the parameter $\theta_{0}$. Since the distribution of $v_{it}\mid Z_{ist}$ is the same over time, then the bounds of the distribution across time should have intersections. This restriction serves as the identifying restriction for $\theta_{0}$. Next, we show how to derive bounds on the distribution of $v_{it}\mid Z_{ist}$.
Let $F_{v_{t}\mid W_{st}}(\cdot\mid w_{st})$ denote the conditional CDF of $v_{it}$ given $W_{ist}=w_{st}$. For any fixed point $a\in\mathcal{R}$, we first derive bounds for the conditional distribution $F_{v_{t}\mid W_{st}}(a\mid w_{st})$, and then establish bounds on $F_{v_{t}\mid Z_{st}}\left(a\mid z_{st}\right)$ by taking expectation over the covariate $X_{ist}$ given $Z_{ist}=z_{st}$.
Lower bound. The conditional choice probability given $W_{ist}=w_{st}$ is given as
$\Pr\left(Y_{it}=1\mid W_{ist}=w_{st}\right)=\Pr(v_{it}\leq w_{t}^{\prime}\theta_{0}\mid W_{ist}=w_{st})=F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta_{0}\mid w_{st}).$
For any fixed point $a$, by the monotonicity of a distribution, $F_{v_{t}\mid W_{st}}(a\mid w_{st})$ can be bounded below using the observed choice probability:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\Pr(Y_{it}=1\mid W_{ist}=w_{st})\mathbbm{1}\{w_{t}^{\prime}\theta_{0}\leq a\}.$
The above inequality holds since either $a$ is larger than the covariate index $w_{t}^{\prime}\theta_{0}$ so we know $F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta_{0}\mid w_{st})$ or the right hand side is equal to zero when $a$ is smaller than $w_{t}^{\prime}\theta_{0}$.
As Assumption 1 is on the distribution of $v_{it}$ given $Z_{ist}$, we can establish the lower bound for $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ by taking expectation over covariates $X_{ist}$ given $Z_{ist}=z_{st}$:
[…]
$=\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\leq a\mid z_{st}).$
The above lower bound only depends on observed variables $(W_{ist},Y_{it})$ and $\theta_{0}$, so it is identified up to the parameter $\theta_{0}$. This bound will help construct identifying restrictions on $\theta_{0}$.
Upper bound. The idea of constructing the upper bound is similar. For any $a\in\mathcal{R}$, the conditional distribution $F_{v_{t}\mid W_{st}}$ can be bounded above:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\leq\Pr(Y_{it}=1\mid W_{ist}=w_{st})\mathbbm{1}\{w_{t}^{\prime}\theta_{0}\geq a\}+\mathbbm{1}\{w_{t}^{\prime}\theta_{0}<a\}.$
Then, taking expectation over covariates $X_{ist}$ given $Z_{ist}=z_{st}$ leads to the upper bound for $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$:
[…]
$=\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\Pr(z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st})$
$=1-\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st}).$
Given the lower and upper bound for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, we are ready to establish identifying conditions for $\theta_{0}$. Since Assumption 1 requires that $v_{t}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{s}\mid Z_{ist}$, it implies the upper bound for the distribution of $v_{s}\mid Z_{ist}$ must be larger than the lower bound for the distribution of $v_{t}\mid Z_{ist}$ otherwise Assumption 1 cannot hold. This restriction characterizes an identified set for $\theta_{0}$, presented in the following proposition.
Proposition 1.
Under Assumption 1, an identified set $\Theta_{I,1}$ for $\theta_{0}$ is the set of parameters that satisfy the following conditions:
$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st}),$ (7)
for any $a\in\mathcal{R}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Proposition 1 characterizes an identified set for $\theta_{0}$, using the conditional joint distribution of $(Y_{it},X_{it})$ given $Z_{ist}$ across any pair of periods $(s,t)$. The identification result allows for endogeneity in the covariate $X_{ist}$ and does not rely on any excluded instruments. [^7] When a sub-vector $Z_{it}^{exc}$ of $Z_{it}$ is assumed to be excluded from the binary choice model, then the identifying restriction in (7) is adjusted by setting the coefficient of $Z_{it}^{exc}$ as zero: $\beta_{0}^{exc}=0$.
The identifying condition in (7) can be reformulated as conditional moment inequalities, by substituting the probability $\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})$ with the conditional expectation $\mathbb{E}\left[\mathbf{\mathbbm{1}}\{Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\}\mid z_{st}\right]$. Consequently, the estimation and inference for $\theta_{0}$ can be conducted using established methods such as Chernozhukov, Hong, and Tamer (2007), Andrews and Shi (2013), and Chernozhukov, Lee, and Rosen (2013).
Remark 1.
Our “partial stationarity” condition accommodates the “full stationarity” condition (2) as a special case. Under full stationarity, or in other words, when all covariates are exogenous, Proposition 1 specializes to the following “maximum-score-type” identifying restrictions in Manski (1987):
[…]
Remark 2.
In the special case where the only endogenous covariate is one lagged outcome variable, as studied in Khan, Ponomareva, and Tamer (2023), the identifying restrictions yielded by Proposition 1 become
$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta_{0}+Y_{i,s-1}\gamma_{0}\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq a\mid z_{st}),$
for any $a\in\{z_{t}^{\prime}\beta_{0},z_{t}^{\prime}\beta_{0}+\gamma_{0},z_{s}^{\prime}\beta_{0},z_{s}^{\prime}\beta_{0}+\gamma_{0}\}$. As shown in Appendix A.4, the above identifying condition on $\theta_{0}$ implies the identifying conditions in Khan, Ponomareva, and Tamer (2023, Theorem 1, p. 8-9), under different values of $(z_{s}^{\prime}\beta_{0},z_{t}^{\prime}\beta_{0},\gamma_{0})$.
Proposition 1 does not impose restrictions on the support of the endogenous covariate $X_{it}$, allowing it to be either discrete or continuous. When $X_{it}$ is continuous, then Proposition 1 requires condition (7) to hold for any $a\in\mathcal{R}$ so it exploits the whole distribution information of $(X_{it},Y_{it})$ at any value $a$. For the estimation and inference of $\theta_{0}$, we can follow the conventional inference method to discretize the space of $\mathcal{R}$ and pick $K_{n}$ points for a. In the case of discrete $X_{it}$, the dimension of $a$ can be further reduced, and the identified set $\Theta_{I,1}$ can be characterized by finite number of restrictions, shown in the following corollary.
Corollary 1.
When the endogenous covariate $X_{it}\in\{a_{1},...,a_{K}\}$ only takes finite number of values, then the identified set $\Theta_{I,1}$ is the set of parameters $\theta$ that satisfy condition (7) for any $a\in\{z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma,z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma\}_{k=1}^{K}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Corollary 1 reduces the number of restrictions in Proposition 1 to $2K$ number of moment restrictions for any pair of two periods $s,t$. For the dynamic model where $X_{it}$ includes one lagged dependent variable $X_{it}=Y_{i,t-1}\in\{0,1\}$, there are only four conditional restrictions for $\theta_{0}$. When $X_{it}$ includes $j$ lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2},...Y_{i,t-j})$, then there will be $2^{j+1}$ restrictions.
We show that $\Theta_{I,1}$ is an identified set for $\theta_{0}$, however, it is still unclear whether condition (7) exploits all the available information for $\theta_{0}$. The following theorem establishes the sharpness of the identified set $\Theta_{I,1}$ when $X_{it}$ takes finite number of values.
Theorem 1 (Sharpness).
Under Assumption 1, the identified set $\Theta_{I,1}$ is sharp when $X_{it}$ takes finite number of values for any $t\leq T$.
Proposition 1 provides identification results without restrictions on the support of $X_{it}$, and Theorem 1 shows that the results already exhausted all available information for $\theta_{0}$ when $X_{it}$ only takes finite number of values. This results implies that for dynamic models with any number of lagged dependent variables, the identified set $\Theta_{I,1}$ is the smallest set we can obtain.
2.3 Point Identification
Proposition 1 characterizes the sharp identified set for $\theta_{0}$ by only imposing Assumption 1. This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$ under support conditions on the exogenous covariate $Z_{it}$. We focus on the scenario where the endogenous covariate $X_{it}$ is discrete $X_{it}\in\mathcal{X}=\{a_{1},a_{2},...,a_{K}\}$ and there are only two periods $T=2$.
To point identify $\beta_{0}$, the first step is to determine the sign of the covariate index $(Z_{i2}-Z_{i1})^{\prime}\beta_{0}$ under certain variation of observed choice probability. To identify the sign of $(Z_{i2}-Z_{i1})^{\prime}\beta_{0}$, we define the following two sets:
$\mathcal{Z}_{1}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\Pr(Y_{i1}=0,X_{i1}=x\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=x\mid z_{12})\Big{\}},$
$\mathcal{Z}_{2}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\Pr(Y_{i1}=1,X_{i1}=x\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=x\mid z_{12})\Big{\}}.$
Let $\mathcal{Z}:=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Let $\Delta Z_{i}=Z_{i2}-Z_{i1}$ and $\Delta\mathcal{Z}$ be defined as
$\Delta\mathcal{Z}:=\Big{\{}\Delta z:=z_{2}-z_{1}\mid(z_{1},z_{2})\in\mathcal{Z}\Big{\}}.$
As shown in Appendix A.3, when $\Delta z$ satisfies $\Delta z\in\Delta\mathcal{Z}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified. In the definition of the two sets $\mathcal{Z}_{1},\mathcal{Z}_{2}$, we only need the existence of one value in the support of $\mathcal{X}$ such that the choice probability in the two sets are observed. When observing such choice probability, the sign of $\Delta z^{\prime}\beta_{0}$ is identified. Then $\beta_{0}$ can be identified up to scale under the standard large support condition on $\Delta z$.
Let $\Delta z^{j}$ denote the $j$th element of $\Delta z$. The following is the support condition on the covariate.
Assumption 2 (Support Condition).
(1) $\Delta\mathcal{Z}$ is not contained in any proper linear subspace of $\mathcal{R}^{d_{z}}$; (2) for any $\Delta z\in\Delta\mathcal{Z}$, there exists one element $\Delta z^{j^{*}}$ such that $\beta_{0}^{j*}\neq 0$, and the conditional support of $\Delta z^{j^{*}}$ is $\mathcal{R}$ given $\Delta z\setminus\Delta z^{j^{*}}$, where $\Delta z\setminus\Delta z^{j^{*}}$ denote the remaining components of $\Delta z$ besides $\Delta z^{j^{*}}$.
Proposition 2.
Under Assumptions 1-2, the parameter $\beta_{0}$ is point identified up to scale.
We provide point identification for $\beta_{0}$ with two periods $T=2$. When there are more than two periods, then we only require the existence of two periods, satisfying Assumption 2. As shown in Manski (1987), the large support assumption is necessary to point identify $\beta_{0}$, as without it, there exists some $b\neq k\beta_{0}$ such that $\Delta z^{\prime}b$ has the same sign with $\Delta z^{\prime}\beta_{0}$ when $\Delta z$ has bounded support.
The parameter $\gamma_{0}$ in general can be only partially identified given potential endogeneity of $X_{it}$ and flexible structures on $(\alpha_{i},\epsilon_{it})$. Nevertheless, we can still bound the value $(x_{1}-x_{2})^{\prime}\gamma_{0}$ and identify the sign of $\gamma_{0}$ under certain choice probabilities. We derive the sufficient conditions to identify the sign of $\gamma_{0}$.
Let $x^{j}$ denote the $j$-th element of $x$ and $\gamma_{0}^{j}$ denote the $j$-th coefficient of $\gamma_{0}$. We define the following two sets:
[…]
$1-\Pr(Y_{i1}=0,X_{i1}=x_{1}\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=x_{2}\mid z_{12})\Big{\}};$
[…]
$1-\Pr(Y_{i1}=1,X_{i1}=x_{1}\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=x_{2}\mid z_{12})\Big{\}}.$
From the identifying results in Proposition 1, the value of $(x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}$ can be bounded when $(z_{1},z_{2})$ belong to the two sets:
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{3}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0},$
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{4}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}.$
Then the sign of $\gamma_{0}^{j}$ is identified if either the sign of $\Delta z^{\prime}\beta_{0}$ is identified as negative when $(z_{1},z_{2})\in\mathcal{Z}_{2}$ or as positive when $(z_{1},z_{2})\in\mathcal{Z}_{1}$.
Proposition 3.
Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}^{j}_{3}\cap\mathcal{Z}_{2}\neq\emptyset$ or $\mathcal{Z}^{j}_{4}\cap\mathcal{Z}_{1}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
When the endogenous variable $X_{it}$ is a scalar, e.g., the lagged dependent variable $X_{it}=Y_{i,t-1}$, then the definition of the two sets $\mathcal{Z}^{j}_{3},\mathcal{Z}^{j}_{4}$ can be simplified as there existing $x_{1}\neq x_{2}$ such that the corresponding choice probability is observed. Besides the sign of $\gamma_{0}$, the identification results can also bound the value of $\gamma_{0}$ from variation in the exogenous covariates.
When $X_{it}$ is multi-dimensional such as including two lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2})$ with $\gamma_{0}=(\gamma_{0}^{1},\gamma_{0}^{2})$, then $\gamma_{0}^{1}$ is identified when the required choice probability in the two sets $\mathcal{Z}^{1}_{3},\mathcal{Z}^{1}_{4}$ are observed for $(Y_{i,1},Y_{i,0})=(1,1),(Y_{i,2},Y_{i,1})=(0,1)$ or $(Y_{i,1},Y_{i,0})=(0,0),(Y_{i,2},Y_{i,1})=(1,0)$. We provide general sufficient conditions to identify the sign of $\gamma_{0}$, which may be stronger than necessary and can be relaxed in certain scenarios. For example, when we know that $\gamma_{0}^{1}+\gamma_{0}^{2}>0$ while $\gamma_{0}^{1}<0$, we can infer that $\gamma_{0}^{2}>0$ without requiring additional assumptions on the two sets $\mathcal{Z}^{2}_{3},\mathcal{Z}^{2}_{4}$.
3 Ordered Response Model
3.1 Model and Identification
Consider a setting where the outcome variable $Y_{it}$ is discrete and takes $J$ values: $Y_{it}\in\{y_{1},..,y_{J}\}$. For example, $Y_{it}$ can be different categories of income, health outcomes, or educational attainment. We study the following panel ordered response model:
$Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it},$ (8)
$Y_{it}$ $=\sum_{j=1}^{J}y_{j}\mathbbm{1}\{b_{j}<Y_{it}^{*}\leq b_{j+1}\},$
where $Y_{it}^{*}$ denotes the latent dependent variable, $b_{1}=-\infty,b_{J+1}=+\infty$, and the remaining threshold parameters $b_{j}$ (where $b_{j}\leq b_{j+1}$) can be either known or unknown for $12\leq j\leq J-$ 1. The binary choice model in (1) is nested with $J=2$ and $b_{2}=0$. Similar to model (1), $X_{it}$ is the potentially endogenous covariate that can be arbitrarily dependent on $(\alpha_{i},\epsilon_{it})$, such as the lagged dependent variable $Y_{i,t-1}$, while $Z_{it}$ is the exogenous covariate that satisfies Assumption 1:
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
We now show how our key identification strategy based on partial stationarity can also be exploit to partially identify $\theta_{0}$. Similar to Section 2, we still seek to establish both lower and upper bounds for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, where $v_{it}:=\epsilon_{it}+\alpha_{i}$. The distinction is that in the ordered choice setting we can obtain $J$ different (upper and lower) bounds, which can then be aggregated over to form tighter bounds as shown below.
1. Lower bound. By the monotonicity of a distribution, we first derive a lower bound for the distribution $v_{t}\mid w_{st}$ conditional on all covariates $W_{ist}=w_{st}$, given as
$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $=\Pr(v_{t}\leq a\mid w_{st})$
$\geq\Pr(v_{t}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
where the above lower bound holds for any choice $j\leq J$. Since the interval $(-\infty,b_{j+1}-w_{t}^{\prime}\theta_{0}]$ can be expressed as the union of multiple intervals: $(-\infty,b_{j+1}-w_{t}^{\prime}\theta_{0}]=\bigcup\limits_{k=1}^{j}(b_{k}-w_{t}^{\prime}\theta_{0},b_{k+1}-w_{t}^{\prime}\theta_{0}]$, the above lower bound can be also written as
$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $\geq\sum_{k=1}^{j}\Pr(b_{k}-w_{t}^{\prime}\theta_{0}<v_{t}\leq b_{k+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}$
$=\sum_{k=1}^{j}\Pr(Y_{it}=y_{k}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
for any $j\leq J$. Fixing the constant $a$, there exists a maximum choice $j^{*}_{a}$, defined as $j^{*}_{a}:=\max\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}$. Then, the largest lower bound is the sum of choice probabilities up to the maximum choice $j^{*}_{a}$:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\sum_{k=1}^{j^{*}_{a}}\Pr(Y_{it}=y_{k}\mid w_{st})\mathbbm{1}\{b_{j^{*}_{a}+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$
An equivalent expression for this lower bound is given as:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
as the indicator function is zero $\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}=0$ for any choice $j>j^{*}_{a}$.
2. Upper bound. Similarly, we can derive the upper bound for $F_{v_{t}\mid w_{st}}(a\mid w_{st})$ using the sum of conditional choice probabilities, given as follows:
$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $=1-\Pr(v_{t}>a\mid w_{st})$
$\leq 1-\sum_{j=1}^{J}\Pr(b_{j}-w_{t}^{\prime}\theta_{0}<v_{t}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j}-w_{t}^{\prime}\theta_{0}\geq a\}$
$\leq 1-\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{t}^{\prime}\theta_{0}\geq a\}.$
Then, the upper bound for the distribution $F_{v_{t}\mid z_{st}}(a\mid z_{st})$ is derived as follows by taking expectation over covariate $X_{ist}$ given $z_{st}$:
$F_{v_{t}\mid z_{st}}(a\mid z_{st})\leq 1-\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j}-z_{t}^{\prime}\beta_{0}-X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st}).$
Given the established lower and bounds on the distribution of $v_{t}\mid z_{st}$, an identified set for $\theta_{0}$ is characterized in the following proposition.
Proposition 4.
Under Assumption 1, the identified set $\Theta_{I,2}$ for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following conditions:
$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st}),$ (9)
for any $a\in\mathcal{R}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Proposition 4 characterizes the identified set for $\theta_{0}$ for the general ordered response model with potential endogeneity. This result allows for both static and dynamic models, as discussed in Section 3.2. In comparison to the binary choice model, Proposition 4 aggregates information from all $J$ ordered responses across different time periods to identify $\theta_{0}$. When there are only two choices ($J=2$), this result simplifies to the identifying condition (7) in Proposition 1, with $y_{1}=0,y_{2}=1$ and changing $a$ to $-a$. Again, the result in Proposition 4 accommodates both discrete and continuous endogenous covariates. When the endogenous $X_{it}$ only takes a finite number of values, such as lagged dependent variable, we can reduce the number of restrictions in Proposition 4 to a finite number, as shown in the next corollary.
Corollary 2.
Under Assumption 1 and $X_{it}$ only takes finite number of values $X_{it}\in\{a_{1},...,a_{K}\}$, the identified set $\Theta_{I,2}$ is the set of parameters $\theta$ that satisfy condition (9) for any $a\in\mathcal{Q}_{2}:=\{b_{j}-z_{s}^{\prime}\beta-a_{1}^{\prime}\gamma,...,b_{j}-z_{s}^{\prime}\beta-a_{K}^{\prime}\gamma,b_{j}-z_{t}^{\prime}\beta-a_{1}^{\prime}\gamma,...,b_{j}-z_{t}^{\prime}\beta-a_{K}^{\prime}\gamma\}_{j=2}^{J}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Similar to Corollary 1, Corollary 2 reduces the number of conditional moment inequalities to $2K\times(J-1)$ numbers for any pair of two periods. For the dynamic model with one lagged dependent variable $X_{it}=Y_{i,t-1}\in\{y_{1},...,y_{J}\}$, then there will be $2J(J-1)$ number of restrictions.
3.2 Applications
In this section, we apply the results in Proposition 4 to explore two panel ordered choice model. The first one examines a static model where all covariates are exogenous, while the other studies a dynamic model with one lagged dependent variable. To the best of our knowledge, these two scenarios have not been explored in the literature within the framework of the stationarity assumption.
Static model: we consider a static model where Assumption 1 holds conditional on all covariates $W_{ist}$. The identifying restriction in Proposition 4 is given as
$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{s}^{\prime}\theta_{0}\geq a\}\\ \geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$
We can further simplify the above condition by getting rid of the parameter $a$ and transform it into finite number of conditional moment inequalities. The following lemma presents the identified set for $\Theta_{I,2}$ for this static model.
Lemma 1.
Assume that $\epsilon_{is}\mid(W_{ist},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{ist},\alpha_{i})$, the identified set $\Theta_{I,2}$ for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following conditions:
$b_{j_{1}+1}-w_{s}^{\prime}\theta\geq b_{j_{2}+1}-w_{t}^{\prime}\theta\Longrightarrow\sum_{j=1}^{j_{1}}\Pr(Y_{is}=y_{j}\mid w_{st})\geq\sum_{j=1}^{j_{2}}\Pr(Y_{it}=y_{j}\mid w_{st}),$
for any $11\leq j_{1},j_{2}\leq J-$ 1, $s,t\leq T$, and $w_{st}\in\mathcal{R}^{d}\times\mathcal{R}^{d}$.
In the static model, variation in conditional choice probability only comes from changes in the covariate $W_{it}$. Thus, we can directly establish the relationship between the variation in choice probability and the changes in covariates over time, eliminating the parameter $a$. Different from the binary choice model, the results in Lemma 1 also exploit intertemporal variation in the sum of multiple choices rather than investigating a single choice. Moreover, we can utilize variations in the sum of different choices across various periods to identify $\theta_{0}$.
Besides the static model, the results in Proposition 4 can also be applied to study a dynamic ordered response model, where people’s choice at the current period ($t$) can depend on their choice in the last period ($t-1$).
Dynamic model: consider the following dynamic ordered response model with one lagged dependent variable:
$Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+Y_{it-1}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\sum_{j=1}^{J}y_{j}\mathbbm{1}\{b_{j}<Y_{it}^{*}\leq b_{j+1}\}.$
In this example, the endogenous covariate is the lagged dependent variable $X_{it}=Y_{it-1}\in\{y_{1},...,y_{J}\}$. The identifying restriction in Proposition 4 is presented as follows:
$1-\sum_{j=1}^{J}\Pr\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\geq a\mid z_{st}\right)\\ \geq\sum_{j=1}^{J}\Pr\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq a\mid z_{st}\right),$
for any $a\in\{b_{j}-z_{s}^{\prime}\beta-y_{1}\gamma,...,b_{j}-z_{s}^{\prime}\beta-y_{J}\gamma,b_{j}-z_{t}^{\prime}\beta-y_{1}\gamma,...,b_{j}-z_{t}^{\prime}\beta-y_{J}\gamma\}_{j=2}^{J}$.
For the dynamic model, the identified set $\Theta_{I,2}$ is characterized by $2J(J-1)$ number of conditional inequalities for any pair of two periods $(s,t)$. This approach can also allow for dynamic models with more than one lagged dependent variable, such as $X_{it}=(Y_{i,t-1},Y_{i,t-2})$. Furthermore, it can accommodate dynamic models that allow for heterogeneous effects from different choices in the last period, as given below:
$Y_{it}^{*}=Z_{it}^{\prime}\beta_{0}+\sum_{j=2}^{J}\mathbf{\mathbbm{1}}\{Y_{it-1}=j\}\gamma_{0,j}+\alpha_{i}+\epsilon_{it}.$
3.3 Point Identification
This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$. Similar to Section 2.3, we still focus on the scenario where the endogenous covariate $X_{it}$ only takes finite number of values $X_{it}\in\mathcal{X}=\{a_{1},a_{2},...,a_{K}\}$, and we illustrate the concept using two periods ($T=2$).
To identify $\beta_{0}$, we define the following two sets:
[…]
$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x\mid z_{12})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Big{\}};$
[…]
$1-\sum_{j=1}^{k-1}\Pr(Y_{i1}=y_{j},X_{i1}=x\mid z_{12})<\sum_{j=k}^{J}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Big{\}}.$
Let $\mathcal{Z}_{order}:=\mathcal{Z}_{1,order}\cup\mathcal{Z}_{2,order}$. Let $\Delta Z_{i}=Z_{i2}-Z_{i1}$ and $\Delta\mathcal{Z}_{order}$ be defined as
$\Delta\mathcal{Z}_{order}:=\Big{\{}\Delta z:=z_{2}-z_{1}\mid(z_{1},z_{2})\in\mathcal{Z}_{order}\Big{\}}.$
Assumption 3 (Support Condition).
(1) $\Delta\mathcal{Z}_{order}$ is not contained in any proper linear subspace of $\mathcal{R}^{d_{z}}$; (2) for any $\Delta z\in\Delta\mathcal{Z}_{order}$, there exists at least one least one element $\Delta z^{j^{*}}$ with $\beta_{0}^{j*}\neq 0$, and the conditional support of $\Delta z^{j^{*}}$ is $\mathcal{R}$ given $\Delta z\setminus\Delta z^{j^{*}}$, where $\Delta z\setminus\Delta z^{j^{*}}$ denote the remaining components of $\Delta z$ besides $\Delta z^{j^{*}}$.
Proposition 5.
Under Assumptions 1 &\&& 3, the parameter $\beta_{0}$ is point identified up to scale.
Similar to Proposition 2, the strategy is to first identify the sign of $\Delta z^{\prime}\beta_{0}$ from certain variation in the choice probability, as defined in the two sets $(\mathcal{Z}_{1,order},\mathcal{Z}_{2,order})$. Given the sign of $\Delta z^{\prime}\beta_{0}$ is identified, then the parameter $\beta_{0}$ is point identified up to scale under the conventional large support condition on $\Delta z$. For the static model (with $\gamma_{0}=0$), all coefficients $\beta_{0}$ are point identified under the support condition.
To identify the sign of $\gamma_{0}$, we define the following two sets: for $j\leq d_{x}$,
[…]
$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}\mid z_{12})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Big{\}};$
[…]
$1-\sum_{j=1}^{k-1}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}\mid z_{12})<\sum_{j=k}^{J}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Big{\}}.$
Proposition 6.
Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}\neq\emptyset$ or $\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
For the dynamic model with one lagged dependent variable $X_{it}=Y_{i,t-1}$, then the joint choice probabilities defined in the two sets $\mathcal{Z}^{j}_{3,order},\mathcal{Z}^{j}_{4,order}$ can be simplified as taking different values for the lagged dependent variable: $Y_{i0}=y_{j_{1}},Y_{i1}=y_{j_{2}}$ with $y_{j_{1}}\neq y_{j_{2}}$. The identifying restriction in Proposition 4, in general, can bound the value of $(y_{j}-y_{k})\gamma_{0}$ using the variation in $Z_{it}$. Proposition 6 ensures that its bound excludes zero for some $(y_{j},y_{k})$, thereby identifying the sign of $\gamma_{0}$. When there are more than two periods ($T>2$), then point identification of $\beta_{0}$ and the sign of $\gamma_{0}$ is achieved when there exists two periods such that the assumptions are satisfied.
4 Multinomial Choice Model
4.1 Model and Identification
This section applies our key identification strategy to panel multinomial choice model with endogeneity. Specifically, consider a set of choice alternatives $\mathcal{J}=\{0,1,...,J\}$. Let $u_{ijt}$ denote the utility for individual $i$ for choice $j$ at time $t$, which depends on three components: observed covariate $W_{ijt}=(Z_{ijt}^{\prime},X_{ijt}^{\prime})^{\prime}$, unobserved fixed effects $\alpha_{ij}$, and unobserved time-varying preference shock $\epsilon_{ijt}$. The utility of choice 0 (outside option) is normalized to zero: $u_{i0t}=0$. Let $Y_{it}\in\mathcal{J}$ denote individual $i$’s choice at time $t$. We study the following panel multinomial choice model:
$u_{ijt}$ $=Z_{ijt}^{\prime}\beta_{0}+X_{ijt}^{\prime}\gamma_{0}+\alpha_{ij}+\epsilon_{ijt},$
$Y_{it}$ $=\arg\max_{j\in\mathcal{J}}u_{ijt},$
where $X_{ijt}$ denotes the potentially endogenous covariate, such as the lagged dependent variable or endogenously determined prices, and $Z_{ijt}$ denotes the exogenous covariates that satisfies the partial stationarity condition (Assumption 1)
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
with $Z_{it}=\{Z_{ijt}\}_{j\in\mathcal{J}}$,$\alpha_{i}=\{\alpha_{ij}\}_{j\in\mathcal{J}}$ and $\epsilon_{it}=\{\epsilon_{ijt}\}_{j\in\mathcal{J}}$ defined to collect terms across all $J$ choice alternatives.
The identification of $\theta_{0}$ has been studied under the standard full stationarity condition given all covariates $W_{ist}$ in different models, including Pakes and Porter (2022), Shi, Shum, and Song (2018), Gao and Li (2020), and Wang (2022). The main focus of this paper is to derive identification for $\theta_{0}$, while allowing for the endogenous covariate $X_{it}$. As discussed in Section 4.2, our approach also accommodates the standard stationarity assumption, and the results are consistent with those in Pakes and Porter (2022).
Let $v_{ijt}=\alpha_{ij}+\epsilon_{ijt}$ and $v_{it}=\{v_{ijt}\}_{j\in\mathcal{J}}$. The identification strategy is still to establish bounds on the distribution of $v_{it}\mid Z_{ist}$ and derive identifying restrictions on $\theta_{0}$ based on partial stationarity. Here, the distinction is that the error term $v_{it}$ is multi-dimensional instead of one-dimensional in binary or ordered choice models. Moreover, consumers’ choice involves the comparison between different choices.
The observed choice probability of selecting $j$ at period $t$ is given as
$\Pr(Y_{it}=j\mid w_{st})=\Pr\left(w_{jt}^{\prime}\theta_{0}+v_{ijt}\geq w_{kt}^{\prime}\theta_{0}+v_{ikt}\quad\forall k\in\mathcal{J}\mid w_{st}\right).$
The above probability restricts the distribution $\Pr(v_{ikt}-v_{ijt}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \forall k\neq j\mid w_{st})$. From observed choice probability, we can derive bounds for the distribution $\Pr(v_{ikt}-v_{ijt}\leq a_{jk}\ \forall k\neq j\mid z_{st})$ as a function of $\theta_{0}$ and observed variable $(X_{it},Z_{it},Y_{it})$. Therefore, the identifying condition for $\theta_{0}$ is characterized by the restriction that the bounds over different periods must intersect. More generally, we observe the conditional probability $\Pr(Y_{it}\in K\mid w_{st})$ for any set $K\subset\mathcal{J}$ and can bound the probability $\Pr(v_{ikt}-v_{ijt}\leq a_{jk}\ \exists j\in K,\forall k\in K^{c}\mid z_{st})$, where $K^{c}:=\mathcal{J}\setminus K$. In the case where $K$ is a singleton ($K=\{j\}$), it then becomes the conditional probability of selecting a single choice $j$. The following proposition characterizes the identification results for $\theta_{0}$.
Proposition 7.
Under Assumption 1, an identified set $\Theta_{I,3}$ for $\theta_{0}$ is the set of parameters $\theta=(\beta,\gamma)$ that satisfy the following condition:
$1-\Pr(Y_{is}\in K^{c},(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st})\\ \geq\Pr(Y_{it}\in K,(z_{jt}-z_{kt})^{\prime}\beta+(X_{ijt}-X_{ikt})^{\prime}\gamma\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st}),$ (10)
for any subset $K\subset\mathcal{J}$, any $a_{jk}\in\mathcal{R}$ with $j\in K$ and $k\in K^{c}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Corollary 3.
Under Assumption 1, and when $X_{ijt}-X_{ikt}$ only takes finite number of values $X_{ijt}-X_{ikt}\in\mathcal{X}_{jk}:=\{x_{jk,1},...,x_{jk,M}\}$, then the identified set $\Theta_{I,3}$ is the set of parameters $\theta$ that satisfy condition (10) for any subset $K\subset\mathcal{J}$, any $a_{jk}\in\{(z_{js}-z_{ks})^{\prime}\beta+x_{jk}^{\prime}\gamma,(z_{jt}-z_{kt})^{\prime}\beta+x_{jk}^{\prime}\gamma\}_{x_{jk}\in\mathcal{X}_{jk}}$ with $j\in K$ and $k\in K^{c}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Proposition 7 provides general identification results for multinomial choice models with discrete or continuous endogenous covariates. Corollary 3 further reduces the number of restrictions in Proposition 7 to a finite number when the difference in the endogenous covariate between choices, $X_{ijt}-X_{ikt}$, only takes finite number of values. The results accommodate both static models and dynamic models with any number of lagged dependent variables. Point identification can be achieved under conditions similar to those presented in Section 2.3 and 3.3, so further analysis is omitted here.
4.2 Applications
We examine the two applications of Proposition 7 for the panel multinomial choice model. The first one studies the static model, where all covariates $W_{it}$ are exogenous ($\epsilon_{is}\mid W_{ist},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{ist},\alpha_{i}$), which is the same setting as Pakes and Porter (2022). In this scenario, we show that Proposition 7 boils down to the identifying results in Pakes and Porter (2022). Another one explores the dynamic multinomial choice model, incorporating one lagged dependent variable $Y_{i,t-1}$.
Static model: Pakes and Porter (2022) characterizes the sharp identified set for $\theta_{0}$ under the stationarity assumption given all covariates: $\epsilon_{is}\mid W_{is},W_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{is},W_{it},\alpha_{i}$. Now, we show that the identifying restriction in (10) is consistent with the results in Pakes and Porter (2022).
Under full stationarity (given all covariates $W_{it}$), condition (10) becomes
$1-\Pr(Y_{is}\in K^{c},(w_{js}-w_{ks})^{\prime}\theta_{0}\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid w_{st})\\ \geq\Pr(Y_{it}\in K,(w_{jt}-w_{kt})^{\prime}\theta_{0}\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid w_{st}).$
The above equation is only informative when $a_{jk}\in[(w_{jt}-w_{kt})^{\prime}\theta_{0},(w_{js}-w_{ks})^{\prime}\theta_{0}]$ for any $j\in K,k\in K^{c}$; otherwise either the upper bound becomes one or the lower bound becomes zero in the above restriction. When $a\in[(w_{jt}-w_{kt})^{\prime}\theta_{0},(w_{js}-w_{ks})^{\prime}\theta_{0}]$, the above restriction is equivalent to the following condition: for any $K\subset\mathcal{J}$,
$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\geq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in K^{c}\\ \Longrightarrow 1-\Pr(Y_{is}\in K^{c}\mid w_{st})=\Pr(Y_{is}\in K\mid w_{st})\geq\Pr(Y_{it}\in K\mid w_{st}),$
which aligns with the identification result in Pakes and Porter (2022) (Proposition 1, P. 12).
Dynamic model: Proposition 7 is also applicable to the following dynamic panel multinomial choice model:
$u_{ijt}=Z_{ijt}^{\prime}\beta_{0}+\mathbbm{1}\{Y_{i,t-1}=j\}\gamma_{0}+\alpha_{ij}+\epsilon_{ijt}.$
The above model allows individuals’ current utility at time $t$ to depend on their choices at the last period $t-1$. In this model, the endogenous variable $X_{ijt}$ is the lagged dependent variable $Y_{i,t-1}$, and the difference of the endogenous covariate between choices only takes three values: $X_{ijt}-X_{ikt}\in\mathcal{X}_{jk}:=\{1,-1,0\}$. Then the identified set for $\Theta_{I,3}$ is characterized by the condition in Corollary 2 with $a_{jk}\in\{(z_{js}-z_{ks})^{\prime}\beta+\gamma,(z_{js}-z_{ks})^{\prime}\beta-\gamma,(z_{js}-z_{ks})^{\prime}\beta,(z_{jt}-z_{kt})^{\prime}\beta+\gamma,(z_{jt}-z_{kt})^{\prime}\beta-\gamma,(z_{jt}-z_{kt})^{\prime}\beta\}$.
5 Censored Dependent Variable Model
The previous sections primarily investigate discrete choice models, while our approach also applies to models with continuous dependent variables, including those with censored or interval outcomes. To illustrate, we focus on censored outcomes models below.
Consider the following panel models with censored outcomes:
$Y^{*}_{it}$ $=Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\max\{Y^{*}_{it},0\},$
where $X_{it}$ represents the endogenous covariate, $Z_{it}$ stands for the exogenous covariate that satisfies Assumption 1, $Y^{*}_{it}$ denotes the latent outcome not observed in the data, and $Y_{it}$ represents the observed outcome, which is censored at zero. The threshold for censoring can be replaced with other nonzero constants. The identification strategy is still to derive conditions for $\theta_{0}$ based on the stationarity assumption. Here the conditional distribution of $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ is not identified since we only observe the censoring outcome $Y_{it}$. While we can still bound the conditional distribution $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ using variation in the joint distribution $(Y_{it},Z_{it},X_{it})$. The following proposition presents the identification results of $\theta_{0}$ with censoring outcomes.
Proposition 8.
Under Assumption 1, an identified set for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following condition:
$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a\mid z_{st}),$
for any $a\in\mathcal{R}$, any $(s,t)\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Similar to discrete choice models studied in previous sections, Proposition 8 characterizes an identified set for $\theta_{0}$ by exploiting the variation in the joint distribution $(Y_{it},X_{it})\mid Z_{ist}$ over time and the variation in the exogenous covariates $Z_{ist}$. The bounds on the distribution $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ can be derived either from the probability $\Pr(0<Y_{it}\leq y\mid z_{st})$ or $\Pr(Y_{it}=0\mid z_{st})$, depending on the value of the covariate index $z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma$. This result still accommodates both static and dynamic models with censored outcomes.
Static model: consider that the standard stationarity assumption holds, where $\epsilon_{it}\mid\alpha_{i},W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{ist}$. Then, the identifying condition in Proposition 8 is given as
$\Pr(0<Y_{is}\leq w_{s}^{\prime}\theta-a)+\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$ (11)
The above restriction is informative only when $w_{t}^{\prime}\theta\geq a$, otherwise the lower bound becomes zero. We discuss two cases for the constant $a$. When $w_{s}^{\prime}\theta\leq a\leq w_{t}^{\prime}\theta$, then condition (11) becomes
$\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$
When $a$ satisfies $a\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$, condition (11) transforms into
$\Pr(0<Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})+\Pr(Y_{is}=0\mid w_{st})=\Pr(Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})\\ \geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$
Since the above condition needs to hold for any $(s,t)$ and is symmetric in $(s,t)$, it becomes equalities after exchanging $s$ and $t$.
The following lemma derives the identified set for $\theta_{0}$ using both conditional moment inequalities and equalities.
Lemma 2.
Assuming that $\epsilon_{it}\mid\alpha_{i},W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{ist}$, an identified set for $\theta_{0}$ is characterized by the set of parameters $\theta$ that satisfy the following conditions:
$&\text{If }w_{s}^{\prime}\theta\leq a\leq w_{t}^{\prime}\theta\Longrightarrow\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}),\\ &\text{If }a\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}\Longrightarrow\Pr(Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})=\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}),\end{aligned}\right.$
for any $s,t\leq T$ and any $w_{st}\in\mathcal{R}^{d}\times\mathcal{R}^{d}$.
Dynamic model: Proposition 8 also accommodates the following dynamic model with the lagged outcome $Y_{i,t-1}$:
$Y^{*}_{it}$ $=Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\max\{Y^{*}_{it},0\}.$
In this model, since the endogenous variable $X_{it}=Y_{i,t-1}\in[0,\infty)$ can be continuous, we are not able to further simplify the identifying condition in Proposition 8. Appendix A.10 also studies dynamic models with the latent lagged outcome $Y^{*}_{i,t-1}$. Consequently, the results in Proposition 8 need to be adjusted as the endogenous variable $X_{it}=Y^{*}_{i,t-1}$ is not observed.
6 Generalization to Nonseparable Models
The key idea underlying our identification strategy generalizes further beyond the models considered before. In this section, we show that our approach relies only on two conditions: partial stationarity and “index monotonicity.” Additional model specifications and restrictions, such as the scalar-additivity of fixed effects and time-varying errors, are not needed for our identification strategy.
To illustrate, consider the following nonseparable semiparametric model:
$Y_{it}$ $=G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right),$
where $\alpha_{i}$ is the individual fixed effect of arbitrary dimension, $\epsilon_{it}$ is the time-varying error of arbitrary dimension, $W_{it}$ is a vector of observable covariates, $\theta_{0}\in\mathcal{R}^{d}$ is a conformable vector of parameter, and the function $G$ is allowed to be unknown, nonseparable but assumed to be weakly monotone in the the parametric index:
Assumption 4 (Monotonicity).
The mapping $\delta\longmapsto G\left(\delta,\alpha,\epsilon\right)$ is weakly increasing in $\delta\in\mathcal{R}$ for each realization of $\left(\alpha,\epsilon\right)$.
Note that, by setting $\alpha_{i},\epsilon_{it}$ to be scalar-valued, and $G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\mathbf{\mathbbm{1}}\left\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\}$, we obtain the dynamic binary choice model in Section 2, where $G$ is by construction weakly increasing in $W_{it}^{{}^{\prime}}\theta_{0}$.
As before, we decompose $W_{it}$, and correspondingly $\theta_{0}$, into two components, $W_{it}=\left(Z_{it}^{{}^{\prime}},X_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $\theta_{0}=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$so that
$W_{it}^{{}^{\prime}}\theta_{0}=Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},$
where $Z_{it}$ denotes exogenous covariates while $X_{it}$ denotes endogenous covariates, with the precise definition of exogeneity encoded by the partial stationarity in Assumption 1.
Now we show how partial stationarity can be exploited in conjunction with weak monotonicity (Assumption 4) to obtain identifying restrictions in the presence of endogeneity.
Let $\mathcal{Y}$ denote the support of $Y_{it}$. For any $a\in\mathcal{R}$ and $y\in\mathcal{Y}$, conditional on $W_{ist}=w_{st}$, we can bound $\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$ below as
$\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$ $\geq\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}\leq a\right\}\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$
$\geq\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}\leq a\right\}\mathbf{\mathbbm{1}}\left\{G\left(w_{t}^{\prime}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\},$
and above as
$\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$
[…]
=\ = $1-\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}>a\right\}\mathbf{\mathbbm{1}}\left\{G\left(w_{t}^{\prime}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)<y\right\}.$
Then, the conditional probability $\mathbb{P}\left(G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\mid w_{st}\right)$ given $W_{ist}=w_{st}$ can be bounded as follows:
[…]
Taking conditional expectation of the above with respect to $X_{ist}$ given $Z_{ist}=z_{st}$ yields
$\mathbb{P}\left(Y_{is}\geq y,W_{is}^{\prime}\theta_{0}\leq a\mid z_{st}\right)$
[…]
The key difference of (6) relative to the corresponding identifying restrictions in previous sections lies in that the “middle term” in (6) is no longer the conditional CDF of $\alpha_{i}+\epsilon_{it}$, but a conditional probability about the event $G\left(a,\,\alpha_{i},\,\epsilon_{is}\right)\geq y$, with the latter representation not dependent on scalar-additivity of fixed effects and time-varying errors.
We summarize the identifying restriction above by the following proposition:
Proposition 9.
Under Assumptions 1 and 4, the following condition holds:
$1-\mathbb{P}\left(Y_{it}<y,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}>a\mid z_{st}\right)\geq\mathbb{P}\left(Y_{is}\geq y,z_{s}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\leq a\mid z_{st}\right),$ (12)
for any $a\in\mathcal{R},$ $y\in\mathcal{Y}$, $s\neq t\leq T$, and $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Note that in the binary choice setting of Section 2, it suffices to set $y=1$ in (12), which then coincides with the identifying condition (7) in Proposition 1. It is also straightforward to show that our identification results in Lemma 9 can be adapted to the ordered response model in Section 3, the multinomial choice model in Section 4, and the censored outcome model in Section 5 without the scalar additive specifications.
7 Simulation
This section examines the finite sample performance of our identification approaches using Monte Carlo simulations. Since the literature has extensively studied binary choice models, we focus on the static and dynamic ordered choice models explored in Section 3 as examples to illustrate the approach. We implement the kernel-based CLR inference approach proposed in the papers by Chernozhukov, Lee, and Rosen (2013) and Chen and Lee (2019), developed to construct confidence interval based on general conditional moment inequalities.
7.1 Static Ordered Choice Model
This section explores a static ordered choice model with three choices $Y_{it}\in\{1,2,3\}$. We consider the following two-period model with $T=2$, and the latent dependent variable $Y_{it}^{*}$ is generated as:
$Y_{it}^{*}=Z_{it}^{1}\beta_{01}+Z_{it}^{2}\beta_{02}+\alpha_{i}+\epsilon_{it},$
[…]
The observed dependent variable $Y_{it}$ is given as
$Y_{it}=1*(Y_{it}^{*}\leq b_{2})+2*(0\leq Y_{it}^{*}\leq b_{2})+3*(Y_{it}^{*}>b_{3}),$
where $b_{2}=-1$ and $b_{3}=1$.
We consider the covariates $Z_{it}$ to be exogenous and Lemma 1 characterizes the identified set for $\beta_{0}$ using the following conditional moment inequalities:
$E[g(Z_{ist},Y_{ist};\beta_{0})\mid z_{st}]\geq 0,$
where
$g(Z_{ist},Y_{ist};\beta_{0})=\left\{\begin{aligned} &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\}).\end{aligned}\right.$
The first element $\beta_{01}$ of the parameter $\beta_{0}$ is normalized to one, and we are interested in conducting inference for the parameter $\beta_{02}$ using the CLR approach. Tables 1 and 2 report the average confidence interval (CI) for $\beta_{02}$, the coverage probability (CP), the average length of the CI (length), the power of the test at zero (power), and the mean absolute deviation of the lower bound ($l_{MAD}$) and upper bound ($u_{MAD}$) of the CI.
Table 1: Performance of $\beta_{02}$ under different values of $\sigma_{z}$ ($\rho=0.25$)
$\sigma_{z}$ CI CP length Power $l_{MAD}$ $u_{MAD}$
$N=2000$
$\sigma_{z}=1$ [0.537, 1.760] 0.876 1.222 1.000 0.476 0.784
$\sigma_{z}=1.5$ [0.556, 1.768] 0.934 1.212 1.000 0.454 0.773
$\sigma_{z}=2$ [0.567, 1.791] 0.950 1.224 1.000 0.440 0.796
$N=8000$
$\sigma_{z}=1$ [0.570, 1.532] 0.939 0.962 1.000 0.439 0.548
$\sigma_{z}=1.5$ [0.607, 1.561] 0.975 0.954 1.000 0.398 0.563
$\sigma_{z}=2$ [0.618, 1.571] 0.985 0.953 1.000 0.383 0.573
Table 2: Performance of $\beta_{02}$ under different values of $\rho$ ($\sigma_{z}=1$)
$\rho$ CI CP length Power $l_{MAD}$ $u_{MAD}$
$N=2000$
$\rho=0$ [0.537, 1.755] 0.895 1.218 1.000 0.476 0.773
$\rho=0.25$ [0.537, 1.760] 0.876 1.222 1.000 0.476 0.784
$\rho=0.5$ [0.511, 1.765] 0.909 1.254 1.000 0.497 0.785
$N=8000$
$\rho=0$ [0.584, 1.553] 0.933 0.969 1.000 0.436 0.568
$\rho=0.25$ [0.570, 1.532] 0.939 0.962 1.000 0.439 0.548
$\rho=0.5$ [0.573, 1.526] 0.934 0.954 1.000 0.442 0.541
As shown in Tables 1 and 2, our approach exhibits robust performance across various specifications of standard deviation $\sigma$ and correlation coefficients $\rho$. The coverage probabilities of the 95% confidence interval (CI) for $\beta_{02}$ are close to the nominal level, the length of the CI is reasonably small, and the CI consistently excludes zero. When the sample size increases, there is a significant decrease in CI length, an improvement in coverage probability, and a reduction of the mean absolute deviation (MAD) for the lower and upper bounds of the CI. Overall, these results demonstrate the good performance of our approach in different DGP designs.
7.2 Dynamic Ordered Choice Model
In this section, we investigate a dynamic ordered choice model with one lagged dependent variable $Y_{i,t-1}$. The latent dependent variable $Y_{it}^{*}$ is generated as follows:
$Y_{it}^{*}=Z_{it}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}.$
[…]
The observed dependent variable $Y_{it}$ is given as
$Y_{it}=1*(Y_{it}^{*}\leq b_{2})+2*(0\leq Y_{it}^{*}\leq b_{2})+3*(Y_{it}^{*}>b_{3}),$
for $1\leq t\leq T$. The initial value $Y_{i0}\in\{1,2,3\}$ is generated independently of all variables, and follow the distribution $\Pr(Y_{i0}=1)=0.6,\Pr(Y_{i0}=2)=\Pr(Y_{i0}=3)=0.2$.
In this dynamic model, the covariates $Z_{i}:=(Z_{it})_{t=1}^{T}$ and the initial value $Y_{i0}$ are exogenous, while the lagged variable $Y_{i,t-1}$ is endogenous. Proposition 4 characterizes the identified set for $\theta_{0}$ with the following conditional moment inequalities:
  • (1)
    When $s\in\{2,3\}$,
    $1-\sum_{j=2}^{3}\Pr\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}$$\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma\geq a\}$
    $\geq\sum_{j=1}^{2}\Pr\left(Y_{is}=y_{j},b_{j+1}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\leq a\mid z,y_{0}\right),$
    $1-\sum_{j=2}^{3}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq a\mid z,y_{0})$
    $\geq\sum_{j=1}^{2}\Pr$ $\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}\{b_{j+1}-z_{1}^{\prime}\beta-y_{0}\gamma\leq a\},$
    for any $a\in\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma,b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma\}_{j=2}^{T}$;
  • (2)
    When $s,t\in\{2,3\}$,
    $1-\sum_{j=2}^{3}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq a\mid z,y_{0})$
    $\geq\sum_{j=1}^{2}$ $\Pr\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq a\mid z,y_{0}\right),$
    for any $a\in\{b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma,b_{j}-z_{t}^{\prime}\beta-\gamma,b_{j}-z_{t}^{\prime}\beta-2\gamma,b_{j}-z_{t}^{\prime}\beta-3\gamma\}_{j=2}^{3}$.
We normalize the first parameter $\beta_{0}$ to one, and report the performance of the coefficient $\gamma_{0}$ for the lagged dependent variable. Tables 3 and 4 illustrate that our approach yields robust and informative results for the dynamic ordered choice model across various DGP specifications. The coverage probability of the CI nearly reaches 95%, and the CI consistently excludes zero, producing significant coefficients. These results remain similar across different values of correlation coefficients. When the standard deviation $\sigma_{z}$ increases, the length of the CI also experiences a slight increase. This phenomenon occurs because, in the dynamic model, only partial identification is achieved, and the bound for $\gamma_{0}$ depends on the variation in $\Delta z^{\prime}\beta_{0}$. A larger variation in $\Delta z^{\prime}\beta_{0}$ may result in a wider identified set in this specification, but it still provides informative results. As the sample size increases, the confidence interval shrinks, and concurrently, the coverage probability improves in all specifications.
Table 3: Performance of $\gamma_{0}$ under different values of $\sigma_{z}$ ($\rho=0.25$)
$\sigma_{z}$ CI CP length Power $l_{MAD}$ $u_{MAD}$
$N=2000$
$\sigma_{z}=1$ [0.446, 1.606] 0.935 1.160 1.000 0.565 0.625
$\sigma_{z}=1.5$ [0.375, 1.673] 0.959 1.298 1.000 0.629 0.693
$\sigma_{z}=2$ [0.311, 1.730] 0.960 1.418 1.000 0.700 0.739
$N=8000$
$\sigma_{z}=1$ [0.529, 1.495] 0.969 0.966 1.000 0.473 0.504
$\sigma_{z}=1.5$ [0.460, 1.559] 0.965 1.100 1.000 0.548 0.564
$\sigma_{z}=2$ [0.427, 1.585] 0.985 1.158 1.000 0.573 0.589
Table 4: Performance of $\gamma_{0}$ under different values of $\rho$ ($\sigma_{z}=1$)
$\rho$ CI CP length Power $l_{MAD}$ $u_{MAD}$
$N=2000$
$\rho=0$ [0.472, 1.593] 0.932 1.121 1.000 0.550 0.607
$\rho=0.25$ [0.446, 1.606] 0.935 1.160 1.000 0.565 0.625
$\rho=0.5$ [0.457, 1.631] 0.943 1.173 1.000 0.548 0.648
$N=8000$
$\rho=0$ [0.528, 1.472] 0.958 0.945 1.000 0.475 0.487
$\rho=0.25$ [0.529, 1.495] 0.969 0.966 1.000 0.473 0.504
$\rho=0.5$ [0.535, 1.515] 0.975 0.980 1.000 0.467 0.519
8 Empirical Application
In this section, we apply our proposed approach to explore the empirical analysis of income categories using the NLSY79 dataset. The dependent variable is three categories of (log) income, denoted by the three values $\{1,2,3\}$, indicating whether an individual falls within the top 33.3% highest income bracket, the 33.3%-66.6% highest income range, and the lowest 33.3% income tier, respectively. We include two covariates in this analysis: one is tenure, defined as the total duration (in weeks) with the current employer, and the other is the residence indicator for whether one lives in an urban or rural area. [^8] This dataset also contains other crucial factors for income such as gender and race. However, these variables are time-invariant and cannot be included for panel models with fixed effects. We use two periods of panel data from the years 1982 and 1983 as well as the income data from 1981 as initial values, and there are $n=5259$ individuals in each period. The following table presents the summary statistics of these variables.
Table 5: Application: Summary Statistics
income category residence tenure /100
mean 1.990 0.799 0.825
s.d. 0.810 0.401 0.738
25% quantile 1.000 1.000 0.220
median 2.000 1.000 0.605
75% quantile 3.000 1.000 1.280
minimum 1.000 1.000 0.010
maximum 3.000 1.000 4.850
We adopt various ordered response models introduced in Section 3 to analyze the income category. The first model is the standard static model without any endogeneity. The second is the static model, while treating residence as an endogenous covariate. Residence is potentially endogenous since the choice of living area is typically endogenously determined and may be correlated with individuals’ unobserved ability or preference. The last model considers the dynamic model with one lagged dependent variable, allowing people’s income in current periods to depend on their income in the last period. All three models allow for individual fixed effects and do not impose any parametric distributions on time-changing shocks. Proposition 4 characterizes the identified set of the model coefficients for these three models using conditional moment inequalities. Similar to Section 7, we exploit the kernel-based CLR inference method to construct confidence intervals. The coefficient of the variable ‘residence’ is normalized to one. Table 6 reports the confidence intervals for the coefficients of the covariate ‘tenure’ and the lagged dependent variable (when applicable).
Table 6: Application: Income Categories
$\beta_{0,1}$ (residence) $\beta_{0,2}$ (tenure) $\gamma_{0}$ (lag)
exogenous static model 1 [0.612, 0.939] -
endogenous static model 1 [0.041, 0.939] -
dynamic model 1 [0.531, 0.694] [0.286, 0.612]
As shown in Table 6, tenure exhibits a significantly positive effect on the income category across all specifications. When allowing for the endogeneity of residence, the confidence interval for tenure becomes wider, as we need to account for all possible correlations between residence and unobserved heterogeneity. The results from the dynamic model show that the income category in the current period is also positively affected by last period’s income, and this effect is significant. Furthermore, this analysis demonstrates the flexibility of our approach, which can not only allow for endogeneity introduced by dynamics but also account for contemporary endogeneity.
9 Conclusion
We introduce a general method to identify nonlinear panel data models based on a partial stationarity condition. This approach accommodates dynamic models with an arbitrary finite number of lagged outcome variables and other types of endogenous covariates. We demonstrate how our key identification strategy can be applied to obtain informative identifying restrictions in various limited dependent variable models, including binary choice, ordered response, multinomial choice, as well as censored dependent variable models. Finally, we further extend this approach to study general nonseparable models.
There are some natural directions for follow-up research. In this paper we focus on the identification of model parameters, but it would also be interesting to investigate how our identification strategy can be exploited to obtain informative bounds on average marginal effects and other counterfactual parameters, say, following the approach proposed in Botosaru and Muris (2022). [^9] Botosaru and Muris (2022) proposes an approach to obtain bounds on counterfactual CCPs in semiparametric dynamic panel data models, assuming that the index parameters are (partially) identified. Additionally, the idea of bounding an endogenous object (parametric index in our case) by an arbitrary constant so as to obtain an object free of endogeneity issues may have broader applicability beyond the models studied in this work, and it remains to see whether our key identification strategy can be further adapted to other structures.
Appendix A Appendix
A.1 Proof of Lemma 1
Proof.
The identifying condition (7) for the parameter $\theta$ in Proposition 1 is given as
$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st}).$ (13)
Now we show that when $X_{it}\in\{a_{1},...,a_{K}\}$, the above condition is informative only at the $2K$ points $a\in\mathcal{Q}_{1}=\{z_{t}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma,z_{s}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{s}^{\prime}\beta+a_{K}^{\prime}\gamma\}$. Let $q_{k}\in\mathcal{Q}_{1}$ denote an element in $\mathcal{Q}$. Without loss of generalization, we can rank $q_{k}$ as $q_{1}\leq q_{2}\leq...\leq q_{2K}$.
When $a>q_{2K}$, then $a>z_{s}^{\prime}\beta+\max\limits_{k}{a_{k}^{\prime}\gamma}$. Therefore, the upper bound becomes $11-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})=$ 1 and the identifying condition holds for any $\theta$. Therefore, there is no information for $\theta$ when $a>q_{2K}$.
When $a<q_{1}$, similarly the lower bound is zero $\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st})=0$ and the inequality has no identifying power for $\theta$.
When $a\in(q_{k},q_{k+1})$ for some $k$, the identifying restriction at $a$ can be implied by the restriction at $a=q_{k+1}$. It is because the upper bound at $q_{k+1}$ is the same as $a$ since there does not exist any point $z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$ that lies in $(q_{k},q_{k+1})$. The lower bound $\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st})$ is weakly larger at $q_{k+1}$ than $a$ since it is weakly increasing. Then, any parameter that satisfies condition (13) evaluated at $q_{k+1}$ also satisfies the one at $a$, so condition (13) evaluated at $a$ does not contain additional information for $\theta$.
A.2 Proof of Theorem 1
To prove sharpness of the identified set $\Theta_{I,1}$, we seek to show that, for any $\theta\in\Theta_{I,1}$, there exists an underlying DGP characterized by the conditional distribution of $(v_{is},v_{it})\mid W_{ist}$) that produces the same observed choice probability and satisfies Assumption 1. The following construction is for one candidate parameter $\theta\in\Theta_{I,1}$. We set the fixed effect to be zero $\alpha_{i}=0$, so that $v_{it}:=-(\epsilon_{it}+\alpha_{i})=-\epsilon_{it}$.
Let $P_{st}\left(j,k\mid w_{st}\right):=\Pr\left(Y_{is}=j,Y_{it}=k\mid W_{ist}=w_{st}\right)$ denote the observed joint probability of selecting $j$ at period $s$ and selecting $k$ at period $t$. The first requirement for sharpness is that the constructed distribution of $(v_{is},v_{it})\mid W_{ist}$ matches the observed joint choice probability according to the model: given $W_{ist}=w_{st}$,
$P_{st}\left(1,1\mid w_{st}\right)$ $=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$ (14)
$P_{st}\left(1,0\mid w_{st}\right)$ $=\Pr(v_{is}\leq w_{s}^{\prime}\theta\mid w_{st})-\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$
$P_{st}\left(0,1\mid w_{st}\right)$ $=\Pr(v_{it}\leq w_{t}^{\prime}\theta\mid w_{st})-\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$
and $P_{st}\left(0,0\mid w_{st}\right)$ can be matched automatically when the above conditions hold.
Another condition for sharpness requires the marginal distribution of $v_{it}\mid W_{ist}$ to satisfy the following stationarity condition in Assumption 1:
$F_{v_{t}\mid Z_{st}}(a\mid z_{st})=F_{v_{s}\mid Z_{st}}(a\mid z_{st})\quad\forall a\in\mathcal{R}.$ (15)
We can focus on a reduced problem, which is to find a marginal distribution $v_{it}\mid W_{ist}$ that matches the marginal choice probability and satisfies the stationarity assumption. Let $P_{t}(w_{st}):=\Pr(Y_{it}=1\mid W_{ist}=w_{st})$ and $P_{s}(w_{st}):=\Pr(Y_{is}=1\mid W_{ist}=w_{st})$ denote the observed marginal choice probability at $t$ and $s$, given as follows:
$P_{t}(w_{st})=\Pr\left(v_{it}\leq w_{t}^{\prime}\theta\mid W_{ist}=w_{st}\right),$ (16)
$P_{s}(w_{st})=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta\mid W_{ist}=w_{st}\right).$
If we can construct a marginal distribution satisfying conditions (15) and (16), then the joint choice probability in (14) is matched by setting the joint distribution of $(v_{is},v_{it})\mid W_{ist}$ at $(w_{s}^{\prime}\theta,w_{t}^{\prime}\theta)$ as $P_{st}\left(1,1\mid w_{st}\right)=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right)$. This construction is feasible since there is no other assumptions on the joint distribution $(v_{is},v_{it})\mid W_{ist}$. Moreover, other choice probabilities such as $P_{st}\left(1,0\mid w_{st}\right)$ can be computed using the marginal probability $P_{s}(w_{st})$ subtracting $P_{st}\left(1,1\mid w_{st}\right)$.
Now, for any $\theta\in\Theta_{I,1}$, our objective is to construct the distribution of $v_{it}\mid W_{st}$ and $v_{is}\mid W_{st}$ such that conditions (15) and (16) hold. Consider that the endogenous covariate $X_{it}$ only takes $K$ values: $X_{it}\in\mathcal{X}=\{a_{1},...,a_{K}\}$ for any $t\leq T$.  [^10] The covariate $X_{it}$ is allowed to have different support across time $t$, then $\mathcal{X}$ is the union of the support of $X_{it}$ over all periods. Given $Z_{it}=z_{t}$, the covariate index $z_{t}^{\prime}\beta+x_{t}^{\prime}\gamma$ belongs to the $K$ points $\{z_{t}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma\}$. Without loss of generalization, we can rank the $K$ points as $\alpha_{1}^{\prime}\gamma\leq a_{2}^{\prime}\gamma\leq...\leq a_{K}^{\prime}\gamma$ since we can reindex the $K$ values of $\{a_{1},...,a_{K}\}$ such that this ranking is satisfied. Let $m_{k}:=z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma$ denote the $K$ points for which $w_{t}^{\prime}\theta$ can take, and $n_{k}:=z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$ for $1\leq k\leq K$ denote the $K$ points $w_{s}^{\prime}\theta$ can take. Let $\bar{m}=\max\{z_{t}^{\prime}\beta,z_{s}^{\prime}\beta\}+a_{K}^{\prime}\gamma+\epsilon$ for some $\epsilon>0$.
We first look at the stationarity condition in (15). Section 2.2 derives bounds for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, given as follows:
$F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ $\leq\sum_{x_{st}\in\mathcal{X}}\Big{\{}F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta\geq a\}+\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta<a\}\Big{\}}\Pr(X_{st}=x_{st}\mid z_{st})$
$:=U_{t}(a\mid z_{st}),$
$F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ $\geq\sum_{x_{st}\in\mathcal{X}}F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta\leq a\}\Pr(X_{st}=x_{st}\mid z_{st}):=L_{t}(a\mid z_{st}).$
Given the definition of the identified set $\Theta_{I,1}$, for $\theta\in\Theta_{I,1}$, the following condition holds for any $a\in\mathcal{R}$,
$\max_{s,t}\{L_{s}(a\mid z_{st}),L_{t}(a\mid z_{st})\}\leq\min_{s,t}\{U_{t}(a\mid z_{st}),U_{s}(a\mid z_{st})\},$
Let $\tilde{L}(a\mid z_{st})=\max_{s,t}\{L_{t}(a\mid z_{st}),L_{s}(a\mid z_{st})\}$, the stationarity condition (15) is satisfied by constructing the conditional distribution of $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ as follows:
$F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})=\left\{\begin{aligned} &0\ &\text{if }a<m_{1},\\ \tilde{L}&(a\mid z_{st})\ &\text{if }m_{1}\leq a<\bar{m},\\ &1\ &\text{if }\bar{m}\leq a.\end{aligned}\right.$ (17)
The above construction guarantees that the marginal distribution $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})$ lies between the bounds $[L_{s}(a\mid z_{st}),U_{s}(a\mid z_{st})]$. From the definition of $L_{t}(a)$, we know that $L_{t}(a\mid z_{st})\geq L_{t}(a^{\prime}\mid z_{st})$ for any $a\geq a^{\prime}$. Therefore, the monotonicity of the above distribution is satisfied as $\max\limits_{s,t}\{L_{t}(a\mid z_{st}),L_{s}(a\mid z_{st})\}\geq\max\limits_{s,t}\{L_{t}(a^{\prime}\mid z_{st}),L_{s}(a^{\prime}\mid z_{st})\}$ for any $a\geq a^{\prime}$. Additionally, the above construction satisfies the stationarity assumption given $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ for any $a$.
We transform the two requirements of sharpness into finding a distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ to generate the marginal distribution $F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ in (17) and also match the marginal choice probability given in condition (16).
Construction
We focus on the construction for $v_{it}\mid W_{ist}$, and the construction for $v_{is}\mid W_{ist}$ at period $s$ is similar so it is omitted here. The construction proceeds in four steps. We first constructs a distribution to match the lower and upper bounds $L_{t}(a\mid z_{st})$ and $U_{t}(a\mid z_{st})$. Then, we construct the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ at the points $\{m_{1},...,m_{K}\}$ with $m_{k}=z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma$. The third step look at the construction at the points $\{n_{1},...,n_{K}\}$ with $n_{k}=z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$. The final step examines the points $a$ that are not in $\{m_{1},...,m_{K},n_{1},...,n_{K}\}$.
Step 1:
We first show that the bounds $U_{t}(a\mid z_{st})$ and $L_{t}(a\mid z_{st})$ can be achieved under a distribution of $v_{it}\mid W_{ist}$ that matches the marginal choice in (16). From its definition, the upper bound $U_{t}(a\mid z_{st})$ can be achieved under the following distribution $F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st})$:
$F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st})=\left\{\begin{aligned} &0\qquad&\text{if }a<m_{1},\\ &P_{t}(w_{st})\qquad&\text{if }m_{1}\leq a<w_{t}^{\prime}\theta+\epsilon,\\ &1\qquad&\text{if }w_{t}^{\prime}\theta+\epsilon\leq a,\end{aligned}\right.$
where $\epsilon>0$ is used to satisfy the right-continuity property of the distribution.
Similarly, the lower bound function $L_{t}(a\mid z_{st})$ can be achieved under the following distribution $F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})$:
$F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})=\left\{\begin{aligned} &0\qquad&\text{if }a<w_{t}^{\prime}\theta,\\ &P_{t}(w_{st})\qquad&\text{if }w_{t}^{\prime}\theta\leq a<\bar{m},\\ &1&\text{if }\bar{m}\leq a\end{aligned}\right.$
From the definition of the distributions $F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st}),F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})$, they both satisfy condition (16) which requires $F^{U}_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})=F^{L}_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})=P_{t}(w_{st})$.
Step 2: this step establishes the construction of the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ at the points $m_{k}$. The observed marginal distribution given in condition (16) only restricts the distribution of $v_{it}\mid W_{ist}$ at each point $m_{k}$, given $w_{st}=(z_{st},x_{s},a_{k})$,
$P_{t}(w_{st})=\Pr(v_{it}\leq m_{k}\mid z_{st},x_{s},a_{k}).$
(i) We start with the largest value $m_{K}=z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma$. From the definition of $U_{t}(a\mid z_{st}),L_{t}(a\mid z_{st})$, we know that $U_{t}(m_{1}\mid z_{st})=L_{t}(m_{K}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})\leq U_{t}(m_{K}\mid z_{st})$, which is described in the following graph:
$U_{t}(m_{1}\mid z_{st})$$=L_{t}(m_{K}\mid z_{st})$$U_{t}(m_{K-1}\mid z_{st})$$\tilde{L}(m_{K}\mid z_{st})$$U_{t}(m_{K}\mid z_{st})$
Given the $K$ points $U_{t}(m_{1}\mid z_{st})\leq U_{t}(m_{2}\mid z_{st})\leq...\leq U_{t}(m_{K}\mid z_{st})$, we can find $2\leq j_{1}\leq K$ such that $U_{t}(m_{j_{1}-1}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})\leq U_{t}(m_{j_{1}}\mid z_{st})$. Then we can express $\tilde{L}(m_{K}\mid z_{st})$ as follows: for some $\alpha_{1}\in[0,1]$,
$\tilde{L}(m_{K}\mid z_{st})=\alpha_{1}U_{t}(m_{j_{1}-1}\mid z_{st})+(1-\alpha_{1})U_{t}(m_{j_{1}}\mid z_{st}).$
Accordingly, we construct the distribution $F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$ as follows:
$F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})=\alpha_{1}F^{U}_{t}(m_{j_{1}-1}\mid w_{st})+(1-\alpha_{1})F^{U}_{t}(m_{j_{1}}\mid w_{st}).$
The above construction generates the marginal distribution $\tilde{L}(m_{K}\mid z_{st})$. Moreover, it matches the marginal choice probability when $x_{t}=a_{K}$ in condition (16) since $F^{U}_{t}(m_{j}\mid w_{st})=P_{t}(w_{st})$ if $x_{t}=a_{K}$ for any $1\leq j\leq K$ from the definition of $F^{U}_{t}(m_{j}\mid w_{st})$.
(ii) Now consider that $a=m_{K-1}$. Given that $L_{t}(m_{K-1}\mid z_{st})\leq\tilde{L}(m_{K-1}\mid z_{st})\leq U_{t}(m_{K-1}\mid z_{st})$, there either exists $11\leq j_{2}\leq K-$ 1 such that $\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})]$ or $\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})]$, given in the following graph:
$L_{t}(m_{K-1}\mid z_{st})$$U_{t}(m_{1}\mid z_{st})$$=L_{t}(m_{K}\mid z_{st})$$U_{t}(m_{K-1}\mid z_{st})$$\tilde{L}(m_{K-1}\mid z_{st})$$U_{t}(m_{K}\mid z_{st})$
Then $\tilde{L}(m_{K-1}\mid z_{st})$ can be expressed as follows, for $\alpha_{2}\in[0,1]$ and $\tilde{\alpha}_{2}\in[0,1]$,
$\tilde{L}(m_{K-1}\mid z_{st})$
$=$ $&\alpha_{2}U_{t}(m_{j_{2}-1}\mid z_{st})+(1-\alpha_{2})U_{t}(m_{j_{2}}\mid z_{st})\ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})],\\ &\tilde{\alpha}_{2}L_{t}(m_{K-1}\mid z_{st})+(1-\tilde{\alpha}_{2})U_{t}(m_{1}\mid z_{st})\ \ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})].\end{aligned}\right.$
Accordingly, the distribution $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$ at $m_{K-1}$ is constructed as
$F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$
$=$ $&\alpha_{2}F^{U}_{t}(m_{j_{2}-1}\mid w_{st})+(1-\alpha_{2})F^{U}_{t}(m_{j_{2}}\mid w_{st})\ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})],\\ &\tilde{\alpha}_{2}F^{L}_{t}(m_{K-1}\mid z_{st})+(1-\tilde{\alpha}_{2})F^{U}_{t}(m_{1}\mid w_{st})\ \ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})].\end{aligned}\right.$
Condition (16) is satisfied since $U_{t}(m_{j}\mid w_{st})=P_{t}(w_{st})$ when $x_{t}=a_{K-1}$ for any $11\leq j\leq K-$ 1.
Now we still need to show that the constructed distribution satisfies monotonicity. If $\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})]$, then $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{U}_{t}(m_{1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$. Otherwise, by the monotonicity of $\tilde{L}(a\mid z_{st})$ in $a$, we know that $j_{2}\leq j_{1}$. If $j_{2}\leq j_{1}-1$, then monotonicity is satisfied since $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{U}_{t}(m_{j_{2}}\mid w_{st})\leq F^{U}_{t}(m_{j_{1}-1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$. If $j_{1}-1<j_{2}=j_{1}$, then it implies that $\alpha_{2}\leq\alpha_{1}$ since $\tilde{L}(m_{K-1}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})$. Therefore, we have $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$.
The distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ for the remaining points $a\in\{m_{1},...,m_{K-2}\}$ can be constructed in the same way as $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$, so it is omitted here.
Step 3: we show the construction at the points $a=n_{k}$ for $1\leq k\leq K$. For points $a=n_{k}$, we only need to construct a proper distribution to be consistent with Equation (17). We still start with the largest point $n_{K}$. The distribution $\tilde{L}(n_{K}\mid z_{st})$ can be expressed as follows in different scenarios: for some $2\leq j_{3}\leq K$
$\tilde{L}(n_{K}\mid z_{st})=\left\{\begin{aligned} &\alpha_{3}\tilde{L}(m_{K}\mid z_{st})+(1-\alpha_{3})*1&\text{ if }n_{K}>m_{K},\\ &\alpha_{4}\tilde{L}(m_{j_{3}-1}\mid z_{st})+(1-\alpha_{4})\tilde{L}(m_{j_{3}}\mid z_{st})&\text{ if }m_{j_{3}-1}<n_{K}<m_{j_{3}},\\ &\alpha_{5}*0+(1-\alpha_{5})\tilde{L}(m_{1}\mid z_{st})&\text{ if }n_{K}<m_{1},\end{aligned}\right.$
where $\alpha_{3},\alpha_{4},\alpha_{5}\in[0,1]$.
Accordingly, the distribution $F^{*}_{v_{t}\mid W_{st}}(n_{k}\mid w_{st})$ is constructed as follows,
$F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})=\left\{\begin{aligned} &\alpha_{3}F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})+(1-\alpha_{3})*1&\text{ if }n_{K}>m_{K},\\ &\alpha_{4}F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}-1}\mid w_{st})+(1-\alpha_{4})F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}}\mid w_{st})&\text{ if }m_{j_{3}-1}<n_{K}<m_{j_{3}},\\ &\alpha_{5}*0+(1-\alpha_{5})F^{*}_{v_{t}\mid W_{st}}(m_{j_{1}}\mid w_{st})&\text{ if }n_{K}<m_{1}.\end{aligned}\right.$
When $a=n_{K-1}$, the distribution $\tilde{L}(n_{K-1}\mid z_{st})$ can be expressed as follows in different scenarios: for some $11\leq j_{4}\leq j_{3}-$ 1,
$\tilde{L}(n_{K-1}\mid z_{st})=\left\{\begin{aligned} &\alpha_{6}\tilde{L}(m_{K}\mid z_{st})+(1-\alpha_{6})\tilde{L}(n_{K}\mid z_{st})&\text{ if }m_{K}<n_{K-1}<n_{K},\\ &\alpha_{7}\tilde{L}(m_{j_{3}-1}\mid z_{st})+(1-\alpha_{7})\tilde{L}(n_{K}\mid z_{st})&\text{ if }m_{j_{3}-1}<n_{K-1}<n_{K}<m_{j_{3}},\\ &\alpha_{8}\tilde{L}(m_{j_{4}-1}\mid z_{st})+(1-\alpha_{8})\tilde{L}(m_{j_{4}}\mid z_{st})&\text{ if }m_{j_{4}-1}<n_{K-1}<m_{j_{4}}\leq m_{j_{3}-1},\\ &\alpha_{9}*0+(1-\alpha_{9})\tilde{L}(n_{K}\mid z_{st})&\text{ if }n_{K-1}<n_{K}<m_{1},\end{aligned}\right.$
where $\alpha_{6},\alpha_{7},\alpha_{8},\alpha_{9}\in[0,1]$.
The distribution $F^{*}_{v_{t}\mid W_{st}}(n_{K-1}\mid w_{st})$ is constructed accordingly:
$F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})$
$=$ $&\alpha_{6}F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})+(1-\alpha_{6})F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})&\text{ if }m_{K}<n_{K-1}<n_{K},\\ &\alpha_{7}F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}-1}\mid w_{st})+(1-\alpha_{7})F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})&\text{ if }m_{j_{3}-1}<n_{K-1}<n_{K}<m_{j_{3}},\\ &\alpha_{8}F^{*}_{v_{t}\mid W_{st}}(m_{j_{4}-1}\mid w_{st})+(1-\alpha_{8})F^{*}_{v_{t}\mid W_{st}}(m_{j_{4}}\mid w_{st})&\text{ if }m_{j_{4}-1}<n_{K-1}<m_{j_{4}}\leq m_{j_{3}-1},\\ &(1-\alpha_{9})F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})&\text{ if }n_{K-1}<n_{K}<m_{1}.\end{aligned}\right.$
The monotonicity of the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ among the $2K$ points $\{m_{1},...,m_{K},n_{1},...,n_{K}\}$ is satisfied by construction. Furthermore, the construction for the rest of points $\{n_{1},...,n_{K-2}\}$ are similar so it is omitted here.
Step 4: For any $a\notin\{m_{1},...,m_{K},n_{1},...,n_{K}\}$, The distribution is constructed as $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})=0$ if $a<m_{1}$ and $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})=1$ if $a\geq\bar{m}$. When $a\in[m_{1},\bar{m}]$, there must exist some $q_{j}\in\{m_{1},...,m_{K},n_{1},...,n_{K}\}$ such that $q_{j}<a$ and $\tilde{L}(a\mid z_{st})=\tilde{L}(q_{j}\mid z_{st})$. Then we set $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})=F^{*}_{v_{t}\mid W_{st}}(m_{j}\mid w_{st})$. This completes the construction for the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ for any $a$.
A.3 Proof of Proposition 2 and 3
Proof.
The proof for the point identification of $\beta_{0}$ consists of two steps: we first show that when $\Delta z\in\Delta\mathcal{Z}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified from the identifying condition (7) in Proposition 1. Then, the large support condition in Assumption 2 ensures that $\beta_{0}$ is point identified up to scale.
When $X_{it}$ is discrete and there are two periods $T=2$, the identifying condition (7) is given as
$1-\Pr(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{12})\geq\Pr(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12}),$
for $a\in\{z_{1}^{\prime}\beta_{0}+a_{1}^{\prime}\gamma_{0},...,z_{1}^{\prime}\beta_{0}+a_{K}^{\prime}\gamma_{0},z_{2}^{\prime}\beta_{0}+a_{1}^{\prime}\gamma_{0},...,z_{2}^{\prime}\beta_{0}+a_{K}^{\prime}\gamma_{0}\}$, and another identifying condition switches the order of period 1 and 2.
Let $a=z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$, [^11] The value of $a=z_{2}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$ leads to the same identifying condition., then the above upper bound can be further bounded as
$1-\Pr(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}\mid z_{12})\leq 1-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12}).$
When $z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0$ which implies $z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}\geq z_{2}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$, then the lower bound can be bounded below as
$\Pr(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{\mid}z_{12})\leq\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12}).$
Combining the above results leads to
If $\text{If }z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12})\geq\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12}).$
The contraposition of the above inequality yields
$01-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}>$ 0 .
Switching the order of the time period leads to another identifying restriction as follows:
$01-\Pr(Y_{i1}=1,X_{i1}=a_{k}\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=a_{k}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
Therefore, when $\Delta z\in\Delta\mathcal{Z}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified.
Next, we show that $\beta_{0}$ is point identified under the large support assumption. To prove it, we will show that for any $\beta\neq k\beta_{0}$ for some $k$, there exists some value $\Delta z$ such that $\Delta z^{\prime}b$ has different signs from $\Delta z^{\prime}\beta_{0}$.
From Assumption 2, the conditional support of $\Delta z^{j^{*}}$ is $\mathcal{R}$ and $\beta_{0}^{j^{*}}\neq 0$. We focus on the case where $\beta_{0}^{j^{*}}>0$, and the analysis also applies to the other case. Let $\Delta\tilde{z}:=\Delta z\setminus\Delta z^{j^{*}}$ denote the remaining covariates in $\Delta z$ and $\tilde{\beta}_{0}$ denote its coefficient. For any candidate $b$, we discuss three cases: $b^{j^{*}}<0$, $b^{j^{*}}=0$, and $b^{j^{*}}>0$.
Case 1: $b^{j^{*}}<0$. When the covariate $\Delta z^{j^{*}}$ takes a large positive value $\Delta z^{j^{*}}\rightarrow+\infty$ and the remaining covariates take bounded values in their support, it implies that $\Delta z^{\prime}\beta_{0}>0$ and $\Delta z^{\prime}b<0$.
Case 2: $b^{j^{*}}=0$. For any value $\Delta z$, the value of $\Delta z^{\prime}b$ is either positive or nonpositive. When $\Delta z^{\prime}b>0$ is positive, then let $\Delta z^{j*}$ take a large negative value $\Delta z^{j*}\rightarrow-\infty$ such that $\Delta z^{\prime}\beta_{0}<0$, which has a different sign from $\Delta z^{\prime}b$. Similarly, if $\Delta z^{\prime}b\leq 0$, there exists $\Delta z^{j^{*}}\rightarrow+\infty$ such that $\Delta z^{\prime}\beta_{0}>0$.
Case 3: $b^{j^{*}}>0$. Assumption 2 requires that $\Delta\mathcal{Z}$ is not contained in any proper linear subspace, so there exists $\Delta z$ such that $\Delta\tilde{z}^{\prime}\tilde{\beta}_{0}/\beta_{0}^{j^{*}}\neq\Delta\tilde{z}^{\prime}\tilde{b}/b^{j^{*}}$. Suppose that $\Delta\tilde{z}^{\prime}\tilde{\beta}_{0}/\beta_{0}^{j^{*}}-\Delta\tilde{z}^{\prime}\tilde{b}/b^{j^{*}}=k>0$, then when the covariate takes the value $\Delta Z_{i}=-\Delta\tilde{z}^{\prime}\tilde{b}/b^{j^{*}}-\epsilon$ with $0<\epsilon<k$. The sign of the covariate index satisfies: $\Delta z^{\prime}\beta_{0}=\beta_{0}^{j^{*}}(k-\epsilon)>0$ and $\Delta z^{\prime}b=-b^{j^{*}}\epsilon<0$. The construction is similar when $k<0$.
For the identification of $\gamma_{0}$, under the similar analysis for $\beta_{0}$, we have
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{3}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0},$
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{4}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}.$
As previously shown, when $(z_{1},z_{2})\in\mathcal{Z}_{2}$, it implies that $\Delta z^{\prime}\beta_{0}<0$. Therefore, when $(z_{1},z_{2})\in\mathcal{Z}_{2}\cap\mathcal{Z}^{j}_{3}$, we have $(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0}<0$ and the sign of $\gamma^{j}_{0}$ is identified given $x_{1}^{j}\neq x_{2}^{j}$. Similarly, when $(z_{1},z_{2})\in\mathcal{Z}_{1}\cap\mathcal{Z}^{j}_{4}$, the sign of $\gamma^{j}_{0}$ is also identified given $(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}>0$. Proposition 3 requires that for any $j\leq d_{x}$, either $\mathcal{Z}_{2}\cap\mathcal{Z}^{j}_{3}\neq\emptyset$ or $\mathcal{Z}_{1}\cap\mathcal{Z}^{j}_{4}\neq\emptyset$ so that the sign of $\gamma_{0}^{j}$ is identified for any $j$.
A.4 Relationship to Khan, Ponomareva, and Tamer (2023)
Lemma 3.
Under Assumption 1 and $X_{it}=Y_{i,t-1}$, condition (7) implies the identifying restriction Khan, Ponomareva, and Tamer (2023), presented as follows:
$(i)$ $P\left(y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+|\gamma_{0}|>0,$
$(ii)$ $P\left(y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0,$
$(iii)$ $P\left(y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}>0,$
$(iv)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}>0,$
$(v)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}>0,$
$(vi)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\gamma_{0}>0,$
$(vii)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0,$
$(viii)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\gamma_{0}>0,$
$(ix)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}>0.$
Proof.
When the endogenous regressor is the lagged dependent variable $X_{it}=Y_{it-1}$, the identifying condition in Proposition 1 is given as
$U_{s}(a):=1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta_{0}+Y_{i,s-1}\gamma_{0}\geq a\mid z_{st})\\ \geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq a\mid z_{st}):=L_{t}(a).$
Since $Y_{i,t-1}$ only takes two values $Y_{i,t-1}\in\{0,1\}$, we can express the probability in the above bounds as a mixture of $Y_{i,t-1}=1$ and $Y_{i,t-1}=0$:
$U_{s}(a)=1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\mathbf{\mathbbm{1}}\{z_{s}^{\prime}\beta_{0}$ $+\gamma_{0}\geq a\}$
$-\Pr(Y_{is}=0,Y_{i,s-1}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{s}^{\prime}\beta_{0}\geq a\}.$
Similarly, the lower bound is given as
$L_{t}(a)=\Pr(Y_{it}=1,Y_{i,t-1}=1\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}$ $+\gamma_{0}\leq a\}$
$+\Pr(Y_{it}=1,Y_{i,t-1}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}\leq a\}.$
Now we discuss different scenarios for the value of $(z_{t}^{\prime}\beta_{0},z_{s}^{\prime}\beta_{0},\gamma_{0})$, leading to various identifying restrictions in Khan, Ponomareva, and Tamer (2023) described in Lemma 3.
Case 1: consider that $z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}\leq z_{s}^{\prime}\beta_{0}+\min\{0,\gamma_{0}\}$. Let $a=z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}$, the bounds can be simplified as
$U_{s}(a)$ $=1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})-\Pr(Y_{is}=0,Y_{i,s-1}=0\mid z_{st})=\Pr(Y_{is}=1\mid z_{st}),$
$L_{t}(a)$ $=\Pr(Y_{it}=1,Y_{i,t-1}=1\mid z_{st})+\Pr(Y_{it}=1,Y_{i,t-1}=0\mid z_{st})=\Pr(Y_{it}=1\mid z_{st}).$
By exploiting the condition $\max\{0,\gamma_{0}\}-\min\{0,\gamma_{0}\}=|\gamma_{0}|$, the identifying condition is given as
$(z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0})+|\gamma_{0}|\leq 0\Longrightarrow U_{s}(a)=\Pr(Y_{is}=1\mid z_{st})\geq L_{t}(a)=\Pr(Y_{it}=1\mid z_{st}).$
The contraposition of the above condition yields condition (i) in Lemma 3:
$\Pr(Y_{it}=1\mid z_{st})>\Pr(Y_{is}=1\mid z_{st})\Longrightarrow(z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0})+|\gamma_{0}|>0.$
Case 2: consider that $z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}\leq z_{s}^{\prime}\beta_{0}+\gamma_{0}$. Let $a=z_{s}^{\prime}\beta_{0}+\gamma_{0}$, the bounds are simplified as
$U_{s}(a)\leq 1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st}),\quad L_{t}(a)=\Pr(Y_{it}=1\mid z_{st}).$
Given that $\max\{0,\gamma_{0}\}-\gamma_{0}=-\min\{0,\gamma_{0}\}$, we have the following identifying condition:
$z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}\leq 0\Longrightarrow 1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\geq\Pr(Y_{it}=1\mid z_{st}).$
The contraposition of this condition leads to condition (ii) in Lemma 3:
$\Pr(Y_{it}=1\mid z_{st})>1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\Longrightarrow z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0.$
The rest of conditions in Lemma 3 can be derived in a similar way so we omit the analysis here.
A.5 Proof of Proposition 2
Proof.
The identifying restriction (9) in Proposition 4 is given as follows:
$U_{t,order}(a\mid z_{st})$ $:=1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})$ (18)
$\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st}):=L_{t,order}(a\mid z_{st}).$
The proof idea of Corollary 2 is the same as Appendix 1 for Corollary 1. Let $q_{k}\in\mathcal{Q}_{2}$ denote a point in $\mathcal{Q}_{2}$, and we can rank the $2K(J-1)$ points as $q_{1}\leq q_{2}...\leq q_{2K(J-1)}$.
When $a>q_{2K(J-1)}$, then $\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})=0$ for any $j$ such that the upper bound is one $U_{t,order}(a\mid z_{st})=1$. Therefore, condition (18) holds for any $\theta$ and it is not informative for $\theta$.
When $a<q_{1}$, then $\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st})=0$ for any $2\leq j\leq K$ so that the lower bound is zero $L_{t,order}(a\mid z_{st})=0$ and it is not informative for $\theta$.
When $a\in(q_{k},q_{k+1})$, we show that the identifying condition (18) at $q_{k+1}$ contains more information than the one at $a$. The lower bound $L_{t,order}(a\mid z_{st})$ is increasing in $a$, therefore the lower bound evaluated at $q_{k+1}$ is larger than the one evaluated at $a$. The upper bound is the same when evaluated at $q_{k+1}$ and at $a$, as there does not exist any point $b_{j}-z_{s}^{\prime}\beta-a_{k}^{\prime}\gamma$ that lies in $[a,q_{k+1})$ from the definition of $q_{k}$. Therefore, if the parameter $\theta$ satisfies condition (18) evaluated at $q_{k+1}$, it must satisfy condition (18) evaluated at $a$. The identifying condition at $a$ does not provide additional information for $\theta$.
A.6 Proof of Lemma 1
Proof.
When all regressors are exogenous in the sense $\epsilon_{is}\mid(W_{ist},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{ist},\alpha_{i})$, the identifying restriction in Proposition 4 is given as
$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{s}^{\prime}\theta_{0}\geq a\}\\ \geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$ (19)
For any $a\in\mathcal{R}$, the above identifying restriction is only informative when there exists $2\leq j_{1}\leq J$ and $11\leq j_{2}\leq J-$ 1 such that
$b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq a,\quad b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\leq a.$ (20)
It is because if the above condition does not hold, then either the upper bound in condition (19) becomes one or the lower bound becomes zero. Then condition (19) holds for any $\theta$ and there is no identifying power.
Condition (20) can be satisfied for some $a$ if and only if there exists $2\leq j_{1}\leq J$ and $11\leq j_{2}\leq J-$ 1 such that $b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}$. In this case, the identifying condition (19) becomes
if $\text{if }b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\Longrightarrow\sum_{j=1}^{j_{1}-1}\Pr(Y_{is}=y_{j}\mid w_{st})=1-\sum_{j=j_{1}}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\\ \geq\sum_{j=1}^{j_{2}}\Pr(Y_{it}=y_{j}\mid w_{st}).$
Replacing $j_{1}$ with $\tilde{j}_{1}=j_{1}-1$ yielding the results in Proposition 4.
A.7 Proof of Proposition 5 and Proposition 6
Proof.
We first prove that $\beta_{0}$ is point identified up to scale by showing that for any $\Delta z\in\Delta\mathcal{Z}_{order}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified. The identifying condition for $\theta_{0}$ in Proposition 4 is given as
$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{12})\geq\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12}),$
and the same condition also holds when changing the order of period 1 and 2.
Let $a=b_{k}-z_{1}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}$ for $2\leq k\leq J$, then we know that $b_{j}-z_{1}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}\geq a$ if $j\geq k$. Therefore, the upper bound becomes
$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{st})\geq 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x).$
Moreover, if $-z_{2}^{\prime}\beta_{0}\leq-z_{1}^{\prime}\beta_{0}$, then $b_{j+1}-z_{2}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}\leq a$ for $j+1\leq k$. The lower bound can be bounded below by
$\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
Combining the results yields the following conditions:
$\Delta z^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x)\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
The contraposition of the above condition generates
$01-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x)\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
Therefore, when $z\in\mathcal{Z}_{1,order}$, the sign of $\Delta z^{\prime}\beta_{0}<0$ is identified. Similarly, the sign of $\Delta z^{\prime}\beta_{0}>0$ is also identified when $z\in\mathcal{Z}_{2,order}$. Given that the sign of $\Delta z^{\prime}\beta_{0}$ is identified, the parameter $\beta_{0}$ is point identified up to scale under the large support condition $\Delta z$. The analysis is the same as proof A.3 for Proposition 2, so it is omitted here.
Now we look at the result for $\gamma_{0}$ in Proposition 6. Let $a=b_{k}-z_{1}^{\prime}\beta_{0}-x_{1}^{\prime}\gamma_{0}$, then the upper bound becomes
$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{st})\geq 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}).$
If $b_{k}-z_{2}^{\prime}\beta_{0}-x_{2}^{\prime}\gamma_{0}\leq a$, the lower bound becomes
$\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12}).$
Combining the results leads to
if $\text{if }-z_{2}^{\prime}\beta_{0}-x_{2}^{\prime}\gamma_{0}\leq-z_{1}^{\prime}\beta_{0}-x_{1}^{\prime}\gamma_{0}\Longrightarrow\\ 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
Similarly, the contraposition of the above condition gives
$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<(x_{1}-x_{2})^{\prime}\gamma_{0}.$
Therefore, when $\Delta z\in\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}$, it implies that
$0<\Delta z^{\prime}\beta_{0}<(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}.$
Given that $x_{1}^{j}-x_{2}^{j}\neq 0$, it implies that the sign of $\gamma^{j}_{0}$ is identified. Similarly when $\Delta z\in\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}$, we have $0<\Delta z^{\prime}\beta_{0}<(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}$. The sign of $\gamma^{j}_{0}$ is also identified. For any $j\leq d_{x}$, Proposition 6 requires that either $\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}\neq\emptyset$ or $\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}\neq\emptyset$, so the sign of $\gamma^{j}_{0}$ is identified for any $j$.
A.8 Proof of Proposition 7
Proof.
For any set $K\subset\mathcal{J}$, the conditional probability of selecting a choice $j\in K$ given $W_{ist}=w_{st}$ is:
$\Pr(Y_{it}\in K\mid w_{st})=\Pr\left(\exists j\in K\text{ s.t. }w_{ijt}^{\prime}\theta_{0}+v_{ijt}\geq w_{ikt}^{\prime}\theta_{0}+v_{ikt}\ \forall k\in K^{c}\mid w_{st}\right).$
The above observed probability provides information regarding the distribution of $v_{ikt}-v_{ijt}$. Therefore, we can use observed data to bound the following distribution: for $a_{jk}\in\mathcal{R}$,
$Q_{t}(a_{jk}\mid w_{st}):=\Pr\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq a_{jk}\ \forall k\in K^{c}\mid w_{st}\right).$
When $a_{jk}$ satisfies $a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ for any $(j,k)$, then the above probability can be bounded below as
$Q_{t}(a_{jk}\mid w_{st})$ $\geq\Pr\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall k\in K^{c}\mid w_{st}\right)$
$=\Pr(Y_{it}\in K\mid w_{st}).$
Therefore, the lower bound for $Q_{t}(a_{jk}\mid w_{st})$ is established as
$Q_{t}(a_{jk}\mid w_{st})\geq\Pr(Y_{it}\in K\mid w_{st})\mathbf{\mathbbm{1}}\{a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}.$
The above inequality holds since either $a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the lower bound is zero.
Since Assumption 1 is only conditioning on the exogenous covariate $z_{st}$, we can bound the conditional distribution $Q_{t}(a_{jk}\mid z_{st})$ by taking expectation of $X_{it}$ given $z_{st}$:
$Q_{t}(a_{jk}\mid z_{st})\geq\Pr\left(Y_{it}\in K,a_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
Similarly, an upper bound for the conditional probability $Q_{t}(a_{jk}\mid w_{st})$ is derived as follows:
$Q_{t}(a_{jk}\mid w_{st})\leq\Pr(Y_{it}\in K\mid w_{st})\mathbf{\mathbbm{1}}\{a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}+\\ 1-\mathbf{\mathbbm{1}}\{a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}.$
The above inequality holds since either $a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ and the probability $Q_{t}(a_{jk}\mid w_{st})$ is weakly increasing or the upper bound is one.
After taking expectation of $X_{ist}$ given $z_{st}$, the upper bound for $Q_{t}(a_{jk}\mid z_{st})$ is obtained as
$Q_{t}(a_{jk}\mid z_{st})\leq\Pr\left(Y_{it}\in K,a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right)\\ +1-\Pr\left(a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
Rearranging the above formula yields
$Q_{t}(a_{jk}\mid z_{st})\leq 1-\Pr\left(Y_{it}\in K^{c},a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
Under Assumption 1, the conditional probability $Q_{t}(a_{jk}\mid z_{st})$ is the same over the two periods $(s,t)$. Therefore, the upper bound at period $s$ of $Q_{t}(a_{jk}\mid z_{st})$ should be larger than the lower bound at period $t$, yielding the identifying condition (10) as follows:
$1-\Pr(Y_{is}\in K^{c},(z_{js}-z_{ks})^{\prime}\beta_{0}+(X_{ijs}-X_{iks})^{\prime}\gamma_{0}\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st})\\ \geq\Pr(Y_{it}\in K,(z_{jt}-z_{kt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st}).$
A.9 Proof of Proposition 8
Proof.
Since the observed outcome $Y_{it}$ is censored, we either observe $Y_{it}=y>0$ or $Y_{it}=0$. Let $v_{it}=-(\alpha_{i}+\epsilon_{it})$, the conditional probability of $Y_{it}=0$ is given as,
$\Pr(Y_{it}=0\mid w_{st})=\Pr(Y^{*}_{it}\leq 0\mid w_{st})=\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w_{st}).$
When $y>0$, the conditional distribution is given as
$\Pr(Y_{it}\leq y\mid w_{st})$ $=\Pr(Y^{*}_{it}\leq 0,Y_{it}\leq y\mid w_{st})+\Pr(0<Y^{*}_{it},Y_{it}\leq y\mid w_{st})$
$=\Pr(Y^{*}_{it}\leq 0\mid w_{st})+\Pr(0<Y^{*}_{it}\leq y\mid w_{st})$
$=\Pr(Y^{*}_{it}\leq y\mid w_{st})$
$=\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-y\mid w_{st}).$
Combining the two scenarios, the full conditional distributional of $Y_{it}\mid W_{ist}$ is characterized as follows:
$\Pr(Y_{it}\leq y\mid w_{st})=\left\{\begin{aligned} \Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+&x_{t}^{\prime}\gamma_{0}-y\mid w_{st})\quad&\text{ if }y\geq 0,\\ &0\quad&\text{ if }y<0.\end{aligned}\right.$
Given observed distribution of $Y_{it}\mid W_{ist}$, the identification approach is to bound the distribution $\Pr(v_{it}\leq a\mid z_{st})$. We first look at the upper bound for the conditional distribution $v_{it}\mid w_{st}$ given all covariates $W_{ist}=w_{st}$.
$\Pr(v_{it}\geq a\mid w_{st})\leq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\geq a\}+\\ \Pr(Y_{it}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a\},$
where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ such that $\Pr(Y_{it}\leq y\mid w_{st})=\Pr(v_{it}\geq a\mid w_{st})$, or $\Pr(v_{it}\geq a\mid w_{st})\leq\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w_{st})=\Pr(Y_{it}=0\mid w_{st})$ when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a$.
Taking expectation over the endogenous covariate $X_{ist}$ yields the upper bound for the distribution $v_{it}\mid z_{st}$:
$\Pr(v_{it}\geq a\mid z_{st})\leq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\\ \Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st}).$
The above upper bound can be also also expressed as
$\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st})$
$=$ $\Pr(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\geq a\mid z_{st})+\Pr(Y_{it}=0\mid z_{st})$
$=$ $\Pr(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{it}=0\mid z_{st}).$
Similarly, the conditional distribution $v_{it}\mid w_{st}$ can be bounded below
$\Pr(v_{it}\geq a\mid w_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a),$
where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ such that $\Pr(Y_{it}\leq y\mid w_{st})=\Pr(v_{it}\geq a\mid w_{st})$, or the lower bound is zero when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a$.
Taking expectation over $X_{ist}$ leads to the following lower bound:
$\Pr(v_{it}\geq a\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a\mid z_{st}).$
The conditional stationarity condition requires $v_{is}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid Z_{ist}$, which implies that the bounds for $v_{it}\mid Z_{ist}$ must have intersections over any pair of periods $(s,t)$. This restriction generates the following identifying condition for $\theta_{0}$:
$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a\mid z_{st})$
for any $a\in\mathcal{R}$, any $s,t$, and any $z_{st}$.
A.10 Dynamic Censored Models with Latent Dependent Variables
Consider the following dynamic censored models with the latent lagged outcome $Y^{*}_{i,t-1}$:
$Y^{*}_{it}$ $=Z_{it}^{\prime}\beta_{0}+Y^{*}_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
$Y_{it}$ $=\max\{Y^{*}_{it},0\},$
In this model, the endogenous variable $X_{it}$ is the lagged outcome: $X_{it}=Y^{*}_{i,t-1}$. However, the variable $Y^{*}_{i,t-1}$ is not observed in data, so the results in Proposition 8 cannot be directly applied here. Due to this feature in the dynamic model, we need to adjust the results in Proposition 8. Given that $Y^{*}_{i,t-1}=Y_{i,t-1}$ when $Y_{i,t-1}>0$, we can further relax the lower and upper bounds in (8) to identify $\theta_{0}$.
The lower bound in condition (8) can be bounded below as follows:
$\Pr(Y_{it}\leq z_{t}^{\prime}\beta+Y^{*}_{i,t-1}\gamma-a\mid z_{st})\\ \geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}\gamma-a,Y_{i,t-1}>0\mid z_{st}):=L_{t,cen}(a\mid z_{st};\theta).$
Similarly, the upper bound in condition (8) can be further bounded above
$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+Y^{*}_{i,s-1}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\leq U_{s,cen}(a\mid z_{st};\theta),$
where $U_{s,cen}(a\mid z_{st};\theta)$ is defined as
$U_{s,cen}(a\mid z_{st};\theta):=\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}\gamma-a,Y_{i,s-1}>0\mid z_{st})\\ +\Pr(Y_{is}>0,Y_{i,s-1}=0\mid z_{st})+\Pr(Y_{is}=0\mid z_{st}).$
For the dynamic model, an identified set for $\theta_{0}$ is characterized by the following lemma:
Lemma 4.
Under Assumption 1 and $X_{it}=Y_{i,t-1}^{*}$, the identified set for $\theta_{0}$ is characterized by the set of parameters $\theta$ that satisfy the following conditions:
$U_{s,cen}(a\mid z_{st};\theta)\geq L_{t,cen}(a\mid z_{st};\theta).$
for any $a\in\mathcal{R}$, $s,t\leq T$, and $z_{st}$.

### v3

@@
-Identification of Dynamic Nonlinear Panel Models
+Identification of Nonlinear Dynamic Panels
 Abstract
-This paper studies identification for a wide range of nonlinear panel data models, including binary choice, ordered repsonse, and other types of limited dependent variable models.
+This paper provides a general identification approach for a wide range of nonlinear panel data models, including binary choice, ordered response, and other types of limited dependent variable models.
 Our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of (potentially contemporary) endogeneity.
@@
 We demonstrate the robust finite-sample performance of our approach using Monte Carlo simulations, and apply the approach to analyze the empirical application of income categories using various ordered choice models.
-Keywords: Panel Discrete Choice Models; Stationarity; Dynamic Models; Partial Identification; Endogeneity
+Keywords: Dynamic Panel, Discrete Choice, Endogeneity, Stationarity, Partial Identification
 1 Introduction
-This paper studies semiparametric partial identification of a wide range of panel data models with limited dependent variables, including various discrete choice models, censored dependent variable models, and nonseparable models.
-In particular, our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of endogenous covariates.
+Panel models have been a crucial tool in economics, as they not only allow for individual heterogeneity but also for investigating dynamic effects over time.
+However, identifying dynamic panel models can be challenging, especially for models with limited dependent variables.
+Firstly, we may encounter nonseparable and unknown functional forms, rendering it unable to remove individual fixed effects through simple demeaning.
+Furthermore, lagged dependent variables in dynamic models introduce additional endogeneity, posing another challenge for identification.
+This paper provides a general and unified identification approach for a wide range of panel data models with limited dependent variables, including various discrete (binary, multinomial, and ordered) choice models and censored dependent variable models.
+In particular, our approach accommodates dynamic models with any number of lagged dependent variables as well as contemporarily endogenous covariates.
+Moreover, the identification approach does not impose parametric distributions on unobserved heterogeneity, and thus it is robust to distributional specifications.
 To fix ideas, we start with the following dynamic binary choice model, which is on its own of considerable theoretical and applied interest.
-Sections 3-6 generalize the approach to other limited dependent variable models.
+Section 3 generalizes the approach to other limited dependent variable models.
 Specifically, consider
-$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (1)
+$Y_{it}=\mathbf{\mathbbm{1}}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (1)
 where $Y_{it}\in\{0,1\}$ denotes a binary outcome variable for individual $i=1,2,...$ and time $t=1,...,T$, while $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$.
-The objective is to identify the parameter $\theta_{0}:=(\beta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime}$ using a panel of observed variables $(Z_{it},X_{it},Y_{it})_{it}$.
+The objective is to identify the parameter $\theta_{0}:=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$ using a panel of observed variables $\left(Z_{i},X_{i},Y_{i}\right){}_{i=1}^{n}$, where $Z_{i}:=\left(Z_{i1},...,Z_{iT}\right)$, and simiarly $X_{i},Y_{i}$ are vector representations of $X_{it},Y_{it}$.
 We focus on short panels, where the number of time periods $T\geq 2$ is fixed and finite.
@@
 Subsequently, Manski (1987) relaxes the distributional assumption and employs the following conditional stationarity of $\epsilon_{it}$ to achieve identification:
-$\epsilon_{is}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\ \stackrel{{\scriptstyle d}}{{\sim}}\ \epsilon_{it}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (2)
-This condition is also referred to as “group stationarity” or “group homogeneity” and has also been exploited in studies such as Chernozhukov et al.
-(2013), Shi, Shum, and Song (2018) and Pakes and Porter (2022).
-Condition (2) does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows dependence between the fixed effect $\alpha_{i}$ and the covariates $(Z_{it},X_{it})$.
-However, condition (2) does impose substantial restriction on the dependence between $(Z_{it},X_{it})$ and the time-varying error term $\epsilon_{it}$: it effectively requires that all covariates in $(Z_{it},X_{it})$ are exogeneous with respect to the time varying error $\epsilon_{it}$.222 For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=X_{it}^{\prime}\eta$, then the conditional distributions of $\epsilon_{it}$ and $\epsilon_{is}$ cannot be the same as long as $X_{it}^{\prime}\eta\neq X_{is}^{\prime}\eta$, so condition (2) fails in general.
-In many economic applications, certain components of the observable covariates, namely $X_{it}$, may exhibit endogeneity.
-For example, in a dynamic setting where $X_{it}$ includes the lagged outcome variable $Y_{i,t-1}$, then endogeneity of $Y_{i,t-1}$ with respect to $\epsilon_{i,t-1}$ arises immediately.
-For another example, if $X_{it}$ includes “price” or other variables that may be endogenously chosen by economic agents, then the exogeneity restriction imposed by condition (2) will again fail to hold.
-We propose in this paper a general identification approach for various nonlinear panel models in the presence of endogenous regressors, using the following “partial stationarity” condition, which can be viewed as a weaker verision of the stationarity condition (2) above:
-$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\ \stackrel{{\scriptstyle d}}{{\sim}}\ \epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (3)
-Our partial stationary condition (3), as its name suggests, only requires that the errors are stationary across time periods conditional on the realizations of a subvector of the covariates (i.e., the exogenous covariates denoted by $Z$) while allowing the remaining covariates (denoted by $X$) to be endogenous in arbitrary manners.
+$\epsilon_{is}\,\sim\,\epsilon_{it}\mid\alpha_{i},Z_{i},X_{i}\quad\forall s,t=1,...,T$ (2)
+This condition is also referred to as “group stationarity” or “group homogeneity” and has also been exploited in studies such as Chernozhukov, Lee, and Rosen (2013), Shi, Shum, and Song (2018) and Pakes and Porter (2022).
+[^1] To be precise, condition (2) is often stated in the following weaker “pairwise” version in the literature, $\epsilon_{is}\,\sim\,\epsilon_{it}\mid\alpha_{i},Z_{is},Z_{it},X_{is},X_{it},\ \quad\forall s,t=1,...,T,$ where only covariate realizations from the two periods $\left(s,t\right)$ are conditioned on.
+However, the difference between condition (2) and the pairwise version above usually only leads to minor adaption of the results in the aforementioned papers (as well as in the current one).
+See Remark 4 for a follow-up discussion.
+Condition (2) does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows dependence between the fixed effect $\alpha_{i}$ and the covariates $\left(Z_{i},X_{i}\right)$.
+However, condition (2) does impose substantial restriction on the dependence between $\left(Z_{i},X_{i}\right)$ and the time-varying error term $\epsilon_{it}$: it effectively requires that all covariates in $\left(Z_{i},X_{i}\right)$ are exogeneous with respect to the time varying error $\epsilon_{it}$.222 For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=X_{it}^{\prime}\eta$, then the conditional distributions of $\epsilon_{it}$ and $\epsilon_{is}$ cannot be the same as long as $X_{it}^{\prime}\eta\neq X_{is}^{\prime}\eta$, so condition (2) fails in general.
+In many economic applications, certain components of the observable covariates, namely $X_{i}$, may exhibit endogeneity.
+For example, in a dynamic setting where $X_{it}$ includes the lagged outcome variable $Y_{i,t-1}$, then endogeneity of $Y_{i,t-1}$ with respect to $\epsilon_{i,t-1}$ (and all $\epsilon_{i,s}$ with $s\geq t$) arises immediately.
+For another example, if $X_{it}$ includes “price” or other variables that may be endogenously chosen by economic agents after observing $\epsilon_{it}$, then $X_{it}$ would be correlated with contemporary $\epsilon_{it}$, so the exogeneity restriction imposed by condition (2) will again fail to hold.
+In this paper, we instead impose and exploit the following “partial stationarity” condition, which can be viewed as a weaker verision of the condition (2) above:
+$\epsilon_{is}\sim\epsilon_{it}\mid\alpha_{i},Z_{i},\quad\forall s,t=1,...,T.$ (3)
+Our partial stationary condition (3), as its name suggests, only requires that the errors are stationary conditional on the realizations of a subvector of the covariates (i.e., the exogenous covariates denoted by $Z_{i}$) while allowing the remaining covariates (denoted by $X_{i}$) to be endogenous in arbitrary manners.
+[^3] Our identification strategy and results can be easily adapted under the alternative “pairwise parital stationarity” condition $\epsilon_{is}\sim\epsilon_{it}\mid\alpha_{i},Z_{is},Z_{it}.$ See Remarks 4 and 5 for follow-up discussions.
 In short, condition (3) imposes exogeneity conditions only on exogenous covariates.
@@
 [^4] Condition (3) also accommodates the standard stationarity assumption conditional on all covariates.
-We describe how to exploit the partial stationarity condition (3) to derive the identified set on the model parameters $\theta_{0}$ through conditional moment inequalities, which take the form of lower and upper bounds for the conditional distribution $\epsilon_{it}+\alpha_{i}\mid Z_{is},Z_{it}$, solely as functions of observed variables and the model parameters $\theta_{0}$.
-We show that these bounds must have nonzero intersections over time under the partial stationarity assumption, thereby forming identifying restrictions for the parameter $\theta_{0}$.
-While extensive work exists on nonlinear panel models under stationarity, previous studies typically focus on individual models.
+We describe how to exploit the partial stationarity condition (3) to derive the identified set on the model parameters $\theta_{0}$ through a class of conditional moment inequalities, which take the form of lower and upper bounds for the conditional distribution $\epsilon_{it}+\alpha_{i}\mid Z_{i}$, solely as functions of observed variables and the model parameters $\theta_{0}$.
+We show that these bounds must have nonzero intersections over time under the partial stationarity assumption, thereby forming a class of identifying restrictions for the parameter $\theta_{0}$.
+Conditional on the exogenous covariates $Z_{i}$, our class of inequalities is indexed by a scalar $c\in\mathcal{R}$, which implicitly traces out all possible values that the parametrix index $Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}$ can take.
+That said, we show how the effective number of identifying restrictions can be reduced to be finite when $X_{it}$ has finite support, a condition naturally satisfied in the important special case of “$p$-th order autorgressive” dynamic binary choice models, where $X_{it}$ consists of lagged outcome variables $Y_{i,t-1},Y_{i,t-2},...,Y_{i,t-p}$ that are by construction discrete.
+We demonstrate the sharpness of the identified set we derived for binary choice models under the discreteness of the endogenous covariates $X_{it}$.
+More precisely, we show that, for any $\theta$ that satisfies all the conditional moment inequalities we derived, we can construct an observationally equivalent joint distribution of the observed and unobserved variables in our model.
+Our proof of sharpness consists of two steps: first we demonstrate “per-period” sharpness, and then generalize the result to “all-period” sharpness.
+A key innovation in our proof technique is using an explicit, simple, and general construction that shows how marginal/aggregate stationarity restrictions and joint choice probability restrictions can be satisfied simultaneously, which might be of independent and wider use.
+While our main result is about set identification, we also provide sufficient conditions for the point identification of the coefficients on exogenous covariates (under scale normalization) as well as the signs of the coefficients on endogenous covariates.
+Our identification strategy based on partial stationarity applies more broadly beyond the context of dynamic binary choice models.
+In Section 3, we demonstrate its applicability in a general nonseparable semiparametric model under monotonicity, and shows how it can be applied to a range of alternative limited dependent variable models, such as ordered response models, multinomial choice models, and censored outcome models.
+The results of our approach accommodates both static and dynamic settings across all these models.
+While extensive work exists on panel discrete choice or other nonlinear panel data models under stationarity-type conditions, previous studies typically focus on individual models.
 The identification strategies are often context specific and the identification results may have various complicated representations, as seen in studies such as Khan, Ponomareva, and Tamer (2023) and Pakes and Porter (2022).
-Our approach offers the advantage of providing a simple and unified characterization of the identified set for a broad range of static and dynamic panel models, irrespective of the specific types of variables (discrete/continuous outcome and covariates) and forms of endogeneity (lagged/contemporary endogenous regressors).
-Furthermore, our strategy does not rely on scalar-additive structures that are often imposed in various parametric and semiparametric models, and can be further extended to accommodate nonseparable panel data models.
-We demonstrate the sharpness of the identified set for binary choice models when the support of $X_{it}$ is finite.
-More precisely, we show that, for any $\theta$ that satisfies all the conditional moment inequalities we derived, we can construct an observationally equivalent joint distribution of the observed and unobserved variables in our model.
-While our main result is about set identification, we also provide sufficient conditions for the point identification of the coefficients $\beta_{0}$ on exogenous covariates (under scale normalization) as well as the signs of the coefficients on endogenous covariates $\gamma_{0}$.
-Our identification strategy based on partial stationarity can be applied more broadly beyond the context of dynamic binary choice models.
-We demonstrate its applicability to other limited dependent variable models, such as ordered response models, multinomial choice models, and censored outcome models.
-The results of our approach accommodates both static and dynamic settings across all these models.
-Furthermore, we illustrate the adaptability of our key strategy to nonseparable semiparametric models with endogeneity.
+Our approach offers the advantage of providing a simple and unified characterization of the identified set for a broad range of static and dynamic panel models, irrespective of the specific types of variables (discrete/continuous outcome and covariates) and the specific forms of endogeneity (lagged/contemporary endogenous regressors).
 We characterize the identified set using a collection of conditional moment inequalities, based on which estimation and inference can be conducted using established econometric methods in the literature, such as Chernozhukov, Hong, and Tamer (2007), Andrews and Shi (2013), and Chernozhukov, Lee, and Rosen (2013).
@@
 Our paper contributes directly to the line of econometric literature on semiparametric panel discrete choice models.
-Dating back to Manski (1987), a series of work exploits “full” stationarity conditions for identification, such as Chernozhukov et al.
-(2013), Khan, Ponomareva, and Tamer (2016), Shi, Shum, and Song (2018), Pakes and Porter (2022), Khan, Ouyang, and Tamer (2021), Khan, Ponomareva, and Tamer (2023), Gao and Li (2020), and Wang (2022).
+Dating back to Manski (1987), a series of work exploits “full” stationarity conditions for identification, such as Chernozhukov, Lee, and Rosen (2013), Khan, Ponomareva, and Tamer (2016), Shi, Shum, and Song (2018), Pakes and Porter (2022), Khan, Ouyang, and Tamer (2021), Khan, Ponomareva, and Tamer (2023), Gao and Li (2020), and Wang (2022).
 As discussed above, full stationarity conditions given all observable covariates effectively require that all covariates are exogenous with no dynamic effects (i.e., lagged dependent variables).
 In contrast, we exploit the “partial” stationarity condition, thereby allowing for lagged dependent variables as well as other endogenous covariates.
-In the literature on dynamic discrete choice models, our paper is most closely related to Khan, Ponomareva, and Tamer (2023, KPT thereafter), who studies the following dynamic panel binary choice model
-$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (4)
-where the one-period lagged dependent variable $Y_{i,t-1}\in\{0,1\}$ serves as the endogenous covariate, and $Z_{it}$ are exogenous covariates. KPT exploits the following assumption:
-$\epsilon_{is}\mid Z_{it},Z_{is},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{it},Z_{is},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
-which is exactly a “partial stationarity” condition in the specific context of (4), and derives the sharp identified set for $\theta_{0}$ by explicitly enumerating the realizations of the one-period lagged outcome variable $Y_{i,t-1}$ (across two periods $t,s$).
+In the literature on dynamic discrete choice models, our paper is most closely related to Khan, Ponomareva, and Tamer (2023, KPT thereafter), who studies the following dynamic panel binary choice model:
+$Y_{it}=\mathbf{\mathbbm{1}}\{Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (4)
+where the one-period lagged dependent variable $Y_{i,t-1}\in\{0,1\}$ serves as the endogenous covariate, and $Z_{it}$ are exogenous covariates.
+KPT exactly imposes the “partial stationarity” condition (3) in the specific context of (4), and derives the sharp identified set for $\theta_{0}$ by explicitly enumerating the realizations of the one-period lagged outcome variable $Y_{i,t-1}$ (across two periods $t,s$).
 In contrast, our model (1), along with the “partial stationarity” condition, is stated with more general specifications of the endogenous covariates $X_{it}$.
@@
 For example, Aristodemou (2021) exploits a type of “full independence” assumption to identify dynamic binary choice models.
-The “full independence” assumption essentially requires that the time-varying errors from all time periods and the exogenous variables from all time periods are independent (conditional on initial conditions), but does not make intertemporal restrictions on the errors (such as stationarity”.
+The “full independence” assumption essentially requires that the time-varying errors from all time periods and the exogenous variables from all time periods are independent (conditional on initial conditions), but does not make intertemporal restrictions on the errors (such as stationarity).
 Hence, such ‘full independence” assumption and the partial stationarity assumption in our paper do not nest each other as special cases.
@@
 The rest of the paper is organized as follows.
-Section 2 studies the sharp identification of binary choice models with endogenous covariates.
-We provide sufficient conditions to achieve point identification and explore two classic models: static and dynamic models.
-Sections 3-5 apply the identification approach to ordered response models, multinomial choice models, and models with censored dependent variables.
-Section 6 broadens the scope to general nonseparable models.
+Section 2 studies the sharp identification of panel binary choice models with endogenous covariates.
+Sections 3 demonstrates how our key identification strategy generalizes to a wide range of dynamic nonlinear panel data models, such as ordered response models, multinomial choice models, and censored outcome models.
 Section 4 presents simulation results about the finite-sample performances of our approach and Section 5 explores the empirical application of income categories using various ordered response models.
@@
 2.1 Model
-In this section, we focus on the identification of dynamic binary choice models with endogenous covariates. Specifically, consider the following binary choice model:
-$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (5)
-where $Y_{it}\in\{0,1\}$ denotes the binary dependent variable for individual $i\in\{1,2,...\}$ and time $t\leq T$, $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$.
-The objective is to identify the parameter $\theta_{0}:=(\beta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime}$ using a short panel of observed variables $(Z_{it},X_{it},Y_{it})_{it}$, where the number of time periods $T\geq 2$ is fixed and finite.
-The identification of model (5) has been explored in the literature under various assumptions.
-[^5] See Section 1 for more detailed discussions on related literature.
-Specifically, Manski (1987) exploits the following conditional stationarity (also referred to as “group homogeneity”) of $\epsilon_{it}$ to achieve identification:
-$\epsilon_{is}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (6)
-The above condition does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows for dependence of $\epsilon_{it}$ over time.
-However, condition (2) does impose substantial restriction on the dependence between $(Z_{it},X_{it})$ and the time-varying error $\epsilon_{it}$.
-For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=Z_{it}^{\prime}\eta_{1}+X_{it}^{\prime}\eta_{2}$, then condition (2) fails in general.
-Hence, condition (2) can also be interpreted as a form of exogeneity condition on $\epsilon_{it}$ and all covariates $(Z_{it},X_{it})$.
-Importantly, in a dynamic setting where $X_{it}$ includes the lagged dependent variable $Y_{i,t-1}$, then condition (2) is no longer a justifiable assumption to work with.
-Our paper aims to study identification of $\theta_{0}$ robust to endogeneity of $X_{it}$, where $X_{it}$ can be arbitrarily dependent with $(\alpha_{i},\epsilon_{it})$.
-Before introducing our assumptions, we first provide several applications to illustrate this endogeneity.
-Example 1 (Dynamic Models).
-Khan, Ponomareva, and Tamer (2023) studies the following dynamic model with one lagged dependent variable:
-$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+Y_{it-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
-under the stationarity assumption only given exogenous covariates $Z_{it}$: $\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}$.
-In this model, the full stationarity condition may naturally fail due to its dynamic nature.
-For example, suppose $\epsilon_{it}$ is i.i.d.
-across time and independent of ($(\{Z_{it}\}_{t=1}^{T},\alpha_{i})$.
-Given $Z_{is}=Z_{it}=z,\alpha_{i}=a$, it is clear that the conditional distribution of $\epsilon_{it}\mid Y_{it-1},Y_{it},z,a$ will be different from $\epsilon_{it+1}\mid Y_{it-1},Y_{it},z,a\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it+1}$, since $Y_{it}$ is informative for $\epsilon_{it}$.
-Example 2 (Omitted Variable).
-Consider the true model is given as $Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+S_{it}\eta_{0}+\alpha_{i}+v_{it}\geq 0\}$, where $v_{it}$ is independent of $(\alpha_{i},\{Z_{it},X_{it},S_{it}\}_{t=1}^{T})$.
-However, the variable $S_{it}$ is omitted and we instead estimate the following model:
-$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
-where $\epsilon_{it}=v_{it}+S_{it}\eta_{0}$.
-The covariate $X_{it}$ can be correlated with $\epsilon_{it}$ through the omitted variable $S_{it}$, provided that the omitted variable $S_{it}$ is correlated with $X_{it}$.
-Then the conditional distribution of $\epsilon_{it}\mid(Z_{is},Z_{it},X_{is},X_{it},\alpha_{i})$ may differ from $\epsilon_{is}\mid(Z_{is},Z_{it},X_{is},X_{it},\alpha_{i})$ since the distribution of $S_{it}\mid X_{it}$ could be different from the distribution $S_{is}\mid X_{is}$ when $X_{it}\neq X_{is}$.
-Example 3 (Measurement Error).
-Consider the true model is given as $Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{*^{\prime}}\gamma_{0}+\alpha_{i}+v_{it}\geq 0\}$, where $v_{it}$ is independent of $\left(\alpha_{i},\{Z_{it},X_{it}^{*}\}_{t=1}^{T}\right)$.
-However, the covariate $X_{it}^{*}$ may not be observed and we only observe its measurement $X_{it}$ which may be subject to a measurement error $X_{it}=X_{it}^{*}+e_{it}$.
-Suppose we estimate the following model using the measurement $X_{it}$ as a regressor:
-$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$
-where $\epsilon_{it}=v_{it}-e_{it}^{\prime}\gamma_{0}$.
-The error term $\epsilon_{it}$ would be correlated with the measurement $X_{it}$ through the measurement error $e_{it}$.
-Similarly, the conditional stationarity in (2) could fail since $X_{it}$ is informative for $e_{it}$ such that the conditional distribution of $\epsilon_{it}\mid z,X_{is},X_{it},\alpha_{i}$ would vary over time given different covariates $X_{is}\neq X_{it}$ while holding $Z_{is}=Z_{it}=z$.
-Motivated by the above examples, this paper aims to establish identification of $\theta_{0}$, while being robust to endogeneity of $X_{it}$ with respect to unobserved factors $(\alpha_{i},\epsilon_{it})$.
-Below we propose an approach that accommodates various types of endogeneity, regardless of whether the endogenous covariates $X_{it}$ are discrete lagged outcome variables (as in Example 1) or continuously-valued covariates with contemporary endogeneity as in other examples.
-To achieve this, we introduce the following partial conditional stationarity assumption:
+We start with an illustration of our general identification strategy in the binary choice setting. Specifically, consider the following binary choice model:
+$Y_{it}=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\},$ (5)
+where a panel of data $\left(Y_{it},Z_{it},X_{it}\right)$ are observed across individuals $i=1,...,n$ and time periods $t=1,...,T$.
+Here $Y_{it}$ denotes the binary outcome variable, $Z_{it}$ and $X_{it}$ denote two types of observed covariates (the difference between $Z_{it}$ and $X_{it}$ will be clarified below), $\alpha_{i}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$.
+The objective is to identify the unknown parameter $\theta_{0}:=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$, with $\beta_{0}\in\mathcal{R}^{d_{z}}$ and $\gamma_{0}\in\mathcal{R}^{d_{x}}$.
+We focus on the short panel setting, where the number of individuals $n$ is considered to be larg while the number of time periods $T\geq 2$ is fixed and finite.
+Write $Z_{i}:=\left(Z_{i1},...,Z_{iT}\right)$ and $X_{i}:=\left(X_{i1},...,X_{iT}\right)$.
+Throughout this paper, we will refer to $Z_{i}$ as “exogenous covariates”, and refer to $X_{i}$ as “endogenous covariates”.
+The exact difference between $Z_{i}$ and $X_{i}$, as well as the precise meaning of “exogeneity” versus “endogeneity” in our context, are formalized through the following “partial stationarity” assumption: the word “partial” emphasizes that this assumption is imposed on only a part of the observed covariates $Z_{i}$, but not on $X_{i}$.
 Assumption 1 (Partial Stationarity).
-The conditional distribution of $\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}$ is stationary over time:
-$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
-Assumption 1 only assumes the stationarity given the covariate $(Z_{is},Z_{it})$, but does not impose any restrictions on the endogenous covariate $X_{it}$.
-The standard full stationarity condition (2) is nested in Assumption 1, and it is in general stronger than Assumption 1.
-The main distinction is that Assumption 1 allows $X_{it}$ to be arbitrarily dependent with both the fixed effect $\alpha_{i}$ and the time-changing error $\epsilon_{it}$.
-In addition, their correlation can be different across time.
-Assumption 1 also shares similar features with condition (2), in that it does not impose any distributional assumption on $\epsilon_{it}$ and allows $\epsilon_{it}$ to be correlated across time.
-The standard i.i.d.
-assumption of $\epsilon_{it}$ across time or independence of $\epsilon_{it}$ with $(\alpha_{i},Z_{it},X_{it})$ are also nested in Assumption 1 as special cases.
-2.2 Identification Strategy
+The conditional distribution of $\epsilon_{it}\mid Z_{i},\alpha_{i}$ is stationary over time, i.e.,
+$\epsilon_{it}\mid Z_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid Z_{i},\alpha_{i}\quad\forall t,s=1,...,T.$
+Assumption 1 essentially requires that the (conditional) distribution of $\epsilon_{it}$ stays the same across all time periods $t=1,...,T$ even if $Z_{i}$ realize to different values.
+To illustrate, suppose that there are only two periods $t=1,2$, and that $Z_{i1},Z_{i2}$ realize to two values $z_{1},z_{2}$, respectively, with $z_{1}<z_{2}$.
+Then Assumption 1 requires that $\epsilon_{i1}$ and $\epsilon_{i2}$ still have the same (conditional) distributions: in particular, $\epsilon_{i1}$ cannot be stochastically smaller (or larger) than $\epsilon_{i2}$ because of $z_{1}<z_{2}$.
+Hence, Assumption 1 can be thought as a definition of the “exogeneity” of the covariates $Z_{it}$ in our context.
+In contrast, Assumption 1 imposes no such restrictions on the (potentially) endogenous covariates $X_{i}$.
+In fact, since $X_{i}$ does not appear in Assumption 1 at all, here we are completely agnostic about the dependence structure between $\epsilon_{i1},...,\epsilon_{iT}$ and $X_{i}$: in particular, the conditional distribution of $\epsilon_{it}$ is allowed to vary across $t$ arbitrarily for any particular realization of $X_{i}$.
+As a result, different forms of endogeneity in $X_{i}$ can be incorporated under our framework in a unified manner, as we illustrate in the examples below.
+Example 1 (Dynamic Effects via Lagged Outcomes).
+Consider the following “AR(1)-type” dynamic binary choice model studied in Khan, Ponomareva, and Tamer (2023, KPT thereafter):
+$Y_{it}=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\},$
+which is a special case of our model with $X_{it}$ set to be the one-period lagged binary outcome variable $Y_{i,t-1}$.
+Here, $X_{it}$ is endogenous since $X_{it}\equiv Y_{i,t-1}$ and $\epsilon_{i,t-1}$ is by construction positively correlated with $Y_{i,t-1}$ for any $t$, and thus the distribution of $\epsilon_{it}$ cannot be stationary across $t$ when conditioned on the realizations of $Y_{i0},...,Y_{i,T-1}$.
+For example, given $Y_{i0}=Y_{i1}=1,$ $Y_{i2}=0$ (and $Z_{i},\alpha_{i}$), the conditional distribution of $\epsilon_{i1}$ will naturally be different from that of $\epsilon_{i2}$.
+To obtain identification under the endogeneity of $Y_{i,t-1}$, KPT imposes the stationarity of $\epsilon_{it}$ conditional on the exogenous covariates $Z_{i}$ only, which coincides with our “partial stationarity” condition (Assumption 1) when specialized to their setting.
+A natural generalization of the AR(1) model above in KPT is the following “AR($p$)” model, which is again a special case of our model with $X_{it}$ taken to be the vector of $p$ lagged outcomes $Y_{i,t-1},...,Y_{i,t-p}$:
+$Y_{it}=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+\sum_{j=1}^{p}Y_{i,t-j}\gamma_{j}+\alpha_{i}+\epsilon_{it}\geq 0\right\}.$
+Similarly, $X_{it}$ is endogenous here due to dependence on $\epsilon_{i,t-1},...,\epsilon_{t-p}$, which can again be handled in our framework under the “partial stationarity” assumption.
+While it is not clear how the identification results in KPT can be easily generalized to the AR($p$) model above, we show in the next subsection how our identification strategy provides a simple and unified approach to derive moment inequalities regardless of the exact specifications of $X_{it}$.
+Example 2 (Contemporarily Endogenous Covariates).
+Alternatively, consider the following binary choice model with contemporarily endogenous covariates:
+$Y_{it}$ $=\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\},$
+$X_{it}$ $=\phi\left(Z_{it},u_{it}\right)$
+where $\phi$ is an unknown “first-stage” function and the “first-stage error” $u_{it}$ is allowed to be arbitrarily correlated with $\epsilon_{it}$.
+For example, $X_{it}$ may be a “price-type” variable that is strategically chosen by a decision maker after observing the current-period error $\epsilon_{it}$, which generates contemporary dependence between $X_{it}$ and $\epsilon_{it}$.
+Even though contemporary endogeneity of this type is very different in nature from the dynamic endogeneity discussed in the previous example, it also induces non-stationarity of $\epsilon_{it}$ when conditioned on $X_{i}$: for example, if $X_{it}$ and $\epsilon_{it}$ are positively correlated, then, conditional on $X_{i1}<X_{i2}$, it is unreasonable to assume the distribution of $\epsilon_{i1}$ is the same as $\epsilon_{i2}$.
+That said, such type of contemporary endogeneity can also be handled in our framework under the “partial stationarity” condition (Assumption 1).
+Remark 1 (Combination of Dynamic and Contemporary Endogeneity).
+We separately discussed two types of endogenous covariates, dynamic covariates (lagged outcome variables) and contemporarily endogenous covariates, in the two examples above, but our identification strategy also applies if both types of endogenous covariates are present together, since our identification strategy works generally under “partial stationarity”, which does not impose or exploit any restrictions on the form of endogeneity between $\epsilon_{it}$ and $X_{i}$.
+Remark 2 (Full Stationarity as Special Case).
+Obviously, the standard “full stationarity” condition (2) is nested under “partial stationarity” condition (Assumption 1) as a special case, where the endogenous covariate $X_{it}$ contains no variables.
+Hence, “full stationarity” is in general stronger than “partial stationarity”.
+Remark 3 (Focus on Time-Varying Endogeneity).
+Technically, our partial stationarity condition also allows some endogeneity between $\epsilon_{it}$ and $Z_{i}$, as long as such endogeneity is time-invariant.
+This is because Assumption 1 is stated conditional on the full vector $Z_{i}=\left(Z_{i1},...,Z_{iT}\right)$ and the time-invariant fixed effect $\alpha_{i}$.
+Hence, as long as the conditional distribution of $\epsilon_{it}$ depends on $Z_{i1},...,Z_{iT}$ and $\alpha_{i}$ in a time-invariant manner, the stationarity of $\epsilon_{it}$ can still hold.
+That said, since in empirical applications we are mostly interested in “time-varying endogeneity”, such as the dynamic and contemporary endogeneity discussed in the examples above, in this paper we refer to $Z_{i}$ as “exogenous” even though it may be endogenous in a time-invariant manner, and only call $X_{i}$, which features time-varying endogeneity, the “endogenous” covariates.
+Remark 4 (Pairwise Version of Partial Stationarity).
+In Assumption 1, we impose partial stationarity of $\epsilon_{it}$ conditional on $Z_{it}$ from all periods $t=1,...,T$.
+Alternatively, we could impose partial stationarity in a “pairwise” version, conditional on $\left(Z_{it},Z_{is}\right)$ from any pair of time periods $\left(t,s\right)$ only:
+[…]
+Clearly, the “pairwise” version is equivalent to the “all-periods” version when $T=2$, but is weaker when $T\geq 3$.
+Our identification strategy applies under both versions of partial stationarity, though the identification results and the corresponding proofs have slightly different representations.
+Essentially, conditioning on all-period covariate realizations would be replaced with conditioning the realizations in any specific pair of period.
+See Remark 5 at the end of Section 2.2 for a follow-up discussion.
+2.2 Key Identification Strategy
 We now explain our key identification strategy based on partial stationarity.
-While it applies more generally, we first illustrate the key idea in the specific context of the dynamic binary choice model.
-Sections 3-6 provide generalizations of the idea to various models.
-Let $v_{it}:=-(\epsilon_{it}+\alpha_{i})$, $W_{it}:=(Z_{it},X_{it})$, and $W_{ist}:=(W_{is},W_{it})$. The conditional choice probability is given as follows:
-$\Pr\left(Y_{it}=1\mid W_{ist}=w_{st}\right)=\Pr\left(v_{it}\leq w_{t}^{\prime}\theta_{0}\mid W_{ist}=w_{st}\right).$
-When assuming all covariates are exogenous under condition (2), variation in the choice probability is solely due to variation in the covariate index $w_{t}^{\prime}\theta_{0}$.
-[^6] Condition 2 implies the stationarity of $v_{it}$: $v_{is}\mid W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid W_{ist}$.
-Therefore, as shown in Manski (1987), the identified set for $\theta_{0}$ is characterized by the condition: an increase in the value of $w_{t}^{\prime}\theta_{0}$ over time implies an increase in the choice probability over time.
-However, Manski (1987)’s method is not applicable to our setting with Assumption 1, given that the endogenous regressor $X_{it}$ can be correlated with $v_{it}$.
-When the covariate $X_{it}$ changes, the conditional distribution of $v_{it}$ might potentially change as well.
-Variation in the choice probability contains a mixture of changes in the covariate index and the conditional distribution of $v_{it}$.
-Hence, it is not feasible to derive identifying restrictions on $\theta_{0}$ merely by examining the intertemporal variation in the choice probability.
-Our analysis relies on the partial stationarity condition in Assumption 1, which implies the stationarity of $v_{it}$ given $Z_{ist}:=(Z_{is},Z_{it})$:
-$v_{is}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid Z_{ist}.$
-The identification strategy proceeds in two steps.
-We first derive lower and upper bounds for the distribution of $v_{it}\mid Z_{ist}$ at each period $t$.
-The derived bounds at each period $t$ are functions of observed variables and the parameter $\theta_{0}$.
-Since the distribution of $v_{it}\mid Z_{ist}$ is the same over time, then the bounds of the distribution across time should have intersections.
-This restriction serves as the identifying restriction for $\theta_{0}$.
-Next, we show how to derive bounds on the distribution of $v_{it}\mid Z_{ist}$.
-Let $F_{v_{t}\mid W_{st}}(\cdot\mid w_{st})$ denote the conditional CDF of $v_{it}$ given $W_{ist}=w_{st}$.
-For any fixed point $a\in\mathcal{R}$, we first derive bounds for the conditional distribution $F_{v_{t}\mid W_{st}}(a\mid w_{st})$, and then establish bounds on $F_{v_{t}\mid Z_{st}}\left(a\mid z_{st}\right)$ by taking expectation over the covariate $X_{ist}$ given $Z_{ist}=z_{st}$.
-Lower bound. The conditional choice probability given $W_{ist}=w_{st}$ is given as
-$\Pr\left(Y_{it}=1\mid W_{ist}=w_{st}\right)=\Pr(v_{it}\leq w_{t}^{\prime}\theta_{0}\mid W_{ist}=w_{st})=F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta_{0}\mid w_{st}).$
-For any fixed point $a$, by the monotonicity of a distribution, $F_{v_{t}\mid W_{st}}(a\mid w_{st})$ can be bounded below using the observed choice probability:
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\Pr(Y_{it}=1\mid W_{ist}=w_{st})\mathbbm{1}\{w_{t}^{\prime}\theta_{0}\leq a\}.$
-The above inequality holds since either $a$ is larger than the covariate index $w_{t}^{\prime}\theta_{0}$ so we know $F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta_{0}\mid w_{st})$ or the right hand side is equal to zero when $a$ is smaller than $w_{t}^{\prime}\theta_{0}$.
-As Assumption 1 is on the distribution of $v_{it}$ given $Z_{ist}$, we can establish the lower bound for $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ by taking expectation over covariates $X_{ist}$ given $Z_{ist}=z_{st}$:
+Write $v_{it}:=-\left(\epsilon_{it}+\alpha_{i}\right)$ so that model (5) can be rewritten as
+$Y_{it}=\mathbf{\mathbbm{1}}\left\{v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\right\}.$
+For any constant $c\in\mathcal{R}$, consider first the event
+$Y_{it}=1\text{ and }Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c.$
+Whenever the event above happens, we must have $v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c$, implying that $v_{it}\leq c$.
+Formally, the above can be summarized by the following inequality:
+$Y_{it}\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right\}$ $=\mathbf{\mathbbm{1}}\left\{v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\right\}\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right\}$ (7)
+$\leq\mathbf{\mathbbm{1}}\left\{v_{it}\leq c\right\}.$
+Symmetrically, we can also consider the “flipped” event
+$Y_{it}=0\text{ and }Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c,$
+which implies $v_{it}>c$, and similarly
+$\left(1-Y_{it}\right)\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}=$ $\mathbf{\mathbbm{1}}\left\{v_{it}>Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\right\}\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}$
 […]
-$=\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\leq a\mid z_{st}).$
-The above lower bound only depends on observed variables $(W_{ist},Y_{it})$ and $\theta_{0}$, so it is identified up to the parameter $\theta_{0}$.
-This bound will help construct identifying restrictions on $\theta_{0}$.
-Upper bound. The idea of constructing the upper bound is similar. For any $a\in\mathcal{R}$, the conditional distribution $F_{v_{t}\mid W_{st}}$ can be bounded above:
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})\leq\Pr(Y_{it}=1\mid W_{ist}=w_{st})\mathbbm{1}\{w_{t}^{\prime}\theta_{0}\geq a\}+\mathbbm{1}\{w_{t}^{\prime}\theta_{0}<a\}.$
-Then, taking expectation over covariates $X_{ist}$ given $Z_{ist}=z_{st}$ leads to the upper bound for $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$:
+which is equivalent to
+$\mathbf{\mathbbm{1}}\left\{v_{it}\leq c\right\}\leq 1-\left(1-Y_{it}\right)\mathbf{\mathbbm{1}}\left\{Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}.$ (8)
+Taking conditional expectations of (7) and (8) given $Z_{i}=z\equiv\left(z_{t}\right)_{t=1}^{T}$,
+$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right|Z_{i}=z\right)$
 […]
-$=\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\Pr(z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st})$
-$=1-\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st}).$
-Given the lower and upper bound for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, we are ready to establish identifying conditions for $\theta_{0}$.
-Since Assumption 1 requires that $v_{t}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{s}\mid Z_{ist}$, it implies the upper bound for the distribution of $v_{s}\mid Z_{ist}$ must be larger than the lower bound for the distribution of $v_{t}\mid Z_{ist}$ otherwise Assumption 1 cannot hold.
-This restriction characterizes an identified set for $\theta_{0}$, presented in the following proposition.
-Proposition 1.
-Under Assumption 1, an identified set $\Theta_{I,1}$ for $\theta_{0}$ is the set of parameters that satisfy the following conditions:
-$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st}),$ (7)
-for any $a\in\mathcal{R}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Proposition 1 characterizes an identified set for $\theta_{0}$, using the conditional joint distribution of $(Y_{it},X_{it})$ given $Z_{ist}$ across any pair of periods $(s,t)$.
-The identification result allows for endogeneity in the covariate $X_{ist}$ and does not rely on any excluded instruments.
-[^7] When a sub-vector $Z_{it}^{exc}$ of $Z_{it}$ is assumed to be excluded from the binary choice model, then the identifying restriction in (7) is adjusted by setting the coefficient of $Z_{it}^{exc}$ as zero: $\beta_{0}^{exc}=0$.
-The identifying condition in (7) can be reformulated as conditional moment inequalities, by substituting the probability $\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})$ with the conditional expectation $\mathbb{E}\left[\mathbf{\mathbbm{1}}\{Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\}\mid z_{st}\right]$.
-Consequently, the estimation and inference for $\theta_{0}$ can be conducted using established methods such as Chernozhukov, Hong, and Tamer (2007), Andrews and Shi (2013), and Chernozhukov, Lee, and Rosen (2013).
+=\ = $\mathbb{P}\left(\left.v_{is}\leq c\right|Z_{i}=z\right)$
+[…]
+where the middle equality follows from the partial stationarity condition (Assumption 1). Specifically, observe that Assumption 1 implies the partial stationarity of $v_{it}$ given $Z_{i}$, i.e.,
+$v_{it}\mid Z_{i}\stackrel{{\scriptstyle d}}{{\sim}}v_{is}\mid Z_{i},$
+so that $\mathbb{P}\left(\left.v_{it}\leq c\right|Z_{i}=z\right)=\mathbb{P}\left(\left.v_{is}\leq c\right|Z_{i}=z\right)$ for any $c\in\mathcal{R}$.
+Essentially, in the above we exploit the joint occurrence of $v_{it}\leq Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}$ and $Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c$ to deduce an implication on the composite error $v_{it}\leq c$ that is free of the endogenous covariates $X_{it}$, and then leverage the partial stationarity of $v_{it}$ given $Z_{i}$ for intertemporal comparisons.
+Since the lower and upper bounds in (9) hold for any $t$ and $s$, we summarize the identifying restrictions (9) across all time periods in the following proposition.
+Proposition 1 (Identified Set).
+Under model (5) and Assumption 1, $\theta_{0}\in\Theta_{I}$, where the identified set $\Theta_{I}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$\max_{t=1,...,T}\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)\leq 1-\max_{s=1,...,T}\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)$ (10)
+for any $c\in\mathcal{R}$ and any realization $z=\left(z_{1},...,z_{T}\right)$ in the support of $Z_{i}$.
+Proposition 1 characterizes the identified set $\Theta_{I}$ for $\theta_{0}$ as restrictions on the conditional joint distribution of $Y_{it}$ and $X_{it}$ given $Z_{i}=z$.
+More specifically, the restrictions in (10) can be regarded as a collection of conditional moment inequalities that relate $\mathbf{\mathbbm{1}}\left\{Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right\}$ conditional on $Z_{i}=z$.
+Proposition 1 holds regardless of whether the endogenous covariates $X_{it}$ are discrete or continuous.
+When $X_{it}$ are continuous (taking a continuum of values), then Proposition 1 requires that condition (10) hold for a continuum of constants $c\in\mathcal{R}$, so that (the information in) the whole joint distribution of the binary variable $Y_{it}$ and the continuous variable $z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma$ can be captured by the collection of joint distributions of $\left(Y_{it},\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right\}\right)$ across all possible cutoff values $c$.
+However, when $X_{it}$ are discrete, such as in the AR$\left(p\right)$ dynamic model where $X_{it}$ consists of $p$ lagged binary outcome variables, there is no need to evaluate (10) at all possible values of $c\in\mathcal{R}$, since the inequalities in (10) can only bind at finitely many values of $c$.
+We formalize this observation via the following Proposition.
+Proposition 2 (Identified Set with Discrete Endogenous Covariates).
+Suppose that the endogenous covariate $X_{it}$ can only take finite number of values in $\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ across all time periods $t=1,...,T$, then $\Theta_{I}=\Theta_{I}^{disc},$ where $\Theta_{I}^{disc}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ that satisfy condition (10) for any
+$c\in\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K,t=1,...,T\right\},$ (11)
+for any $z=\left(z_{1},...,z_{T}\right)$ in the support of $Z_{i}$.
+Proposition 2 shows that the discreteness of the endogenous covariates $X_{it}$ help reduce the infinite number of inequality restrictions in Proposition 1 to finitely many, or more precisely, $KT$ ones (conditional on $z$).
+The case of discrete $X_{it}$ is conceptually important, since it nests the dynamic AR$\left(p\right)$ model widely studied in the literature as a special case.
+Clearly, when $X_{it}$ consists of $p$ (finitely many) lagged binary outcome variables $Y_{i,t-1},...,Y_{i,t-p}$, then $X_{it}$ by construction can only take $K=2^{p}$ discrete values.
+Specialized further to the AR(1) model in KPT, Proposition 2 shows that the identified set $\Theta_{I}$ is characterized by $2T$ conditional restrictions, which is drastically smaller than the $9T\left(T-1\right)$ conditional restrictions listed out in KPT (even when $T$ is small).
 Remark 5.
-Our “partial stationarity” condition accommodates the “full stationarity” condition (2) as a special case.
-Under full stationarity, or in other words, when all covariates are exogenous, Proposition 1 specializes to the following “maximum-score-type” identifying restrictions in Manski (1987):
+Following up on Remark 4, if pairwise partial stationarity is adopted, then Propositions 1 and 2 continue to hold with (10) adapted to the following “pairwise” version:
+$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i,ts}=z_{ts}\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i,ts}=z_{ts}\right),\forall\,(t,s),$ (12)
+where $Z_{i,ts}:=(Z_{it},Z_{is})$ and $z_{ts}:=(z_{t},z_{s}).$ Relative to (10), the statement in (12) reflects the fact that pairwise partial stationarity is imposed on all pairs of time periods separately instead of all $T$ time periods jointly.
+It is straightforward to verify that the identification arguments above, in particular (7)-(9), carry over with all conditional probabilities/expectations taken conditional on $Z_{i,ts}=z_{ts}$ instead of $Z_{i}=z$.
+2.3 Sharpness of Identified Set
+So far we have only shown that $\Theta_{I}$, which equals to $\Theta_{I}^{disc}$ under discreteness of $X_{it}$, is a valid identified set for $\theta_{0}$.
+However, it is not yet clear whether it has incorporated all the available information for $\theta_{0}$.
+The following theorem establishes the sharpness of the identified set $\Theta_{I}$ under the discreteness of $X_{it}$.
+Theorem 1 (Sharpness under Discreteness).
+When $X_{it}\in\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ for all $t=1,...,T$, then under model (5) and Assumption 1, the identified set $\Theta_{I}$ is sharp.
+The formal definition of sharpness, along with the complete proof of Theorem 1, are available in Appendix A.2.
+In short, we show (by construction) that, for each $\theta\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, there exists a data generating process (DGP) that satisfies Assumption 1 and produces the same joint distribution of observable data $\left(Y_{i},X_{i},Z_{i}\right)$ under model (5) and the discreteness of $X_{it}$.
+Theorem 1 demonstrates that our key identification strategy based on the bounding of (endogenous) parametric index by arbitrary constants, as described in Section 2.2, is able to extract all the available information for $\theta_{0}$ from the model and the observable data, and thus it is impossible to further differentiate $\theta_{0}$ from alternatives in the identified set $\Theta_{I}$ under model (5) and our assumption of partial stationarity (without further restrictions).
+Theorem 1 immediately implies that, in the special case of dynamic AR$\left(p\right)$ models where $X_{it}$ consists of discrete lagged outcomes, our characterization of the identified set $\Theta_{I}^{disc}$ in Proposition 2 is sharp.
+In particular, our result generalizes the corresponding result in KPT, which focuses on the AR(1) model.
+Furthermore, KPT characterizes the sharp identified set via $9T\left(T-1\right)$ conditional restrictions, the derivation of which is based on an exhaustive enumeration of lagged outcome realizations $Y_{i,t-1}$.
+In this paper we adopt an entirely different (and much more general) identification strategy, and arrive at a characterization of the identified set by $2T$ conditional restrictions, which we also show to be sharp by Theorem 1.
+Since our model and assumption specialize exactly to that in KPT under the AR(1) specification, it follows that our $2T$ restrictions must be able to reproduce all the $9T\left(T-1\right)$ restrictions in KPT.
+This demonstrates that our identification strategy not only applies more generally than the one in KPT, but also leads to a more elegant characterization of the sharp identified set with much fewer restrictions.
+We provide a more detailed explanation about this point in the next subsection.
+Another conceptually remarkable, or surprising, feature of Proposition 2 and Theorem 1 is that they are established without reference to the exact nature, or interpretation, of the endogenous covariates $X_{it}$.
+The identified set $\Theta_{I}$ we characterized is valid and sharp regardless of whether $X_{it}$ are specified as lagged outcome variables or contemporarily endogenous covariates, as long as $X_{it}$ is discrete.
+Our proof of sharpness consists of two main lemmas.
+First, we show for each $\theta\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$ how to construct the per-period marginal distributions of errors that match the per-period marginal choice probabilities.
+Second, we show how to combine the $T$ per-period marginal distributions into an all-period joint distribution that matches the all-period joint choice probabilities, so that observational equivalence holds.
+The proof techniques we exploited are also different from, and thus novel relative to, those used in the related work that leverages stationarity-type conditions for partial identification, such as Pakes and Porter (2022) for static multinomial choice model and KPT for dynamic AR(1) model.
+Instead of showing existence only, we provide a more explicit construction of the joint distribution of the latent variables, which is valid regardless of the exact type of endogeneity in $X_{it}$.
+In particular, a key challenge in proving sharpness based on stationarity-type conditions lies in that stationarity imposes only aggregate restrictions (via integrals/sums) of the joint distribution of errors, which is rather implicit to work with.
+A key innovation in our proof technique is to show how marginal/aggregate stationarity restrictions and joint choice probability restrictions can be satisfied simultaneously by an explicit, simple, and general construction, which might be of independent and wider use.
+2.4 Reconciliation with Related Work
+Our identifying restrictions in (10) and (11) have a somewhat “nonstandard” representation in terms of (conditional) joint probabilities instead of $Y_{it}$ and $X_{it}$ (given $Z_{i}$), instead of conditional probabilities of $Y_{it}$ given $X_{it}$ (such as lagged outcomes), which are more usually found in the related literature.
+Hence, we provide a more detailed discussion about the content and interpretation of our identifying restrictions, as well as a more explicit explanation of how they relate to exisiting results in the related literature.
+Reconciliation with Manski (1987)
+Consider first the special case where there are no endogenous covariates $X_{it}$, or in other words, $X_{it}$ is degenerate.
+In this case, our “partial stationarity” condition specializes to the “full stationarity” condition (2) as in Manski (1987).
+However, our identifying restriction (10) still has a different form than the identifying restriction in Manski (1987).
+To illustrate, focus on any two periods $\left(t,s\right)$, and observe that in this case our identifying restriction takes the form of
+$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta_{0}\leq c\right|z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta_{0}\geq c\right|z\right),\ \forall c,$ (13)
+while the “maximum-score-type” identifying restrictions in Manski (1987) are of the form
+$z_{s}^{{}^{\prime}}\beta_{0}\geq z_{t}^{{}^{\prime}}\beta_{0}\ \Leftrightarrow\ \mathbb{P}\left(\left.Y_{is}=1\right|z\right)\geq\mathbb{P}\left(\left.Y_{it}=1\right|z\right).$ (14)
+The “maximum-score-type” identifying restriction (14) has a quite intuitive and interpretable representation: across two periods $\left(t,s\right)$ under full stationarity, the conditional choice probability at period $s$ is larger if and only if the index $z_{s}^{{}^{\prime}}\beta_{0}$ is larger.
+In contrast, our restriction (13) has a somewhat twisted representation even in this simple setting.
+However, a closer look reveals that our (13) is exactly equivalent to Manski’s “maximum-score-type” identifying restrictions in the current context.
+To see this, notice that, by setting $c=z_{t}^{{}^{\prime}}\beta_{0}$ in (13), we obtain
+$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)=\,$ $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq z_{t}^{{}^{\prime}}\beta_{0}\right\}$
 […]
-Remark 2.
-In the special case where the only endogenous covariate is one lagged outcome variable, as studied in Khan, Ponomareva, and Tamer (2023), the identifying restrictions yielded by Proposition 1 become
-$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta_{0}+Y_{i,s-1}\gamma_{0}\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq a\mid z_{st}),$
-for any $a\in\{z_{t}^{\prime}\beta_{0},z_{t}^{\prime}\beta_{0}+\gamma_{0},z_{s}^{\prime}\beta_{0},z_{s}^{\prime}\beta_{0}+\gamma_{0}\}$.
-As shown in Appendix A.4, the above identifying condition on $\theta_{0}$ implies the identifying conditions in Khan, Ponomareva, and Tamer (2023, Theorem 1, p.
-8-9), under different values of $(z_{s}^{\prime}\beta_{0},z_{t}^{\prime}\beta_{0},\gamma_{0})$.
-Proposition 1 does not impose restrictions on the support of the endogenous covariate $X_{it}$, allowing it to be either discrete or continuous.
-When $X_{it}$ is continuous, then Proposition 1 requires condition (7) to hold for any $a\in\mathcal{R}$ so it exploits the whole distribution information of $(X_{it},Y_{it})$ at any value $a$.
-For the estimation and inference of $\theta_{0}$, we can follow the conventional inference method to discretize the space of $\mathcal{R}$ and pick $K_{n}$ points for a.
-In the case of discrete $X_{it}$, the dimension of $a$ can be further reduced, and the identified set $\Theta_{I,1}$ can be characterized by finite number of restrictions, shown in the following corollary.
-Corollary 1.
-When the endogenous covariate $X_{it}\in\{a_{1},...,a_{K}\}$ only takes finite number of values, then the identified set $\Theta_{I,1}$ is the set of parameters $\theta$ that satisfy condition (7) for any $a\in\{z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma,z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma\}_{k=1}^{K}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Corollary 1 reduces the number of restrictions in Proposition 1 to $2K$ number of moment restrictions for any pair of two periods $s,t$.
-For the dynamic model where $X_{it}$ includes one lagged dependent variable $X_{it}=Y_{i,t-1}\in\{0,1\}$, there are only four conditional restrictions for $\theta_{0}$.
-When $X_{it}$ includes $j$ lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2},...Y_{i,t-j})$, then there will be $2^{j+1}$ restrictions.
-We show that $\Theta_{I,1}$ is an identified set for $\theta_{0}$, however, it is still unclear whether condition (7) exploits all the available information for $\theta_{0}$.
-The following theorem establishes the sharpness of the identified set $\Theta_{I,1}$ when $X_{it}$ takes finite number of values.
-Theorem 1 (Sharpness).
-Under Assumption 1, the identified set $\Theta_{I,1}$ is sharp when $X_{it}$ takes finite number of values for any $t\leq T$.
-Proposition 1 provides identification results without restrictions on the support of $X_{it}$, and Theorem 1 shows that the results already exhausted all available information for $\theta_{0}$ when $X_{it}$ only takes finite number of values.
-This results implies that for dynamic models with any number of lagged dependent variables, the identified set $\Theta_{I,1}$ is the smallest set we can obtain.
+Hence, if $z_{s}^{{}^{\prime}}\beta_{0}\geq z_{t}^{{}^{\prime}}\beta_{0}$, i.e., the left-hand side of (14) holds, then the above implies that
+$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0\right|z\right)=\mathbb{P}\left(\left.Y_{is}=1\right|z\right),$
+which becomes exactly the right-hand side of (14).
+Switching $t$ with $s$ in the argument above produces the other implication $z_{s}^{{}^{\prime}}\beta_{0}\leq z_{t}^{{}^{\prime}}\beta_{0}$Rightarrow $\mathbb{P}\left(\left.Y_{is}=1\right|z\right)\leq\mathbb{P}\left(\left.Y_{it}=1\right|z\right)$.
+Together these exactly constitute the “if-and-only-if” restriction in (14).
+Hence, even though our inequality restriction (13) looks different from the more intuitive “maximum-score-type” restriction, they both incorporate the same information.
+Reconciliation with KPT (Khan, Ponomareva, and Tamer, 2023)
+Now, consider the dynamic AR(1) model as studied in KPT, where the only endogenous covariate is the one-period lagged outcome variable, i.e., $X_{it}:=Y_{i,t-1}$.
+To illustrate, first focus on any two periods $\left(t,s\right)$ only, and observe that in this case our identifying restriction becomes
+$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq c\right|z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta_{0}+Y_{i,t-1}\gamma_{0}\geq c\right|z\right),\ \forall c.$ (15)
+Under the same model and assumption, KPT derives the following 9 inequality implications for $\left(t,s\right)$: [^6] We adapt the notation in KPT to our current notation, and state these 9 inequalities as strict inequalities, which lead to a simpler and more focused explanation.
+The equivalence between our restriction and the KPT restrictions still hold if their inequalities are stated in the weak form.
+KPT(i): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\left|\gamma_{0}\right|>0.$
+KPT(ii): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{i,s}=0,Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\min\left\{0,\gamma_{0}\right\}>0.$
+KPT(iii): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{i,s}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\max\left\{0,\gamma_{0}\right\}>0.$
+KPT(iv): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\max\left\{0,\gamma_{0}\right\}>0.$
+KPT(v): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}>0.$
+KPT(vi): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\gamma_{0}>0.$
+KPT(vii): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\min\left\{0,\gamma_{0}\right\}>0.$
+KPT(viii): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0\text{$,$}Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\gamma_{0}>0.$
+KPT(ix): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}>0.$
+In a way, the 9 inequality restrictions in KPT above are similar to the “maximum-score restrictions”, in the sense that all of them take the form of logical implications between intertemporal comparisons of various conditional probabilities and intertemporal differences of covariate indexes.
+Using a very different identification strategy than the one in KPT, we arrived at our inequality restriction (15), which looks very different from the collection of 9 inequality restrictions in KPT.
+At first sight it is not clear how (15) relates to and compares with the 9 KPT restrictions.
+However, a closer look again reveals that our restriction (15) can reproduce all the 9 restrictions in KPT, and thus incorporate all the information therein in a unified format.
+Take KPT(v) as an illustration and suppose that the left-hand side of KPT(v) holds, then it implies
+$\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=1\right|z\right).$ (16)
+With $X_{it}=Y_{i,t-1}$, our inequality restriction (15) can be equivalently rewritten as follows,
+$\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}+\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$
+[…]
+where the realization of $Y_{i,t-1}$ is explicitly enumerated as in KPT.
+Note that we can further relax condition (17) by dropping the two probabilities $\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$ and $\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}\geq c\right\}$ as it makes the lower bound smaller and the upper bound larger:
+$\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}\leq\ \\ 1-\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0}\geq c\right\}.$
+Then, the statement that $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0}\geq c\right\}$ both holds is precisely equivalent to the following statement of
+[…]
+By contraposition, it leads to exactly the same implication of KPT(v):
+[…]
+Hence, we have shown that (17) implies KPT(v).
+Similarly, it is shown in Appendix A.3 that (17) implies all 9 restrictions in KPT.
+In fact, the representation (17) reveals why there are precisely 9 KPT-type restrictions.
+The two period-$t$ indicators $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$ in the upper expression of (17) may take 3 “useful” [^7] The 4th combination, $\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}=\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}=0$, will make the upper expression of (17) equal to 00, so that the inequality (17) holds trivially.
+Hence, this $\left(0,0\right)$ combination is not useful.
+combinations $\left(1,0\right),\left(0,1\right)$ and $\left(1,1\right)$, while the two period-$s$ indicators $\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0}\geq c\right\}$ and $\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}\geq c\right\}$ in the lower expression of (17) may also take 3 useful combinations.
+Consequently, in total there are $3\times 3=9$ useful combinations, which exactly correspond to the $9$ left-hand-side suppositions in the 9 KPT restrictions.
+Hence, while our restriction (15) appears very different from the 9 KPT restrictions, it actually automatically incorporates all the KPT restrictions.
+In particular, by treating the endogenous covariate $X_{it}=Y_{i,t-1}$ as a random variable, our restriction (15) automatically aggregates the identifying information across all possible realizations of $Y_{i,t-1}$, without the need to explicitly consider each possibility separately.
+Now, consider a general setting with $T\geq 2$ periods.
+By our Proposition 2 and Theorem 1, the sharp identified set can be characterized by $2T$ restrictions, which are generated by evaluating (10) at each $c$ of the $2T$ points in $\left\{z_{t}^{{}^{\prime}}\beta,z_{t}^{{}^{\prime}}\beta+\gamma:t=1,...,T\right\}$.
+In contrast, across $T$ periods the KPT approach produces $9T\left(T-1\right)$ restrictions, which are generated by imposing the $9$ KPT restrictions across all ordered time pairs $\left(t,s\right)$.
+Hence, our approach provides a much simpler characterization of the sharp identified set, using a significantly smaller number of restrictions.
+For example, with $T=2$ periods, we have $4$ restrictions while KPT has $18$; with $T=3$, we have $6$ restrictions while KPT has 54.
+Hence, the reduction in the number of restrictions relative to KPT is quite remarkable.
+In summary, while the representation of our identifying restrictions in (10) and (11) may appear somewhat unnatural in the first place, it actually becomes equivalent to more familiar (and intuitive) representations in the specialized settings of Manski (1987) and KPT.
 2.5 Point Identification
@@
 This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$ under support conditions on the exogenous covariate $Z_{it}$.
-We focus on the scenario where the endogenous covariate $X_{it}$ is discrete $X_{it}\in\mathcal{X}=\{a_{1},a_{2},...,a_{K}\}$ and there are only two periods $T=2$.
+We focus on the scenario where the endogenous covariate $X_{it}$ is discrete $X_{it}\in\mathcal{X}\equiv\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ and there are only two periods $T=2$.
 To point identify $\beta_{0}$, the first step is to determine the sign of the covariate index $(Z_{i2}-Z_{i1})^{\prime}\beta_{0}$ under certain variation of observed choice probability.
 To identify the sign of $(Z_{i2}-Z_{i1})^{\prime}\beta_{0}$, we define the following two sets:
-$\mathcal{Z}_{1}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\Pr(Y_{i1}=0,X_{i1}=x\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=x\mid z_{12})\Big{\}},$
-$\mathcal{Z}_{2}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\Pr(Y_{i1}=1,X_{i1}=x\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=x\mid z_{12})\Big{\}}.$
+$\mathcal{Z}_{1}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\mathbb{P}(Y_{i1}=0,X_{i1}=x\mid z)<\mathbb{P}(Y_{i2}=1,X_{i2}=x\mid z)\Big{\}},$
+$\mathcal{Z}_{2}:=\Big{\{}(z_{1},z_{2})\mid\exists x\in\mathcal{X}\text{ s.t. }1-\mathbb{P}(Y_{i1}=1,X_{i1}=x\mid z)<\mathbb{P}(Y_{i2}=0,X_{i2}=x\mid z)\Big{\}}.$
 Let $\mathcal{Z}:=\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$. Let $\Delta Z_{i}=Z_{i2}-Z_{i1}$ and $\Delta\mathcal{Z}$ be defined as
@@
 […]
-$1-\Pr(Y_{i1}=0,X_{i1}=x_{1}\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=x_{2}\mid z_{12})\Big{\}};$
+$1-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{1}\mid z)<\mathbb{P}(Y_{i2}=1,X_{i2}=x_{2}\mid z)\Big{\}};$
 […]
-$1-\Pr(Y_{i1}=1,X_{i1}=x_{1}\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=x_{2}\mid z_{12})\Big{\}}.$
+$1-\mathbb{P}(Y_{i1}=1,X_{i1}=x_{1}\mid z)<\mathbb{P}(Y_{i2}=0,X_{i2}=x_{2}\mid z)\Big{\}}.$
 From the identifying results in Proposition 1, the value of $(x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}$ can be bounded when $(z_{1},z_{2})$ belong to the two sets:
-$(z_{1},z_{2})\in\mathcal{Z}^{j}_{3}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0},$
-$(z_{1},z_{2})\in\mathcal{Z}^{j}_{4}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}.$
+$(z_{1},z_{2})\in\mathcal{Z}_{3}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}<\Delta z^{\prime}\beta_{0},$
+$(z_{1},z_{2})\in\mathcal{Z}_{4}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}>\Delta z^{\prime}\beta_{0}.$
 Then the sign of $\gamma_{0}^{j}$ is identified if either the sign of $\Delta z^{\prime}\beta_{0}$ is identified as negative when $(z_{1},z_{2})\in\mathcal{Z}_{2}$ or as positive when $(z_{1},z_{2})\in\mathcal{Z}_{1}$.
 Proposition 4.
-Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}^{j}_{3}\cap\mathcal{Z}_{2}\neq\emptyset$ or $\mathcal{Z}^{j}_{4}\cap\mathcal{Z}_{1}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
-When the endogenous variable $X_{it}$ is a scalar, e.g., the lagged dependent variable $X_{it}=Y_{i,t-1}$, then the definition of the two sets $\mathcal{Z}^{j}_{3},\mathcal{Z}^{j}_{4}$ can be simplified as there existing $x_{1}\neq x_{2}$ such that the corresponding choice probability is observed.
+Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}_{3}^{j}\cap\mathcal{Z}_{2}\neq\emptyset$ or $\mathcal{Z}_{4}^{j}\cap\mathcal{Z}_{1}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
+When the endogenous variable $X_{it}$ is a scalar, e.g., the lagged dependent variable $X_{it}=Y_{i,t-1}$, then the definition of the two sets $\mathcal{Z}_{3}^{j},\mathcal{Z}_{4}^{j}$ can be simplified as there existing $x_{1}\neq x_{2}$ such that the corresponding choice probability is observed.
 Besides the sign of $\gamma_{0}$, the identification results can also bound the value of $\gamma_{0}$ from variation in the exogenous covariates.
-When $X_{it}$ is multi-dimensional such as including two lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2})$ with $\gamma_{0}=(\gamma_{0}^{1},\gamma_{0}^{2})$, then $\gamma_{0}^{1}$ is identified when the required choice probability in the two sets $\mathcal{Z}^{1}_{3},\mathcal{Z}^{1}_{4}$ are observed for $(Y_{i,1},Y_{i,0})=(1,1),(Y_{i,2},Y_{i,1})=(0,1)$ or $(Y_{i,1},Y_{i,0})=(0,0),(Y_{i,2},Y_{i,1})=(1,0)$.
+When $X_{it}$ is multi-dimensional such as including two lagged dependent variable $X_{it}=(Y_{i,t-1},Y_{i,t-2})$ with $\gamma_{0}=(\gamma_{0}^{1},\gamma_{0}^{2})$, then $\gamma_{0}^{1}$ is identified when the required choice probability in the two sets $\mathcal{Z}_{3}^{1},\mathcal{Z}_{4}^{1}$ are observed for $(Y_{i,1},Y_{i,0})=(1,1),(Y_{i,2},Y_{i,1})=(0,1)$ or $(Y_{i,1},Y_{i,0})=(0,0),(Y_{i,2},Y_{i,1})=(1,0)$.
 We provide general sufficient conditions to identify the sign of $\gamma_{0}$, which may be stronger than necessary and can be relaxed in certain scenarios.
-For example, when we know that $\gamma_{0}^{1}+\gamma_{0}^{2}>0$ while $\gamma_{0}^{1}<0$, we can infer that $\gamma_{0}^{2}>0$ without requiring additional assumptions on the two sets $\mathcal{Z}^{2}_{3},\mathcal{Z}^{2}_{4}$.
-3 Ordered Response Model
-3.1 Model and Identification
-Consider a setting where the outcome variable $Y_{it}$ is discrete and takes $J$ values: $Y_{it}\in\{y_{1},..,y_{J}\}$.
-For example, $Y_{it}$ can be different categories of income, health outcomes, or educational attainment.
-We study the following panel ordered response model:
-$Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it},$ (8)
+For example, when we know that $\gamma_{0}^{1}+\gamma_{0}^{2}>0$ while $\gamma_{0}^{1}<0$, we can infer that $\gamma_{0}^{2}>0$ without requiring additional assumptions on the two sets $\mathcal{Z}_{3}^{2},\mathcal{Z}_{4}^{2}$.
+3 Generalization
+3.1 General Identification Strategy
+The key idea underlying our identification strategy generalizes further beyond the binary choice model.
+In this section, we study a general semiparametric model that can accommodate a wide range of panel data models.
+This general model allows for various types of dependent variables, including ordered, multinomial, and censored outcomes.
+Moreover, we explore more flexible specifications for the dependent variable, allowing for multidimensional fixed effects and time-varying errors, as well as nonseparable interactions between them.
+To illustrate, consider the following nonseparable semiparametric model:
+$Y_{it}$ $=G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right),$ (18)
+where $Y_{it}\in\mathcal{Y}$ can be either a discrete or continuous variable, $\alpha_{i}$ is the individual fixed effect of arbitrary dimension, $\epsilon_{it}$ is the time-varying error of arbitrary dimension, $W_{it}$ is a vector of observable covariates, $\theta_{0}\in\mathcal{R}^{d}$ is a conformable vector of parameter, and the function $G$ is allowed to be unknown, nonseparable but assumed to be weakly monotone in the the parametric index:
+Assumption 3 (Monotonicity).
+The mapping $\delta\longmapsto G\left(\delta,\alpha,\epsilon\right)$ is weakly increasing in $\delta\in\mathcal{R}$ for each realization of $\left(\alpha,\epsilon\right)$.
+Note that, by setting $\alpha_{i},\epsilon_{it}$ to be scalar-valued, and $G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\mathbf{\mathbbm{1}}\left\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\}$, we obtain the binary choice model in Section 2, where $G$ is by construction weakly increasing in $W_{it}^{{}^{\prime}}\theta_{0}$.
+Beyond binary choice model, since the dependent variable $Y_{it}$ is not restricted, we show that this general model can accommodate a wide range of panel models with various types of dependent variables.
+Example 3 (Ordered Response Model).
+Consider that the dependent variable $Y_{it}$ can take $J$ possible ordered values: $Y_{it}\in\{y_{1},...,y_{J}\}$ with $y_{j}<y_{j+1}$, and it is generated as follows:
+$Y_{it}^{*}$ $=W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it},$
 $Y_{it}$ $=\sum_{j=1}^{J}y_{j}\mathbbm{1}\{b_{j}<Y_{it}^{*}\leq b_{j+1}\},$
-where $Y_{it}^{*}$ denotes the latent dependent variable, $b_{1}=-\infty,b_{J+1}=+\infty$, and the remaining threshold parameters $b_{j}$ (where $b_{j}\leq b_{j+1}$) can be either known or unknown for $12\leq j\leq J-$ 1.
+where the threshold parameters satisfy $b_{1}=-\infty,b_{J+1}=\infty$, and $b_{j}\leq b_{j+1}$.
+Then the function $G$ is given as
+$G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\sum_{j=1}^{J}y_{j}\mathbf{\mathbbm{1}}\left\{b_{j}<W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\leq b_{j+1}\right\}$
+and it is weakly monotone in $W_{it}^{{}^{\prime}}\theta_{0}$ since $y_{j}<y_{j+1}$ for any $11\leq j\leq J-$ 1.
+Example 4 (Multinomial Choice Model).
+Consider that the dependent variable $Y_{it}$ takes $J$ possible unordered choices: $Y_{it}\in\mathcal{J}=\{0,1,...,J\}$.
+The latent utility $u_{ijt}$ for each option $j$ and the dependent variable $Y_{it}$ are generated as follows:
+$u_{ijt}$ $=W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt},$
+$Y_{it}$ $=\arg\max_{j\in\mathcal{J}}u_{ijt}.$
+Although the $J$ choices are unordered and cannot be directly compared, we can construct a new variable $\tilde{Y}_{it}^{K}=\mathbf{\mathbbm{1}}\{Y_{it}\in K\}$ for any subset $K\subset\mathcal{J}$, representing individuals’ choice belonging to the set $K$.
+It is equivalent to the event that there exists one choice from the set $K$ yielding higher utility than other choices:
+$\tilde{Y}_{it}^{K}$ $=\mathbf{\mathbbm{1}}\{W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt}\geq W_{ikt}^{\prime}\theta_{0}+\alpha_{ik}+\epsilon_{ikt},\ \forall j\in K,k\in\mathcal{J}\backslash K\}$
+$:=G\left((W_{ijt}-W_{ikt})^{\prime}\theta_{0},\alpha_{ij}-\alpha_{ik},\epsilon_{ijt}-\epsilon_{ikt},\ \forall j\in K,k\in\mathcal{J}\backslash K\right).$
+Given this new variable $\tilde{Y}_{it}^{K}$, the function $G$ is weakly monotone in $(W_{ijt}-W_{ikt})^{\prime}\theta_{0}$ for $j\in K,k\in\mathcal{J}\backslash K$ and for any set $K\subset\mathcal{J}$, and we have a similar monotonicity structure.
+Example 5 (Censored Outcome Model).
+Consider the dependent variable $Y_{it}$ is censored at 0, given as follows:
+$Y^{*}_{it}$ $=W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it},$
+$Y_{it}$ $=\max\{Y^{*}_{it},0\},$
+and the function $G(W_{it}^{{}^{\prime}}\theta_{0},\alpha_{i},\epsilon_{it})=\max\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it},0\}$ clearly satisfies the monotonicity in $W_{it}^{{}^{\prime}}\theta_{0}$.
+We have shown that the semiparametric model with monotonicity can accommodate various types of panel data models.
+Now we describe our identification strategy for this general model.
+As before, we decompose the covariate $W_{it}$, and correspondingly $\theta_{0}$, into two components, $W_{it}=\left(Z_{it}^{{}^{\prime}},X_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $\theta_{0}=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$so that
+$W_{it}^{{}^{\prime}}\theta_{0}=Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},$
+where $Z_{it}$ denotes exogenous covariates while $X_{it}$ denotes endogenous covariates, with the precise definition of exogeneity encoded by the partial stationarity in Assumption 1.
+We show how partial stationarity can be exploited in conjunction with weak monotonicity (Assumption 3) to obtain identifying restrictions in the presence of endogeneity.
+Let $\mathcal{Y}$ denote the support of $Y_{it}$. For any $c\in\mathcal{R}$ and $y\in\mathcal{Y}$, conditional on $W_{ist}=w_{st}$, we consider the event that
+$Y_{it}\leq y,\quad\quad Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c,$
+implying that
+$\mathbf{\mathbbm{1}}\{Y_{it}\leq y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\}$ $=\mathbf{\mathbbm{1}}\left\{G\left(Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},\,\alpha_{i},\,\epsilon_{it}\right)\leq y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\right\}$
+$\leq\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\},$
+where the inequality holds by the monotonicity assumption on the function $G$.
+Symmetrically, we can provide an upper bound for $\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\}$ by looking at the following event:
+$\mathbf{\mathbbm{1}}\{Y_{it}>y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\}$ $=\mathbf{\mathbbm{1}}\left\{G\left(Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},\,\alpha_{i},\,\epsilon_{it}\right)>y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\right\}$
+$\leq\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)>y\}$
+$=1-\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\}.$
+which is equivalent to
+$\mathbf{\mathbbm{1}}\{G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\}\leq 1-\mathbf{\mathbbm{1}}\{Y_{it}>y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\leq c\}.$
+The partial stationarity assumption in 1 implies that
+$\mathbb{P}\left(Y_{it}\leq y,\ Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}\geq c\mid Z_{i}=z\right)$
+$=$ $\mathbb{P}\left(G\left(c,\,\alpha_{i},\,\epsilon_{it}\right)\leq y\mid Z_{i}=z\right)$
+$=$ $\mathbb{P}\left(G\left(c,\,\alpha_{i},\,\epsilon_{is}\right)\leq y\mid Z_{i}=z\right)$
+$\leq$ $1-\mathbb{P}\left(Y_{is}>y,\ Z_{is}^{{}^{\prime}}\beta_{0}+X_{is}^{{}^{\prime}}\gamma_{0}\leq c\mid Z_{i}=z\right).$
+The key difference of the above identification strategy in (LABEL:eq:Bounds_z) relative to the corresponding identifying restrictions in previous sections lies in that the “middle term” in (LABEL:eq:Bounds_z) is no longer the conditional CDF of $\alpha_{i}+\epsilon_{it}$, but a conditional distribution of $\mathbb{P}\left(G\left(c,\,\alpha_{i},\,\epsilon_{is}\right)\leq y\mid Z_{i}=z\right)$, with the latter representation not dependent on scalar-additivity of fixed effects and time-varying errors.
+We summarize the identifying restriction above by the following proposition:
+Proposition 5 (Identified Set).
+Under Assumptions 1 and 3, $\theta_{0}\in\Theta_{I,gen}$, where the identified set $\Theta_{I,gen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$\max_{t=1,...,T}\mathbb{P}\left(Y_{it}\leq y,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c\mid Z_{i}=z\right)\leq 1-\max_{s=1,...,T}\mathbb{P}\left(Y_{is}>y,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\leq c\mid Z_{i}=z\right),$ (19)
+for any $c\in\mathcal{R},$ $y\in\mathcal{Y}$, and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
+Note that in the binary choice setting of Section 2, it suffices to set $y=0$ in (19), which then coincides with the identifying results in Proposition 1.
+More generally, we show that our identification results in Proposition 5 can be adapted to the ordered response model in Section 3.2, the multinomial choice model in Section 3.3, and the censored outcome model in Section 3.4.
+The results in Proposition 5 generally hold regardless of whether the dependent variable and the endogenous covariate are discrete or continuous.
+The next proposition shows that additional discreteness in either the dependent variable or endogenous covariates can further simplify and reduce the number of the identifying conditions in (19).
+Proposition 6 (Discreteness).
+When $X_{it}\in\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ for any $t$, then $\Theta_{I,gen}=\Theta_{I,gen}^{disc_{x}},$ where $\Theta_{I,gen}^{disc_{x}}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}$ that satisfy condition (19) for any $c\in\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K,t=1,...,T\right\}$.
+Moreover, when $Y_{it}\in\left\{\overline{y}_{1},...,\overline{y}_{K}\right\}$ with $\overline{y}_{j}\leq\overline{y}_{j+1}$ for any $t$, then $\Theta_{I,gen}=\Theta_{I,gen}^{disc_{y}},$ where $\Theta_{I,gen}^{disc_{y}}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}$ that satisfy condition (19) for any $y\in\left\{\overline{y}_{1},...,\overline{y}_{K-1}\right\}$.
+Proposition 6 shows that for any discrete choice models with discrete endogenous variables, such as dynamic binary, ordered, and multinomial choice models, the identified set $\Theta_{I,gen}$ is characterized by a finite number of moment inequalities.
+Sections 3.2-3.4 establish general identification results for panel models with various types of dependent variables and endogeneity, and also explore both the static model without endogeneity and the dynamic model with lagged dependent variables.
+3.2 Ordered Response Model
+Consider that the outcome variable $Y_{it}$ takes $J$ ordered values: $Y_{it}\in\{y_{1},..,y_{J}\}$ with $y_{j}\leq y_{j+1}$.
+Examples of such ordered outcomes include various income categories, health outcomes, or levels of educational attainment.
+We study the following panel ordered choice model:
+$Y_{it}^{*}$ $=W_{it}^{\prime}\theta_{0}+\alpha_{i}+\epsilon_{it},$ (20)
+$Y_{it}$ $=\sum_{j=1}^{J}y_{j}\mathbbm{1}\{b_{j}<Y_{it}^{*}\leq b_{j+1}\},$
+where $Y_{it}^{*}$ denotes the latent dependent variable, and $Y_{it}$ denotes the ordered outcome which takes value $y_{j}$ when $Y_{it}^{*}\in(b_{j},b_{j+1}]$.
+The threshold parameters satisfy $b_{1}=-\infty,b_{J+1}=+\infty$, and the remaining threshold parameters $b_{j}$ (where $b_{j}\leq b_{j+1}$) can be either known or unknown for $12\leq j\leq J-$ 1.
 The binary choice model in (1) is nested with $J=2$ and $b_{2}=0$.
-Similar to model (1), $X_{it}$ is the potentially endogenous covariate that can be arbitrarily dependent on $(\alpha_{i},\epsilon_{it})$, such as the lagged dependent variable $Y_{i,t-1}$, while $Z_{it}$ is the exogenous covariate that satisfies Assumption 1:
-$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
-We now show how our key identification strategy based on partial stationarity can also be exploit to partially identify $\theta_{0}$.
-Similar to Section 2, we still seek to establish both lower and upper bounds for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, where $v_{it}:=\epsilon_{it}+\alpha_{i}$.
-The distinction is that in the ordered choice setting we can obtain $J$ different (upper and lower) bounds, which can then be aggregated over to form tighter bounds as shown below.
-1. Lower bound. By the monotonicity of a distribution, we first derive a lower bound for the distribution $v_{t}\mid w_{st}$ conditional on all covariates $W_{ist}=w_{st}$, given as
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $=\Pr(v_{t}\leq a\mid w_{st})$
-$\geq\Pr(v_{t}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
-where the above lower bound holds for any choice $j\leq J$.
-Since the interval $(-\infty,b_{j+1}-w_{t}^{\prime}\theta_{0}]$ can be expressed as the union of multiple intervals: $(-\infty,b_{j+1}-w_{t}^{\prime}\theta_{0}]=\bigcup\limits_{k=1}^{j}(b_{k}-w_{t}^{\prime}\theta_{0},b_{k+1}-w_{t}^{\prime}\theta_{0}]$, the above lower bound can be also written as
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $\geq\sum_{k=1}^{j}\Pr(b_{k}-w_{t}^{\prime}\theta_{0}<v_{t}\leq b_{k+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}$
-$=\sum_{k=1}^{j}\Pr(Y_{it}=y_{k}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
-for any $j\leq J$.
-Fixing the constant $a$, there exists a maximum choice $j^{*}_{a}$, defined as $j^{*}_{a}:=\max\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}$.
-Then, the largest lower bound is the sum of choice probabilities up to the maximum choice $j^{*}_{a}$:
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\sum_{k=1}^{j^{*}_{a}}\Pr(Y_{it}=y_{k}\mid w_{st})\mathbbm{1}\{b_{j^{*}_{a}+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$
-An equivalent expression for this lower bound is given as:
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\},$
-as the indicator function is zero $\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}=0$ for any choice $j>j^{*}_{a}$.
-2. Upper bound. Similarly, we can derive the upper bound for $F_{v_{t}\mid w_{st}}(a\mid w_{st})$ using the sum of conditional choice probabilities, given as follows:
-$F_{v_{t}\mid W_{st}}(a\mid w_{st})$ $=1-\Pr(v_{t}>a\mid w_{st})$
-$\leq 1-\sum_{j=1}^{J}\Pr(b_{j}-w_{t}^{\prime}\theta_{0}<v_{t}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w_{st})\mathbbm{1}\{b_{j}-w_{t}^{\prime}\theta_{0}\geq a\}$
-$\leq 1-\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{t}^{\prime}\theta_{0}\geq a\}.$
-Then, the upper bound for the distribution $F_{v_{t}\mid z_{st}}(a\mid z_{st})$ is derived as follows by taking expectation over covariate $X_{ist}$ given $z_{st}$:
-$F_{v_{t}\mid z_{st}}(a\mid z_{st})\leq 1-\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j}-z_{t}^{\prime}\beta_{0}-X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st}).$
-Given the established lower and bounds on the distribution of $v_{t}\mid z_{st}$, an identified set for $\theta_{0}$ is characterized in the following proposition.
+For this ordered choice model, we observe the conditional probability of each choice $j$:
+$\mathbb{P}\left(Y_{it}\leq y_{j}\mid w\right)=\mathbb{P}\left(\alpha_{i}+\epsilon_{it}\leq b_{j+1}-w_{t}^{\prime}\theta_{0}\mid w\right),$
+and this choice probability is monotone with respect to $b_{j+1}-w_{t}^{\prime}\theta_{0}$.
+Compared to the binary choice model, here we can exploit variations in the conditional probability across all possible choices to provide more informative results for $\theta_{0}$.
+Following the same identification strategy in Sections 2 and 3, we can bound the conditional probability of $\mathbb{P}\left(\alpha_{i}+\epsilon_{it}\leq c\mid w\right)$ below as follows: for any constant $c$,
+$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)\geq\mathbb{P}\left(Y_{it}\leq y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right).$ (21)
+Since the above inequality holds for any choice $j$, we can take the largest lower bound over all possible choices that satisfy $b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c$.
+We define $\overline{j}_{c}$ as $\overline{j}_{c}:=\max\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\}$, and condition (21) leads to:
+$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$ $\geq\sum_{k=1}^{\overline{j}_{c}}\mathbb{P}\left(Y_{it}=y_{k},\ b_{\overline{j}_{c}+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right)$
+$=\sum_{j=1}^{J}\mathbb{P}\left(Y_{it}=y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right).$
+where the last inequality holds since $\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\}=0$ for any choice $j>\overline{j}_{c}$.
+Furthermore, the final expression offers the advantage of avoiding the need to search for the maximum choice $\overline{j}_{c}$ for each constant $c$.
+Similarly, we can derive an upper bound for the probability $\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$:
+$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)\leq 1-\mathbb{P}\left(Y_{it}>y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\geq c\mid w\right).$
+Define $\underline{j}_{c}$ as $\underline{j}_{c}:=\min\{j:b_{j+1}-w_{t}^{\prime}\theta_{0}\geq c\}$, then we have
+$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$ $\leq 1-\mathbb{P}\left(Y_{it}>b_{\underline{j}_{c}},\ b_{\underline{j}_{c}+1}-w_{t}^{\prime}\theta_{0}\geq c\mid w\right)$
+$=1-\mathbb{P}\left(Y_{it}\geq b_{\underline{j}_{c}+1},\ b_{\underline{j}_{c}+1}-w_{t}^{\prime}\theta_{0}\geq c\mid w\right).$
+Furthermore, the above inequality can be equivalently written as
+$\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)\leq 1-\sum_{j=1}^{J}\mathbb{P}\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq c\mid z\right),$
+which holds since for $\mathbf{\mathbbm{1}}\{b_{j+1}-w_{t}^{\prime}\theta_{0}<c\}=0$ for any choice $j<\underline{j}_{c}$.
+Given the established lower and upper bounds on the conditional probability $\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid w)$, we can derive the corresponding bounds for $\mathbb{P}(\alpha_{i}+\epsilon_{it}\leq c\mid z)$ by taking expectation over the endogenous covariate $X$.
+Then, the identifying condition for $\theta_{0}$ is characterized by the restriction that the bounds over different periods must have nonempty intersections, as presented in the following proposition.
 Proposition 7.
-Under Assumption 1, the identified set $\Theta_{I,2}$ for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following conditions:
-$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st}),$ (9)
-for any $a\in\mathcal{R}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Proposition 4 characterizes the identified set for $\theta_{0}$ for the general ordered response model with potential endogeneity.
-This result allows for both static and dynamic models, as discussed in Section 3.2.
-In comparison to the binary choice model, Proposition 4 aggregates information from all $J$ ordered responses across different time periods to identify $\theta_{0}$.
-When there are only two choices ($J=2$), this result simplifies to the identifying condition (7) in Proposition 1, with $y_{1}=0,y_{2}=1$ and changing $a$ to $-a$.
-Again, the result in Proposition 4 accommodates both discrete and continuous endogenous covariates.
-When the endogenous $X_{it}$ only takes a finite number of values, such as lagged dependent variable, we can reduce the number of restrictions in Proposition 4 to a finite number, as shown in the next corollary.
+Under Assumptions 1, $\theta_{0}\in\Theta_{I,order}$, where the identified set $\Theta_{I,order}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$\max_{t=1,...,T}\sum_{j=1}^{J}\mathbb{P}\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq c\mid z\right)\leq\\ 1-\max_{s=1,...,T}\sum_{j=1}^{J}\mathbb{P}\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq c\mid z\right),$ (22)
+for any $c\in\mathcal{R}$ and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
+Proposition 7 characterizes the identified set of $\theta_{0}$ for the general ordered response model with endogeneity.
+In contrast to the binary choice model discussed in Section 2, Proposition 7 exploits information from all possible choices across different time periods to identify $\theta_{0}$.
+This result accommodates both static and dynamic models, and we derive the simplified identifying results for these two types of models as follows.
+Static model: consider that there is no endogeneity and the full stationarity assumption holds conditional on all covariates $W_{i}$:
+$\epsilon_{is}\mid W_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{i},\alpha_{i}.$
+The identifying restriction in Proposition 7 is given as
+$\max_{t=1,...,T}\sum_{j=1}^{J}\mathbb{P}\left(Y_{it}=y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right)\leq 1-\max_{s=1,...,T}\sum_{j=1}^{J}\mathbb{P}(Y_{is}=y_{j},\ b_{j}-w_{s}^{\prime}\theta_{0}\geq c\mid w).$
+The above condition is informative only if there exists $j_{1},j_{2}$ such that $b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\leq c\leq b_{j_{1}}-w_{s}^{\prime}\theta_{0}$, leading to
+$\max_{t=1,...,T}\sum_{j=1}^{j_{2}}\mathbb{P}\left(Y_{it}=y_{j}\mid w_{i}\right)$ $\leq 1-\max_{s=1,...,T}\sum_{j=j_{1}}^{J}\mathbb{P}(Y_{is}=y_{j}\mid w)$
+$=\max_{s=1,...,T}\sum_{j=1}^{j_{1}-1}\mathbb{P}(Y_{is}=y_{j}\mid w).$
+The following proposition presents the identification results for the static ordered choice model by changing $j_{1}-1$ with $j_{1}$.
 Corollary 1.
-Under Assumption 1 and $X_{it}$ only takes finite number of values $X_{it}\in\{a_{1},...,a_{K}\}$, the identified set $\Theta_{I,2}$ is the set of parameters $\theta$ that satisfy condition (9) for any $a\in\mathcal{Q}_{2}:=\{b_{j}-z_{s}^{\prime}\beta-a_{1}^{\prime}\gamma,...,b_{j}-z_{s}^{\prime}\beta-a_{K}^{\prime}\gamma,b_{j}-z_{t}^{\prime}\beta-a_{1}^{\prime}\gamma,...,b_{j}-z_{t}^{\prime}\beta-a_{K}^{\prime}\gamma\}_{j=2}^{J}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Similar to Corollary 1, Corollary 2 reduces the number of conditional moment inequalities to $2K\times(J-1)$ numbers for any pair of two periods.
-For the dynamic model with one lagged dependent variable $X_{it}=Y_{i,t-1}\in\{y_{1},...,y_{J}\}$, then there will be $2J(J-1)$ number of restrictions.
-3.2 Applications
-In this section, we apply the results in Proposition 4 to explore two panel ordered choice model.
-The first one examines a static model where all covariates are exogenous, while the other studies a dynamic model with one lagged dependent variable.
-To the best of our knowledge, these two scenarios have not been explored in the literature within the framework of the stationarity assumption.
-Static model: we consider a static model where Assumption 1 holds conditional on all covariates $W_{ist}$. The identifying restriction in Proposition 4 is given as
-$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{s}^{\prime}\theta_{0}\geq a\}\\ \geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$
-We can further simplify the above condition by getting rid of the parameter $a$ and transform it into finite number of conditional moment inequalities.
-The following lemma presents the identified set for $\Theta_{I,2}$ for this static model.
-Lemma 1.
-Assume that $\epsilon_{is}\mid(W_{ist},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{ist},\alpha_{i})$, the identified set $\Theta_{I,2}$ for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following conditions:
-$b_{j_{1}+1}-w_{s}^{\prime}\theta\geq b_{j_{2}+1}-w_{t}^{\prime}\theta\Longrightarrow\sum_{j=1}^{j_{1}}\Pr(Y_{is}=y_{j}\mid w_{st})\geq\sum_{j=1}^{j_{2}}\Pr(Y_{it}=y_{j}\mid w_{st}),$
-for any $11\leq j_{1},j_{2}\leq J-$ 1, $s,t\leq T$, and $w_{st}\in\mathcal{R}^{d}\times\mathcal{R}^{d}$.
-In the static model, variation in conditional choice probability only comes from changes in the covariate $W_{it}$.
-Thus, we can directly establish the relationship between the variation in choice probability and the changes in covariates over time, eliminating the parameter $a$.
-Different from the binary choice model, the results in Lemma 1 also exploit intertemporal variation in the sum of multiple choices rather than investigating a single choice.
-Moreover, we can utilize variations in the sum of different choices across various periods to identify $\theta_{0}$.
-Besides the static model, the results in Proposition 4 can also be applied to study a dynamic ordered response model, where people’s choice at the current period ($t$) can depend on their choice in the last period ($t-1$).
-Dynamic model: consider the following dynamic ordered response model with one lagged dependent variable:
+Assume that $\epsilon_{is}\mid(W_{i},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{i},\alpha_{i})$, then $\Theta_{I,order}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$b_{j_{1}+1}-w_{s}^{\prime}\theta\geq b_{j_{2}+1}-w_{t}^{\prime}\theta\Longrightarrow\max_{s=1,...,T}\sum_{j=1}^{j_{1}}\mathbb{P}(Y_{is}=y_{j}\mid w)\geq\max_{t=1,...,T}\sum_{j=1}^{j_{2}}\mathbb{P}(Y_{it}=y_{j}\mid w),$
+for any $11\leq j_{1},j_{2}\leq J-$ 1, and any realization $w=(w_{1},...,w_{T})$ in the support of $W_{i}$.
+The results in Corollary 1 are analogous to the maximum-score type result in Manski (1987), with the distinction being that we can exploit variations in the sum of multiple choices rather than investigating a single choice to identify $\theta_{0}$.
+Furthermore, with multiple choices, we can utilize variations in the sum of different choices across various periods for identification.
+Besides the static model, Proposition 7 can also accommodate dynamic ordered choice models with lagged dependent variable.
+Dynamic model: consider the following dynamic ordered choice model with one lagged dependent variable:
 $Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+Y_{it-1}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
 $Y_{it}$ $=\sum_{j=1}^{J}y_{j}\mathbbm{1}\{b_{j}<Y_{it}^{*}\leq b_{j+1}\}.$
-In this example, the endogenous covariate is the lagged dependent variable $X_{it}=Y_{it-1}\in\{y_{1},...,y_{J}\}$. The identifying restriction in Proposition 4 is presented as follows:
-$1-\sum_{j=1}^{J}\Pr\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\geq a\mid z_{st}\right)\\ \geq\sum_{j=1}^{J}\Pr\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq a\mid z_{st}\right),$
-for any $a\in\{b_{j}-z_{s}^{\prime}\beta-y_{1}\gamma,...,b_{j}-z_{s}^{\prime}\beta-y_{J}\gamma,b_{j}-z_{t}^{\prime}\beta-y_{1}\gamma,...,b_{j}-z_{t}^{\prime}\beta-y_{J}\gamma\}_{j=2}^{J}$.
-For the dynamic model, the identified set $\Theta_{I,2}$ is characterized by $2J(J-1)$ number of conditional inequalities for any pair of two periods $(s,t)$.
-This approach can also allow for dynamic models with more than one lagged dependent variable, such as $X_{it}=(Y_{i,t-1},Y_{i,t-2})$.
-Furthermore, it can accommodate dynamic models that allow for heterogeneous effects from different choices in the last period, as given below:
-$Y_{it}^{*}=Z_{it}^{\prime}\beta_{0}+\sum_{j=2}^{J}\mathbf{\mathbbm{1}}\{Y_{it-1}=j\}\gamma_{0,j}+\alpha_{i}+\epsilon_{it}.$
-3.3 Point Identification
-This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$.
-Similar to Section 2.3, we still focus on the scenario where the endogenous covariate $X_{it}$ only takes finite number of values $X_{it}\in\mathcal{X}=\{a_{1},a_{2},...,a_{K}\}$, and we illustrate the concept using two periods ($T=2$).
-To identify $\beta_{0}$, we define the following two sets:
 […]
-$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x\mid z_{12})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Big{\}};$
-[…]
-$1-\sum_{j=1}^{k-1}\Pr(Y_{i1}=y_{j},X_{i1}=x\mid z_{12})<\sum_{j=k}^{J}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Big{\}}.$
-Let $\mathcal{Z}_{order}:=\mathcal{Z}_{1,order}\cup\mathcal{Z}_{2,order}$. Let $\Delta Z_{i}=Z_{i2}-Z_{i1}$ and $\Delta\mathcal{Z}_{order}$ be defined as
-$\Delta\mathcal{Z}_{order}:=\Big{\{}\Delta z:=z_{2}-z_{1}\mid(z_{1},z_{2})\in\mathcal{Z}_{order}\Big{\}}.$
-Assumption 3 (Support Condition).
-(1) $\Delta\mathcal{Z}_{order}$ is not contained in any proper linear subspace of $\mathcal{R}^{d_{z}}$; (2) for any $\Delta z\in\Delta\mathcal{Z}_{order}$, there exists at least one least one element $\Delta z^{j^{*}}$ with $\beta_{0}^{j*}\neq 0$, and the conditional support of $\Delta z^{j^{*}}$ is $\mathcal{R}$ given $\Delta z\setminus\Delta z^{j^{*}}$, where $\Delta z\setminus\Delta z^{j^{*}}$ denote the remaining components of $\Delta z$ besides $\Delta z^{j^{*}}$.
-Proposition 5.
-Under Assumptions 1 &\&& 3, the parameter $\beta_{0}$ is point identified up to scale.
-Similar to Proposition 2, the strategy is to first identify the sign of $\Delta z^{\prime}\beta_{0}$ from certain variation in the choice probability, as defined in the two sets $(\mathcal{Z}_{1,order},\mathcal{Z}_{2,order})$.
-Given the sign of $\Delta z^{\prime}\beta_{0}$ is identified, then the parameter $\beta_{0}$ is point identified up to scale under the conventional large support condition on $\Delta z$.
-For the static model (with $\gamma_{0}=0$), all coefficients $\beta_{0}$ are point identified under the support condition.
-To identify the sign of $\gamma_{0}$, we define the following two sets: for $j\leq d_{x}$,
-[…]
-$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}\mid z_{12})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Big{\}};$
-[…]
-$1-\sum_{j=1}^{k-1}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}\mid z_{12})<\sum_{j=k}^{J}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Big{\}}.$
-Proposition 6.
-Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}\neq\emptyset$ or $\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
-For the dynamic model with one lagged dependent variable $X_{it}=Y_{i,t-1}$, then the joint choice probabilities defined in the two sets $\mathcal{Z}^{j}_{3,order},\mathcal{Z}^{j}_{4,order}$ can be simplified as taking different values for the lagged dependent variable: $Y_{i0}=y_{j_{1}},Y_{i1}=y_{j_{2}}$ with $y_{j_{1}}\neq y_{j_{2}}$.
-The identifying restriction in Proposition 4, in general, can bound the value of $(y_{j}-y_{k})\gamma_{0}$ using the variation in $Z_{it}$.
-Proposition 6 ensures that its bound excludes zero for some $(y_{j},y_{k})$, thereby identifying the sign of $\gamma_{0}$.
-When there are more than two periods ($T>2$), then point identification of $\beta_{0}$ and the sign of $\gamma_{0}$ is achieved when there exists two periods such that the assumptions are satisfied.
-4 Multinomial Choice Model
-4.1 Model and Identification
+3.3 Multinomial Choice Model
 This section applies our key identification strategy to panel multinomial choice model with endogeneity.
-Specifically, consider a set of choice alternatives $\mathcal{J}=\{0,1,...,J\}$.
-Let $u_{ijt}$ denote the utility for individual $i$ for choice $j$ at time $t$, which depends on three components: observed covariate $W_{ijt}=(Z_{ijt}^{\prime},X_{ijt}^{\prime})^{\prime}$, unobserved fixed effects $\alpha_{ij}$, and unobserved time-varying preference shock $\epsilon_{ijt}$.
-The utility of choice 0 (outside option) is normalized to zero: $u_{i0t}=0$.
+Specifically, consider a set of unordered choice alternatives $\mathcal{J}=\{0,1,...,J\}$.
+Let $u_{ijt}$ denote the latent utility for individual $i$ of selecting choice $j$ at time $t$, which depends on the three components: observed covariate $W_{ijt}=(Z_{ijt}^{\prime},X_{ijt}^{\prime})^{\prime}$, unobserved fixed effects $\alpha_{ij}$, and unobserved time-varying preference shock $\epsilon_{ijt}$.
 Let $Y_{it}\in\mathcal{J}$ denote individual $i$’s choice at time $t$.
 We study the following panel multinomial choice model:
-$u_{ijt}$ $=Z_{ijt}^{\prime}\beta_{0}+X_{ijt}^{\prime}\gamma_{0}+\alpha_{ij}+\epsilon_{ijt},$
+$u_{ijt}$ $=W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt},$
 $Y_{it}$ $=\arg\max_{j\in\mathcal{J}}u_{ijt},$
-where $X_{ijt}$ denotes the potentially endogenous covariate, such as the lagged dependent variable or endogenously determined prices, and $Z_{ijt}$ denotes the exogenous covariates that satisfies the partial stationarity condition (Assumption 1)
-$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
-with $Z_{it}=\{Z_{ijt}\}_{j\in\mathcal{J}}$,$\alpha_{i}=\{\alpha_{ij}\}_{j\in\mathcal{J}}$ and $\epsilon_{it}=\{\epsilon_{ijt}\}_{j\in\mathcal{J}}$ defined to collect terms across all $J$ choice alternatives.
-The identification of $\theta_{0}$ has been studied under the standard full stationarity condition given all covariates $W_{ist}$ in different models, including Pakes and Porter (2022), Shi, Shum, and Song (2018), Gao and Li (2020), and Wang (2022).
-The main focus of this paper is to derive identification for $\theta_{0}$, while allowing for the endogenous covariate $X_{it}$.
-As discussed in Section 4.2, our approach also accommodates the standard stationarity assumption, and the results are consistent with those in Pakes and Porter (2022).
-Let $v_{ijt}=\alpha_{ij}+\epsilon_{ijt}$ and $v_{it}=\{v_{ijt}\}_{j\in\mathcal{J}}$.
-The identification strategy is still to establish bounds on the distribution of $v_{it}\mid Z_{ist}$ and derive identifying restrictions on $\theta_{0}$ based on partial stationarity.
-Here, the distinction is that the error term $v_{it}$ is multi-dimensional instead of one-dimensional in binary or ordered choice models.
-Moreover, consumers’ choice involves the comparison between different choices.
-The observed choice probability of selecting $j$ at period $t$ is given as
-$\Pr(Y_{it}=j\mid w_{st})=\Pr\left(w_{jt}^{\prime}\theta_{0}+v_{ijt}\geq w_{kt}^{\prime}\theta_{0}+v_{ikt}\quad\forall k\in\mathcal{J}\mid w_{st}\right).$
-The above probability restricts the distribution $\Pr(v_{ikt}-v_{ijt}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \forall k\neq j\mid w_{st})$.
-From observed choice probability, we can derive bounds for the distribution $\Pr(v_{ikt}-v_{ijt}\leq a_{jk}\ \forall k\neq j\mid z_{st})$ as a function of $\theta_{0}$ and observed variable $(X_{it},Z_{it},Y_{it})$.
-Therefore, the identifying condition for $\theta_{0}$ is characterized by the restriction that the bounds over different periods must intersect.
-More generally, we observe the conditional probability $\Pr(Y_{it}\in K\mid w_{st})$ for any set $K\subset\mathcal{J}$ and can bound the probability $\Pr(v_{ikt}-v_{ijt}\leq a_{jk}\ \exists j\in K,\forall k\in K^{c}\mid z_{st})$, where $K^{c}:=\mathcal{J}\setminus K$.
-In the case where $K$ is a singleton ($K=\{j\}$), it then becomes the conditional probability of selecting a single choice $j$.
-The following proposition characterizes the identification results for $\theta_{0}$.
+and impose the same partial stationarity assumption:
+$\epsilon_{is}\mid Z_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{i},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
+with $Z_{it}:=\{Z_{ijt}\}_{j\in\mathcal{J}}$,$\alpha_{i}:=\{\alpha_{ij}\}_{j\in\mathcal{J}}$ and $\epsilon_{it}:=\{\epsilon_{ijt}\}_{j\in\mathcal{J}}$ defined to collect terms across all $J$ choice alternatives, and $Z_{i}:=(Z_{i1},Z_{i2},...,Z_{iT})$.
+Although the $J$ choices are unordered and not directly comparable, we look at the indicator variable $Y_{it}^{j}:=\mathbf{\mathbbm{1}}\{Y_{it}=j\}$ of choosing option $j$, which maintains a similar monotone structure with Assumption 3:
+$Y_{it}^{j}=1\Longleftrightarrow W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt}\geq W_{ikt}^{\prime}\theta_{0}+\alpha_{ik}+\epsilon_{ikt},\ \forall k\in\mathcal{J},$
+and the new variable $Y_{it}^{j}$ is monotone in $W_{ijt}^{\prime}\theta_{0}-W_{ikt}^{\prime}\theta_{0}$ given $(\alpha_{i},\epsilon_{it})$.
+More generally, for any subset $K\subset\mathcal{J}$, the indicator variable $Y_{it}^{K}:=\mathbf{\mathbbm{1}}\{Y_{it}\in K\}$ represents individual $i$’s choice belonging to the subset $K$, given as follows:
+$Y_{it}^{K}=1\Longleftrightarrow W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt}\geq W_{ikt}^{\prime}\theta_{0}+\alpha_{ik}+\epsilon_{ikt},\ \exists j\in K,\forall k\in\mathcal{J}\setminus K,$
+and the variable $Y_{it}^{K}$ is monotone in $W_{ijt}^{\prime}\theta_{0}-W_{ikt}^{\prime}\theta_{0}$ for any $j\in K$ and $k\in\mathcal{J}\setminus K$.
+Following Proposition 5, the identification results for panel multinomial choice models are presented in the following proposition.
 Proposition 8.
-Under Assumption 1, an identified set $\Theta_{I,3}$ for $\theta_{0}$ is the set of parameters $\theta=(\beta,\gamma)$ that satisfy the following condition:
-$1-\Pr(Y_{is}\in K^{c},(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st})\\ \geq\Pr(Y_{it}\in K,(z_{jt}-z_{kt})^{\prime}\beta+(X_{ijt}-X_{ikt})^{\prime}\gamma\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st}),$ (10)
-for any subset $K\subset\mathcal{J}$, any $a_{jk}\in\mathcal{R}$ with $j\in K$ and $k\in K^{c}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Corollary 3.
-Under Assumption 1, and when $X_{ijt}-X_{ikt}$ only takes finite number of values $X_{ijt}-X_{ikt}\in\mathcal{X}_{jk}:=\{x_{jk,1},...,x_{jk,M}\}$, then the identified set $\Theta_{I,3}$ is the set of parameters $\theta$ that satisfy condition (10) for any subset $K\subset\mathcal{J}$, any $a_{jk}\in\{(z_{js}-z_{ks})^{\prime}\beta+x_{jk}^{\prime}\gamma,(z_{jt}-z_{kt})^{\prime}\beta+x_{jk}^{\prime}\gamma\}_{x_{jk}\in\mathcal{X}_{jk}}$ with $j\in K$ and $k\in K^{c}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Proposition 7 provides general identification results for multinomial choice models with discrete or continuous endogenous covariates.
-Corollary 3 further reduces the number of restrictions in Proposition 7 to a finite number when the difference in the endogenous covariate between choices, $X_{ijt}-X_{ikt}$, only takes finite number of values.
-The results accommodate both static models and dynamic models with any number of lagged dependent variables.
-Point identification can be achieved under conditions similar to those presented in Section 2.3 and 3.3, so further analysis is omitted here.
-4.2 Applications
-We examine the two applications of Proposition 7 for the panel multinomial choice model.
-The first one studies the static model, where all covariates $W_{it}$ are exogenous ($\epsilon_{is}\mid W_{ist},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{ist},\alpha_{i}$), which is the same setting as Pakes and Porter (2022).
-In this scenario, we show that Proposition 7 boils down to the identifying results in Pakes and Porter (2022).
-Another one explores the dynamic multinomial choice model, incorporating one lagged dependent variable $Y_{i,t-1}$.
-Static model: Pakes and Porter (2022) characterizes the sharp identified set for $\theta_{0}$ under the stationarity assumption given all covariates: $\epsilon_{is}\mid W_{is},W_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{is},W_{it},\alpha_{i}$.
-Now, we show that the identifying restriction in (10) is consistent with the results in Pakes and Porter (2022).
-Under full stationarity (given all covariates $W_{it}$), condition (10) becomes
-$1-\Pr(Y_{is}\in K^{c},(w_{js}-w_{ks})^{\prime}\theta_{0}\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid w_{st})\\ \geq\Pr(Y_{it}\in K,(w_{jt}-w_{kt})^{\prime}\theta_{0}\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid w_{st}).$
-The above equation is only informative when $a_{jk}\in[(w_{jt}-w_{kt})^{\prime}\theta_{0},(w_{js}-w_{ks})^{\prime}\theta_{0}]$ for any $j\in K,k\in K^{c}$; otherwise either the upper bound becomes one or the lower bound becomes zero in the above restriction.
-When $a\in[(w_{jt}-w_{kt})^{\prime}\theta_{0},(w_{js}-w_{ks})^{\prime}\theta_{0}]$, the above restriction is equivalent to the following condition: for any $K\subset\mathcal{J}$,
-$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\geq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in K^{c}\\ \Longrightarrow 1-\Pr(Y_{is}\in K^{c}\mid w_{st})=\Pr(Y_{is}\in K\mid w_{st})\geq\Pr(Y_{it}\in K\mid w_{st}),$
-which aligns with the identification result in Pakes and Porter (2022) (Proposition 1, P. 12).
-Dynamic model: Proposition 7 is also applicable to the following dynamic panel multinomial choice model:
+Under Assumption 1, $\theta_{0}\in\Theta_{I,mul}$, where the identified set $\Theta_{I,mul}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$\max_{t=1,...,T}\mathbb{P}(Y_{it}^{K}=0,(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\geq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z)\\ \leq 1-\max_{s=1,...,T}\mathbb{P}\left(Y_{is}^{K}=1,(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\leq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right),$ (23)
+for any subset $K\subset\mathcal{J}$, any $c_{jk}\in\mathcal{R}$, any $j\in K$ and $k\in\mathcal{J}\setminus K$, and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
+Proposition 8 provides general identification results for multinomial choice models with endogenous covariates.
+It accommodates static, dynamic multinomial choice models with any number of lagged dependent variables, as well as other types of endogeneity.
+For the dynamic model, similar to Proposition 2, the results can be simplified to hold for a finite number of values of $c_{jk}$, which is the collection of values the covariate index $(z_{jt}-z_{kt})^{\prime}\beta+(X_{ijt}-X_{ikt})^{\prime}\gamma$ can take.
+Reconciliation with Pakes and Porter (2022)
+Next, we show that our results specialize to those in Pakes and Porter (2022), who focuses on the static panel multinomial choice model without any endogeneity.
+Formally, Pakes and Porter (2022) characterizes the sharp identified set for $\theta_{0}$ under the full stationarity assumption given all covariates:
+$\epsilon_{is}\mid W_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{i},\alpha_{i}.$
+Under full stationarity, for any two periods $(s,t)$, our identifying condition in (23) is simplified as
+$\mathbb{P}\left(Y_{it}^{K}=0,(w_{jt}-w_{kt})^{\prime}\theta_{0}\geq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid w\right)\\ \leq 1-\mathbb{P}\left(Y_{is}^{K}=1,(w_{js}-w_{ks})^{\prime}\theta_{0}\leq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid w\right)$ (24)
+The above equation is only informative when $(w_{js}-w_{ks})^{\prime}\theta_{0}\leq c_{jk}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}$ for any $j\in K,k\in k\in\mathcal{J}\setminus K$; otherwise either the upper bound becomes one or the lower bound becomes zero so condition (24) holds for any $\theta$.
+There exists one value $c_{jk}$ satisfying the condition $(w_{js}-w_{ks})^{\prime}\theta_{0}\leq c_{jk}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}$ is equivalent to $(w_{js}-w_{ks})^{\prime}\theta_{0}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}$, generating the following inequality: for any $K\subset\mathcal{J}$,
+$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in\mathcal{J}\setminus K\\ \Longrightarrow\mathbb{P}\left(Y_{it}^{K}=0\mid w\right)\leq 1-\mathbb{P}\left(Y_{is}^{K}=1\mid w\right),$
+which becomes the same result in Pakes and Porter (2022) (Proposition 1, P. 12):
+$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in\mathcal{J}\setminus K\\ \Longrightarrow\mathbb{P}\left(Y_{is}\in K\mid w\right)\leq\mathbb{P}\left(Y_{it}\in K\mid w\right),$
+since $Y_{it}^{K}=1$ is equivalent to $Y_{it}\in K$ by the definition.
+Beyond the static model, Proposition 8 allows for any type of endogeneity including dynamic multinomial models with lagged dependent variable. For example, consider the following dynamic model:
 $u_{ijt}=Z_{ijt}^{\prime}\beta_{0}+\mathbbm{1}\{Y_{i,t-1}=j\}\gamma_{0}+\alpha_{ij}+\epsilon_{ijt}.$
-The above model allows individuals’ current utility at time $t$ to depend on their choices at the last period $t-1$.
-In this model, the endogenous variable $X_{ijt}$ is the lagged dependent variable $Y_{i,t-1}$, and the difference of the endogenous covariate between choices only takes three values: $X_{ijt}-X_{ikt}\in\mathcal{X}_{jk}:=\{1,-1,0\}$.
-Then the identified set for $\Theta_{I,3}$ is characterized by the condition in Corollary 2 with $a_{jk}\in\{(z_{js}-z_{ks})^{\prime}\beta+\gamma,(z_{js}-z_{ks})^{\prime}\beta-\gamma,(z_{js}-z_{ks})^{\prime}\beta,(z_{jt}-z_{kt})^{\prime}\beta+\gamma,(z_{jt}-z_{kt})^{\prime}\beta-\gamma,(z_{jt}-z_{kt})^{\prime}\beta\}$.
-5 Censored Dependent Variable Model
+where individual $i$’s utility at time $t$ can potentially depend on their choices in the previous period $t-1$.
+In this model, the endogenous variable $X_{ijt}$ is whether option $j$ is chosen in the previous period $\mathbbm{1}\{Y_{i,t-1}=j\}$.
+Then, the difference in the endogenous covariate between choices only takes three values: $X_{ijt}-X_{ikt}\in\{1,-1,0\}$, and the identified set for $\Theta_{I,mul}$ is characterized by the condition in (23) with $c_{jk}\in\{(z_{jt}-z_{kt})^{\prime}\beta+\gamma,(z_{jt}-z_{kt})^{\prime}\beta-\gamma,(z_{jt}-z_{kt})^{\prime}\beta\}_{t=1}^{T}$.
+3.4 Censored Outcome Model
 The previous sections primarily investigate discrete choice models, while our approach also applies to models with continuous dependent variables, including those with censored or interval outcomes.
-To illustrate, we focus on censored outcomes models below.
-Consider the following panel models with censored outcomes:
+To illustrate, we focus on the following panel model with censored outcomes:
 $Y^{*}_{it}$ $=Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
 $Y_{it}$ $=\max\{Y^{*}_{it},0\},$
-where $X_{it}$ represents the endogenous covariate, $Z_{it}$ stands for the exogenous covariate that satisfies Assumption 1, $Y^{*}_{it}$ denotes the latent outcome not observed in the data, and $Y_{it}$ represents the observed outcome, which is censored at zero.
+where $Y^{*}_{it}$ denotes the latent outcome which is not observed in the data, and $Y_{it}$ represents the observed outcome, censored at zero.
 The threshold for censoring can be replaced with other nonzero constants.
-The identification strategy is still to derive conditions for $\theta_{0}$ based on the stationarity assumption.
-Here the conditional distribution of $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ is not identified since we only observe the censoring outcome $Y_{it}$.
-While we can still bound the conditional distribution $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ using variation in the joint distribution $(Y_{it},Z_{it},X_{it})$.
-The following proposition presents the identification results of $\theta_{0}$ with censoring outcomes.
+The identification strategy is still to exploit the partial stationarity assumption and bound the conditional distribution of $\alpha_{i}+\epsilon_{it}\mid Z_{i}=z$.
+This censored outcome model imposes an additional structure between the outcome and the parametric index: when $Y_{it}>0$, we have $Y_{it}=Y^{*}_{it}$ and
+$\alpha_{i}+\epsilon_{it}\leq c\Longleftrightarrow Y_{it}-Z_{it}^{\prime}\beta_{0}-X_{it}^{\prime}\gamma_{0}\leq c.$
+This specific structure can be exploited to further tighten the identified set for $\theta_{0}$, and we provide the details of the identification strategy in Appendix A.7.
+The following proposition presents the identification results of $\theta_{0}$ with censored outcomes.
 Proposition 9.
-Under Assumption 1, an identified set for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following condition:
-$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a\mid z_{st}),$
-for any $a\in\mathcal{R}$, any $(s,t)\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Similar to discrete choice models studied in previous sections, Proposition 8 characterizes an identified set for $\theta_{0}$ by exploiting the variation in the joint distribution $(Y_{it},X_{it})\mid Z_{ist}$ over time and the variation in the exogenous covariates $Z_{ist}$.
-The bounds on the distribution $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ can be derived either from the probability $\Pr(0<Y_{it}\leq y\mid z_{st})$ or $\Pr(Y_{it}=0\mid z_{st})$, depending on the value of the covariate index $z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma$.
-This result still accommodates both static and dynamic models with censored outcomes.
-Static model: consider that the standard stationarity assumption holds, where $\epsilon_{it}\mid\alpha_{i},W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{ist}$.
+Under Assumption 1, $\theta_{0}\in\Theta_{I,cen}$, where the identified set $\Theta_{I,cen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$\max_{t=1,...,T}\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-c\mid z)\leq\max_{s=1,...,T}\left\{\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-c\mid z)+\mathbb{P}(Y_{is}=0\mid z)\right\},$
+for any $c\in\mathcal{R}$ and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
+Similar to discrete choice models studied in previous sections, Proposition 9 characterizes an identified set for $\theta_{0}$ by exploiting the variation in the joint distribution of $(Y_{it},X_{it})\mid Z_{i}$ over time and the variation in the exogenous covariates $Z_{i}$.
+The bounds on the distribution $\alpha_{i}+\epsilon_{it}\mid Z_{i}=z$ can be derived either from the probability $\mathbb{P}(0<Y_{it}\leq y\mid z)$ or $\mathbb{P}(Y_{it}=0\mid z)$, depending on the value of the covariate index $z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}$.
+This result still accommodates both static and dynamic models with censored outcomes, and we provide the simplified results for each model.
+Static model: consider that the full stationarity assumption holds, i.e., $\epsilon_{it}\mid\alpha_{i},W_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{i}$.
 Then, the identifying condition in Proposition 9 is given as
-$\Pr(0<Y_{is}\leq w_{s}^{\prime}\theta-a)+\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$ (11)
-The above restriction is informative only when $w_{t}^{\prime}\theta\geq a$, otherwise the lower bound becomes zero.
-We discuss two cases for the constant $a$.
-When $w_{s}^{\prime}\theta\leq a\leq w_{t}^{\prime}\theta$, then condition (11) becomes
-$\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$
-When $a$ satisfies $a\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$, condition (11) transforms into
-$\Pr(0<Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})+\Pr(Y_{is}=0\mid w_{st})=\Pr(Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})\\ \geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$
+$\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(0<Y_{is}\leq w_{s}^{\prime}\theta-c\mid w)+\mathbb{P}(Y_{is}=0\mid w).$ (25)
+The above restriction is informative only when $w_{t}^{\prime}\theta-c\geq 0$, otherwise the lower bound becomes zero.
+We discuss two cases for the constant $c$: $w_{s}^{\prime}\theta\leq c\leq w_{t}^{\prime}\theta$ and $c\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$.
+When $w_{s}^{\prime}\theta\leq c\leq w_{t}^{\prime}\theta$, then condition (25) becomes
+$\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(Y_{is}=0\mid w).$
+When $c$ satisfies $c\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$, condition (25) transforms into
+$\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(0<Y_{is}\leq w_{s}^{\prime}\theta-c\mid w)+\mathbb{P}(Y_{is}=0\mid w)=\mathbb{P}(Y_{is}\leq w_{s}^{\prime}\theta-c\mid w).$
 Since the above condition needs to hold for any $(s,t)$ and is symmetric in $(s,t)$, it becomes equalities after exchanging $s$ and $t$.
-The following lemma derives the identified set for $\theta_{0}$ using both conditional moment inequalities and equalities.
-Lemma 2.
-Assuming that $\epsilon_{it}\mid\alpha_{i},W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{ist}$, an identified set for $\theta_{0}$ is characterized by the set of parameters $\theta$ that satisfy the following conditions:
-$&\text{If }w_{s}^{\prime}\theta\leq a\leq w_{t}^{\prime}\theta\Longrightarrow\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}),\\ &\text{If }a\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}\Longrightarrow\Pr(Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})=\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}),\end{aligned}\right.$
-for any $s,t\leq T$ and any $w_{st}\in\mathcal{R}^{d}\times\mathcal{R}^{d}$.
+The following lemma summarizes the results for the static model.
+Corollary 2.
+Assuming that $\epsilon_{is}\mid\alpha_{i},W_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid\alpha_{i},W_{i}$, the identified set $\Theta_{I,cen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$&\text{If }w_{s}^{\prime}\theta\leq c\leq w_{t}^{\prime}\theta\Longrightarrow\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(Y_{is}=0\mid w);\\ &\text{If }c\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}\Longrightarrow\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)=\mathbb{P}(Y_{is}\leq w_{s}^{\prime}\theta-c\mid w),\end{aligned}\right.$
+for any $c\in\mathcal{R}$, any $(s,t)\leq T$, and any realization $w=(w_{1},...,w_{T})$ in the support of $W_{i}$.
 Dynamic model: Proposition 9 also accommodates the following dynamic model with the lagged outcome $Y_{i,t-1}$:
@@
 Consequently, the results in Proposition 9 need to be adjusted as the endogenous variable $X_{it}=Y^{*}_{i,t-1}$ is not observed.
-6 Generalization to Nonseparable Models
-The key idea underlying our identification strategy generalizes further beyond the models considered before.
-In this section, we show that our approach relies only on two conditions: partial stationarity and “index monotonicity.” Additional model specifications and restrictions, such as the scalar-additivity of fixed effects and time-varying errors, are not needed for our identification strategy.
-To illustrate, consider the following nonseparable semiparametric model:
-$Y_{it}$ $=G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right),$
-where $\alpha_{i}$ is the individual fixed effect of arbitrary dimension, $\epsilon_{it}$ is the time-varying error of arbitrary dimension, $W_{it}$ is a vector of observable covariates, $\theta_{0}\in\mathcal{R}^{d}$ is a conformable vector of parameter, and the function $G$ is allowed to be unknown, nonseparable but assumed to be weakly monotone in the the parametric index:
-Assumption 4 (Monotonicity).
-The mapping $\delta\longmapsto G\left(\delta,\alpha,\epsilon\right)$ is weakly increasing in $\delta\in\mathcal{R}$ for each realization of $\left(\alpha,\epsilon\right)$.
-Note that, by setting $\alpha_{i},\epsilon_{it}$ to be scalar-valued, and $G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\mathbf{\mathbbm{1}}\left\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\}$, we obtain the dynamic binary choice model in Section 2, where $G$ is by construction weakly increasing in $W_{it}^{{}^{\prime}}\theta_{0}$.
-As before, we decompose $W_{it}$, and correspondingly $\theta_{0}$, into two components, $W_{it}=\left(Z_{it}^{{}^{\prime}},X_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $\theta_{0}=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$so that
-$W_{it}^{{}^{\prime}}\theta_{0}=Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0},$
-where $Z_{it}$ denotes exogenous covariates while $X_{it}$ denotes endogenous covariates, with the precise definition of exogeneity encoded by the partial stationarity in Assumption 1.
-Now we show how partial stationarity can be exploited in conjunction with weak monotonicity (Assumption 4) to obtain identifying restrictions in the presence of endogeneity.
-Let $\mathcal{Y}$ denote the support of $Y_{it}$.
-For any $a\in\mathcal{R}$ and $y\in\mathcal{Y}$, conditional on $W_{ist}=w_{st}$, we can bound $\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$ below as
-$\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$ $\geq\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}\leq a\right\}\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$
-$\geq\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}\leq a\right\}\mathbf{\mathbbm{1}}\left\{G\left(w_{t}^{\prime}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\},$
-and above as
-$\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$
-[…]
-[…]
-=\ = $1-\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}>a\right\}\mathbf{\mathbbm{1}}\left\{G\left(w_{t}^{\prime}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)<y\right\}.$
-Then, the conditional probability $\mathbb{P}\left(G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\mid w_{st}\right)$ given $W_{ist}=w_{st}$ can be bounded as follows:
-[…]
-Taking conditional expectation of the above with respect to $X_{ist}$ given $Z_{ist}=z_{st}$ yields
-$\mathbb{P}\left(Y_{is}\geq y,W_{is}^{\prime}\theta_{0}\leq a\mid z_{st}\right)$
-[…]
-[…]
-The key difference of (6) relative to the corresponding identifying restrictions in previous sections lies in that the “middle term” in (6) is no longer the conditional CDF of $\alpha_{i}+\epsilon_{it}$, but a conditional probability about the event $G\left(a,\,\alpha_{i},\,\epsilon_{is}\right)\geq y$, with the latter representation not dependent on scalar-additivity of fixed effects and time-varying errors.
-We summarize the identifying restriction above by the following proposition:
-Proposition 9.
-Under Assumptions 1 and 4, the following condition holds:
-$1-\mathbb{P}\left(Y_{it}<y,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}>a\mid z_{st}\right)\geq\mathbb{P}\left(Y_{is}\geq y,z_{s}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\leq a\mid z_{st}\right),$ (12)
-for any $a\in\mathcal{R},$ $y\in\mathcal{Y}$, $s\neq t\leq T$, and $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
-Note that in the binary choice setting of Section 2, it suffices to set $y=1$ in (12), which then coincides with the identifying condition (7) in Proposition 1.
-It is also straightforward to show that our identification results in Lemma 9 can be adapted to the ordered response model in Section 3, the multinomial choice model in Section 4, and the censored outcome model in Section 5 without the scalar additive specifications.
 4 Simulation
 This section examines the finite sample performance of our identification approaches using Monte Carlo simulations.
-Since the literature has extensively studied binary choice models, we focus on the static and dynamic ordered choice models explored in Section 3 as examples to illustrate the approach.
+We focus on the static and dynamic ordered choice models explored in Section 3.2 as examples to illustrate the approach.
 We implement the kernel-based CLR inference approach proposed in the papers by Chernozhukov, Lee, and Rosen (2013) and Chen and Lee (2019), developed to construct confidence interval based on general conditional moment inequalities.
@@
 where $b_{2}=-1$ and $b_{3}=1$.
-We consider the covariates $Z_{it}$ to be exogenous and Lemma 1 characterizes the identified set for $\beta_{0}$ using the following conditional moment inequalities:
-$E[g(Z_{ist},Y_{ist};\beta_{0})\mid z_{st}]\geq 0,$
+With $Y_{i}:=(Y_{i1},Y_{i2})$ and $Z_{i}:=(Z_{i1},Z_{i2})$, Corollary 1 characterizes the identified set for $\beta_{0}$ using the following conditional moment inequalities: for $s\neq t\leq 2$,
+$E[g(Z_{i},Y_{i};\beta_{0})\mid z]\geq 0,$
 where
-$g(Z_{ist},Y_{ist};\beta_{0})=\left\{\begin{aligned} &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\}).\end{aligned}\right.$
+$g(Z_{i},Y_{i};\beta_{0})=\left\{\begin{aligned} &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\}).\end{aligned}\right.$
 The first element $\beta_{01}$ of the parameter $\beta_{0}$ is normalized to one, and we are interested in conducting inference for the parameter $\beta_{02}$ using the CLR approach.
@@
 $Y_{it}=1*(Y_{it}^{*}\leq b_{2})+2*(0\leq Y_{it}^{*}\leq b_{2})+3*(Y_{it}^{*}>b_{3}),$
-for $1\leq t\leq T$. The initial value $Y_{i0}\in\{1,2,3\}$ is generated independently of all variables, and follow the distribution $\Pr(Y_{i0}=1)=0.6,\Pr(Y_{i0}=2)=\Pr(Y_{i0}=3)=0.2$.
+for $1\leq t\leq T$.
+The initial value $Y_{i0}\in\{1,2,3\}$ is generated independently of all variables, and follow the distribution $\mathbb{P}(Y_{i0}=1)=0.6,\mathbb{P}(Y_{i0}=2)=\mathbb{P}(Y_{i0}=3)=0.2$.
 In this dynamic model, the covariates $Z_{i}:=(Z_{it})_{t=1}^{T}$ and the initial value $Y_{i0}$ are exogenous, while the lagged variable $Y_{i,t-1}$ is endogenous.
@@
     When $s\in\{2,3\}$,
-    $1-\sum_{j=2}^{3}\Pr\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}$$\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma\geq a\}$
-    $\geq\sum_{j=1}^{2}\Pr\left(Y_{is}=y_{j},b_{j+1}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\leq a\mid z,y_{0}\right),$
-    $1-\sum_{j=2}^{3}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq a\mid z,y_{0})$
-    $\geq\sum_{j=1}^{2}\Pr$ $\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}\{b_{j+1}-z_{1}^{\prime}\beta-y_{0}\gamma\leq a\},$
-    for any $a\in\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma,b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma\}_{j=2}^{T}$;
+    $1-\sum_{j=2}^{3}\mathbb{P}\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}$$\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma\geq c\}$
+    $\geq\sum_{j=1}^{2}\mathbb{P}\left(Y_{is}=y_{j},b_{j+1}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\leq c\mid z,y_{0}\right),$
+    $1-\sum_{j=2}^{3}\mathbb{P}(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq c\mid z,y_{0})$
+    $\geq\sum_{j=1}^{2}\mathbb{P}$ $\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}\{b_{j+1}-z_{1}^{\prime}\beta-y_{0}\gamma\leq c\},$
+    for any $c\in\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma,b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma\}_{j=2}^{T}$;
   • (2)
     When $s,t\in\{2,3\}$,
-    $1-\sum_{j=2}^{3}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq a\mid z,y_{0})$
-    $\geq\sum_{j=1}^{2}$ $\Pr\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq a\mid z,y_{0}\right),$
-    for any $a\in\{b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma,b_{j}-z_{t}^{\prime}\beta-\gamma,b_{j}-z_{t}^{\prime}\beta-2\gamma,b_{j}-z_{t}^{\prime}\beta-3\gamma\}_{j=2}^{3}$.
+    $1-\sum_{j=2}^{3}\mathbb{P}(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq c\mid z,y_{0})$
+    $\geq\sum_{j=1}^{2}$ $\mathbb{P}\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq c\mid z,y_{0}\right),$
+    for any $c\in\{b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma,b_{j}-z_{t}^{\prime}\beta-\gamma,b_{j}-z_{t}^{\prime}\beta-2\gamma,b_{j}-z_{t}^{\prime}\beta-3\gamma\}_{j=2}^{3}$.
 We normalize the first parameter $\beta_{0}$ to one, and report the performance of the coefficient $\gamma_{0}$ for the lagged dependent variable.
@@
 maximum 3.000 1.000 4.850
-We adopt various ordered response models introduced in Section 3 to analyze the income category.
+We adopt various ordered response models introduced in Section 3.2 to analyze the income category.
 The first model is the standard static model without any endogeneity.
@@
 Appendix A Appendix
-A.1 Proof of Lemma 1
+A.1 Proof of Proposition 2
 Proof.
-The identifying condition (7) for the parameter $\theta$ in Proposition 1 is given as
-$1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st}).$ (13)
-Now we show that when $X_{it}\in\{a_{1},...,a_{K}\}$, the above condition is informative only at the $2K$ points $a\in\mathcal{Q}_{1}=\{z_{t}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma,z_{s}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{s}^{\prime}\beta+a_{K}^{\prime}\gamma\}$.
-Let $q_{k}\in\mathcal{Q}_{1}$ denote an element in $\mathcal{Q}$.
-Without loss of generalization, we can rank $q_{k}$ as $q_{1}\leq q_{2}\leq...\leq q_{2K}$.
-When $a>q_{2K}$, then $a>z_{s}^{\prime}\beta+\max\limits_{k}{a_{k}^{\prime}\gamma}$.
-Therefore, the upper bound becomes $11-\Pr(Y_{is}=0,z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma\geq a\mid z_{st})=$ 1 and the identifying condition holds for any $\theta$.
-Therefore, there is no information for $\theta$ when $a>q_{2K}$.
-When $a<q_{1}$, similarly the lower bound is zero $\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st})=0$ and the inequality has no identifying power for $\theta$.
-When $a\in(q_{k},q_{k+1})$ for some $k$, the identifying restriction at $a$ can be implied by the restriction at $a=q_{k+1}$.
-It is because the upper bound at $q_{k+1}$ is the same as $a$ since there does not exist any point $z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$ that lies in $(q_{k},q_{k+1})$.
-The lower bound $\Pr(Y_{it}=1,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\leq a\mid z_{st})$ is weakly larger at $q_{k+1}$ than $a$ since it is weakly increasing.
-Then, any parameter that satisfies condition (13) evaluated at $q_{k+1}$ also satisfies the one at $a$, so condition (13) evaluated at $a$ does not contain additional information for $\theta$.
+Clearly, $\Theta_{I}\subseteq\Theta^{disc}$. Below we show $\Theta^{disc}\subseteq\Theta_{I}$ when $X_{it}$ is discrete. Suppose that $\theta$ satisfies condition (10) at all
+$c\in{\cal C}\left(\theta\right):=\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K,t=1,...,T\right\}$
+for any realization $z=\left(z_{1},...,z_{T}\right).$ We seek to show that $\theta$ must also satisfy condition (10) for any $c\in\mathcal{R}\backslash{\cal C}\left(\theta\right)$.
+Without loss of generality, we order elements in ${\cal C}\left(\theta\right)$ from the smallest to the largest as
+$\overline{c}_{1}\leq\overline{c}_{2}\leq...\leq\overline{c}_{KT}.$
+For $c<\overline{c}_{1}$, we must have
+$\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)\equiv 0,$
+so (10) holds trivially. Similarly, for $c>\overline{c}_{KT}$, we must have
+$\mathbb{P}\left(\left.Y_{is}=1,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)\equiv 0,$
+so (10) again holds trivially. For any $c$ s.t. $\overline{c}_{j}<c<\overline{c}_{j+1}$ for some $j$, we have
+$z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\quad\Leftrightarrow\quad z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq\overline{c}_{j}$
+and
+$z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\quad\Leftrightarrow\quad z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j+1}.$
+which implies
+$\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)=$ $\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq\overline{c}_{j}\right|Z_{i}=z\right)$ (26)
+and
+$\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)$ $=\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j+1}\right|Z_{i}=z\right)$
+$\leq\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j}\right|Z_{i}=z\right),$
+or equivalently,
+$1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j}\right|Z_{i}=z\right)\leq 1-\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j+1}\right|Z_{i}=z\right).$ (27)
+Since (10) holds at $\overline{c}_{j}$, we have
+$\max_{t}\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq\overline{c}_{j}\right|Z_{i}=z\right)\leq 1-\max_{s}\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq\overline{c}_{j}\right|Z_{i}=z\right).$
+Combining the above with (26) and (27), we have
+$\max_{t}\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)\leq 1-\max_{s}\mathbb{P}\left(\left.Y_{is}=0,\ z_{s}^{{}^{\prime}}\beta+X_{is}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right).$
 A.2 Proof of Theorem 1
-To prove sharpness of the identified set $\Theta_{I,1}$, we seek to show that, for any $\theta\in\Theta_{I,1}$, there exists an underlying DGP characterized by the conditional distribution of $(v_{is},v_{it})\mid W_{ist}$) that produces the same observed choice probability and satisfies Assumption 1.
-The following construction is for one candidate parameter $\theta\in\Theta_{I,1}$.
-We set the fixed effect to be zero $\alpha_{i}=0$, so that $v_{it}:=-(\epsilon_{it}+\alpha_{i})=-\epsilon_{it}$.
-Let $P_{st}\left(j,k\mid w_{st}\right):=\Pr\left(Y_{is}=j,Y_{it}=k\mid W_{ist}=w_{st}\right)$ denote the observed joint probability of selecting $j$ at period $s$ and selecting $k$ at period $t$.
-The first requirement for sharpness is that the constructed distribution of $(v_{is},v_{it})\mid W_{ist}$ matches the observed joint choice probability according to the model: given $W_{ist}=w_{st}$,
-$P_{st}\left(1,1\mid w_{st}\right)$ $=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$ (14)
-$P_{st}\left(1,0\mid w_{st}\right)$ $=\Pr(v_{is}\leq w_{s}^{\prime}\theta\mid w_{st})-\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$
-$P_{st}\left(0,1\mid w_{st}\right)$ $=\Pr(v_{it}\leq w_{t}^{\prime}\theta\mid w_{st})-\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right),$
-and $P_{st}\left(0,0\mid w_{st}\right)$ can be matched automatically when the above conditions hold.
-Another condition for sharpness requires the marginal distribution of $v_{it}\mid W_{ist}$ to satisfy the following stationarity condition in Assumption 1:
-$F_{v_{t}\mid Z_{st}}(a\mid z_{st})=F_{v_{s}\mid Z_{st}}(a\mid z_{st})\quad\forall a\in\mathcal{R}.$ (15)
-We can focus on a reduced problem, which is to find a marginal distribution $v_{it}\mid W_{ist}$ that matches the marginal choice probability and satisfies the stationarity assumption.
-Let $P_{t}(w_{st}):=\Pr(Y_{it}=1\mid W_{ist}=w_{st})$ and $P_{s}(w_{st}):=\Pr(Y_{is}=1\mid W_{ist}=w_{st})$ denote the observed marginal choice probability at $t$ and $s$, given as follows:
-$P_{t}(w_{st})=\Pr\left(v_{it}\leq w_{t}^{\prime}\theta\mid W_{ist}=w_{st}\right),$ (16)
-$P_{s}(w_{st})=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta\mid W_{ist}=w_{st}\right).$
-If we can construct a marginal distribution satisfying conditions (15) and (16), then the joint choice probability in (14) is matched by setting the joint distribution of $(v_{is},v_{it})\mid W_{ist}$ at $(w_{s}^{\prime}\theta,w_{t}^{\prime}\theta)$ as $P_{st}\left(1,1\mid w_{st}\right)=\Pr\left(v_{is}\leq w_{s}^{\prime}\theta,v_{it}\leq w_{t}^{\prime}\theta\mid w_{st}\right)$.
-This construction is feasible since there is no other assumptions on the joint distribution $(v_{is},v_{it})\mid W_{ist}$.
-Moreover, other choice probabilities such as $P_{st}\left(1,0\mid w_{st}\right)$ can be computed using the marginal probability $P_{s}(w_{st})$ subtracting $P_{st}\left(1,1\mid w_{st}\right)$.
-Now, for any $\theta\in\Theta_{I,1}$, our objective is to construct the distribution of $v_{it}\mid W_{st}$ and $v_{is}\mid W_{st}$ such that conditions (15) and (16) hold.
-Consider that the endogenous covariate $X_{it}$ only takes $K$ values: $X_{it}\in\mathcal{X}=\{a_{1},...,a_{K}\}$ for any $t\leq T$.
-[^10] The covariate $X_{it}$ is allowed to have different support across time $t$, then $\mathcal{X}$ is the union of the support of $X_{it}$ over all periods.
-Given $Z_{it}=z_{t}$, the covariate index $z_{t}^{\prime}\beta+x_{t}^{\prime}\gamma$ belongs to the $K$ points $\{z_{t}^{\prime}\beta+a_{1}^{\prime}\gamma,...,z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma\}$.
-Without loss of generalization, we can rank the $K$ points as $\alpha_{1}^{\prime}\gamma\leq a_{2}^{\prime}\gamma\leq...\leq a_{K}^{\prime}\gamma$ since we can reindex the $K$ values of $\{a_{1},...,a_{K}\}$ such that this ranking is satisfied.
-Let $m_{k}:=z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma$ denote the $K$ points for which $w_{t}^{\prime}\theta$ can take, and $n_{k}:=z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$ for $1\leq k\leq K$ denote the $K$ points $w_{s}^{\prime}\theta$ can take.
-Let $\bar{m}=\max\{z_{t}^{\prime}\beta,z_{s}^{\prime}\beta\}+a_{K}^{\prime}\gamma+\epsilon$ for some $\epsilon>0$.
-We first look at the stationarity condition in (15). Section 2.2 derives bounds for the distribution $F_{v_{t}\mid Z_{st}}(a\mid z_{st})$, given as follows:
-$F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ $\leq\sum_{x_{st}\in\mathcal{X}}\Big{\{}F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta\geq a\}+\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta<a\}\Big{\}}\Pr(X_{st}=x_{st}\mid z_{st})$
-$:=U_{t}(a\mid z_{st}),$
-$F_{v_{t}\mid Z_{st}}(a\mid z_{st})$ $\geq\sum_{x_{st}\in\mathcal{X}}F_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})\mathbf{\mathbbm{1}}\{w_{t}^{\prime}\theta\leq a\}\Pr(X_{st}=x_{st}\mid z_{st}):=L_{t}(a\mid z_{st}).$
-Given the definition of the identified set $\Theta_{I,1}$, for $\theta\in\Theta_{I,1}$, the following condition holds for any $a\in\mathcal{R}$,
-$\max_{s,t}\{L_{s}(a\mid z_{st}),L_{t}(a\mid z_{st})\}\leq\min_{s,t}\{U_{t}(a\mid z_{st}),U_{s}(a\mid z_{st})\},$
-Let $\tilde{L}(a\mid z_{st})=\max_{s,t}\{L_{t}(a\mid z_{st}),L_{s}(a\mid z_{st})\}$, the stationarity condition (15) is satisfied by constructing the conditional distribution of $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ as follows:
-$F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})=\left\{\begin{aligned} &0\ &\text{if }a<m_{1},\\ \tilde{L}&(a\mid z_{st})\ &\text{if }m_{1}\leq a<\bar{m},\\ &1\ &\text{if }\bar{m}\leq a.\end{aligned}\right.$ (17)
-The above construction guarantees that the marginal distribution $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})$ lies between the bounds $[L_{s}(a\mid z_{st}),U_{s}(a\mid z_{st})]$.
-From the definition of $L_{t}(a)$, we know that $L_{t}(a\mid z_{st})\geq L_{t}(a^{\prime}\mid z_{st})$ for any $a\geq a^{\prime}$.
-Therefore, the monotonicity of the above distribution is satisfied as $\max\limits_{s,t}\{L_{t}(a\mid z_{st}),L_{s}(a\mid z_{st})\}\geq\max\limits_{s,t}\{L_{t}(a^{\prime}\mid z_{st}),L_{s}(a^{\prime}\mid z_{st})\}$ for any $a\geq a^{\prime}$.
-Additionally, the above construction satisfies the stationarity assumption given $F^{*}_{v_{s}\mid Z_{st}}(a\mid z_{st})=F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ for any $a$.
-We transform the two requirements of sharpness into finding a distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ to generate the marginal distribution $F^{*}_{v_{t}\mid Z_{st}}(a\mid z_{st})$ in (17) and also match the marginal choice probability given in condition (16).
-Construction
-We focus on the construction for $v_{it}\mid W_{ist}$, and the construction for $v_{is}\mid W_{ist}$ at period $s$ is similar so it is omitted here.
-The construction proceeds in four steps.
-We first constructs a distribution to match the lower and upper bounds $L_{t}(a\mid z_{st})$ and $U_{t}(a\mid z_{st})$.
-Then, we construct the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ at the points $\{m_{1},...,m_{K}\}$ with $m_{k}=z_{t}^{\prime}\beta+a_{k}^{\prime}\gamma$.
-The third step look at the construction at the points $\{n_{1},...,n_{K}\}$ with $n_{k}=z_{s}^{\prime}\beta+a_{k}^{\prime}\gamma$.
-The final step examines the points $a$ that are not in $\{m_{1},...,m_{K},n_{1},...,n_{K}\}$.
+For shorter notation we write $W_{it}:=\left(Z_{it},X_{it}\right)$ for the combination of the exogenous covariates $Z_{it}$ and endogenous covariates $X_{it}$.
+Correspondingly, we write $W_{i}\equiv\left(W_{i1},...,W_{iT}\right)$, $W_{i}^{{}^{\prime}}\theta_{0}\equiv Z_{i}^{{}^{\prime}}\beta_{0}+X_{i}^{{}^{\prime}}\gamma_{0}$ and the lower cases $w\equiv\left(w_{1},...,w_{T}\right)$ for realizations.
+We first clarify the rigorous meaning of “sharpness” in Theorem 1 through the following definition.
+Definition 1.
+We say that $\Theta_{I}$ is sharp under model (5) and Assumption 1 if, for any $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, there exist well-defined latent random variables $\left(\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ such that:
+  • Assumption 1 (partial stationarity) is satisfied, i.e.,
+    $\left.\epsilon_{it}^{*}\sim\epsilon_{is}^{*}\right|Z_{i},\alpha_{i}^{*},\ \forall t,s=1,...,T.$
+  • (CCP-J) $\left(\theta,\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ are observationally equivalent to $\left(\theta_{0},\epsilon_{i},\alpha_{i}\right)$, i.e., formally, $\left(\theta,\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ produces the following conditional choice probabilities under model (5):
+    $\mathbb{P}\left(v_{it}^{*}\leq w_{t}^{{}^{\prime}}\theta\ \forall t\text{ s.t.
+}y_{t}=1,\ v_{is}^{*}>w_{s}^{{}^{\prime}}\theta\ \forall s\text{ s.t.
+}y_{s}=0\mid w\right)=p\left(y\mid w\right),$ (28)
+    where $v_{it}^{*}:=-\left(\epsilon_{it}^{*}+\alpha_{i}^{*}\right)$ and $p\left(\left.\cdot\right|w\right)$ denotes the true conditional probability
+    $p\left(\left.y\right|w\right)$ $:=\mathbb{P}\left(Y_{it}=y_{t}\,\ \forall t=1,...,T\mid W_{i}=w\right)$
+    $\equiv\mathbb{P}\left(v_{it}\leq w_{t}^{{}^{\prime}}\theta_{0}\ \forall t\text{ s.t. }y_{t}=1,\ v_{is}>w_{s}^{{}^{\prime}}\theta\ \forall s\text{ s.t. }y_{s}=0\mid W_{i}=w\right),$
+    for any outcome realization $y\equiv\left(y_{1},...,y_{T}\right)\in\left\{0,1\right\}^{T}$, given any realization $W_{i}=w$.
+We prove Theorem 1, i.e., the sharpness of the identified set $\Theta_{I}$ under discreteness of $X_{it}$ by, for any candidate parameter $\theta\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, we construct the distribution of $\left(\epsilon_{i}^{*},\alpha_{i}^{*}\right)$ to satisfy the above requirements.
+Proof.
+Set $\alpha_{i}^{*}\equiv 0$ and $\epsilon_{i}^{*}:=-v_{i}^{*}$. Then the conclusion follows from Lemma 1 and 2 below. ∎
+Lemma 1 (Per-Period Construction).
+When $X_{it}\in\left\{\overline{x}_{1},...,\overline{x}_{K}\right\}$ for all $t=1,...,T$, then for any $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, there exist well-defined latent random variables $v_{i1}^{*},...,v_{iT}^{*}$ with marginal CDFs $F_{1}^{*},...,F_{T}^{*}$ such that
+$F_{t}^{*}\left(\left.\cdot\right|Z_{i}=z\right)=F_{s}^{*}\left(\left.\cdot\right|Z_{i}=z\right)$ (29)
+and
+$F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right)=p_{t}\left(w\right),\quad\forall t,\forall w,$ (30)
+where
+$p_{t}\left(w\right):=\mathbb{P}\left(\left.Y_{it}=1\right|W_{i}=w\right).$
+Proof.
+For any $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, below we show how to construct $v_{i1}^{*},...,v_{iT}^{*}$, or equivalently, the conditional CDFs $F_{1}^{*}\left(\left.c\right|W_{i}=w\right),...,F_{T}^{*}\left(\left.c\right|W_{i}=w\right)$ for each realization $w$ and each $c\in\mathcal{R}$ so that (i) condition (29) is satisfied so that partial stationarity holds; and (ii) condition (30) is satisfied so that per-period marginal CCPs are matched.
+Fix a specific realization of the exogenous covariates at $z\equiv\left(z_{1},...,z_{T}\right)$.
+We construct the (conditional) CDF $F_{t}^{*}$ of $v_{it}^{*}$ for each $t=1,...,T$ and each given $z$ in the following manner.
+Define
+$L_{t}\left(\left.c\right|z\right)$ $:=\mathbb{P}\left(\left.Y_{it}=1,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right),$
+$U_{t}\left(\left.c\right|z\right)$ $:=1-\mathbb{P}\left(\left.Y_{it}=0,\ z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right),$
+and
+$\overline{L}\left(\left.c\right|z\right):=\max_{s}L_{s}\left(\left.c\right|z\right),\quad$ $\underline{U}\left(\left.c\right|z\right):=\min_{s}U_{s}\left(\left.c\right|z\right).$
+Since $\theta\equiv\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\Theta_{I}\backslash\left\{\theta_{0}\right\}$, by 10 we have,
+$\overline{L}\left(\left.c\right|z\right)\leq\underline{U}\left(\left.c\right|z\right),\quad\forall c\in\mathcal{R}.$
+Observe that both $\overline{L}\left(\left.c\right|z\right)$ and $\underline{U}\left(\left.c\right|z\right)$ are weakly increasing in $c$.
+[…]
+${\cal C}:=\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:t=1,...,T,k=1,...,K\right\}.$
+Let $\delta>0$ be a sufficiently small constant [^10] The small positive constant $\delta>0$ is used to ensure the right continuity of CDFs defined afterwards.
+Let $\underline{\delta}$ to be smallest distance between two distinct points in ${\cal C}$.
+If $\underline{\delta}>0$, then we may set $\delta:=\underline{\delta}/2$.
+If $\underline{\delta}=0$, then $\delta$ can be set as any positive number, say, $\delta:=1$., and define
+$\underline{c}:=\min{\cal C},\quad\overline{c}:=\max{\cal C}+\delta.$
+The parametrix index $w_{t}^{{}^{\prime}}\theta$ must lie within the interval $\left[\underline{c},\overline{c}\right)$ across all $t$ for all possible realization of $x$.
+For each $t=1,...,T,$ we show how to construct $v_{t}^{*}$ with CDF $F_{t}^{*}$ so that
+$F_{t}^{*}\left(\left.c\right|z\right)\equiv\begin{cases}0,&\text{if }c<\underline{c},\\ \overline{L}\left(\left.c\right|z\right),&\text{if }\underline{c}\leq c<\overline{c},\\ 1,&\text{if }c\geq\overline{c},\end{cases}$ (31)
+and
+$F_{t}^{*}\left(\left.w_{t}^{\prime}\theta\right|w\right)=p_{t}\left(w\right).$ (32)
+Clearly, partial stationarity (29) will be satisfied under (31), the right-hand side of which does not depend on the time index $t$.
+Furthermore, (32) is the same as (30), i.e., the marginal CCPs will be matched for each $t$.
+∎
+Proof.
 Step 1:
-We first show that the bounds $U_{t}(a\mid z_{st})$ and $L_{t}(a\mid z_{st})$ can be achieved under a distribution of $v_{it}\mid W_{ist}$ that matches the marginal choice in (16).
-From its definition, the upper bound $U_{t}(a\mid z_{st})$ can be achieved under the following distribution $F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st})$:
-$F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st})=\left\{\begin{aligned} &0\qquad&\text{if }a<m_{1},\\ &P_{t}(w_{st})\qquad&\text{if }m_{1}\leq a<w_{t}^{\prime}\theta+\epsilon,\\ &1\qquad&\text{if }w_{t}^{\prime}\theta+\epsilon\leq a,\end{aligned}\right.$
-where $\epsilon>0$ is used to satisfy the right-continuity property of the distribution.
-Similarly, the lower bound function $L_{t}(a\mid z_{st})$ can be achieved under the following distribution $F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})$:
-$F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})=\left\{\begin{aligned} &0\qquad&\text{if }a<w_{t}^{\prime}\theta,\\ &P_{t}(w_{st})\qquad&\text{if }w_{t}^{\prime}\theta\leq a<\bar{m},\\ &1&\text{if }\bar{m}\leq a\end{aligned}\right.$
-From the definition of the distributions $F^{U}_{v_{t}\mid W_{st}}(a\mid w_{st}),F^{L}_{v_{t}\mid W_{st}}(a\mid w_{st})$, they both satisfy condition (16) which requires $F^{U}_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})=F^{L}_{v_{t}\mid W_{st}}(w_{t}^{\prime}\theta\mid w_{st})=P_{t}(w_{st})$.
-Step 2: this step establishes the construction of the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ at the points $m_{k}$.
-The observed marginal distribution given in condition (16) only restricts the distribution of $v_{it}\mid W_{ist}$ at each point $m_{k}$, given $w_{st}=(z_{st},x_{s},a_{k})$,
-$P_{t}(w_{st})=\Pr(v_{it}\leq m_{k}\mid z_{st},x_{s},a_{k}).$
-(i) We start with the largest value $m_{K}=z_{t}^{\prime}\beta+a_{K}^{\prime}\gamma$.
-From the definition of $U_{t}(a\mid z_{st}),L_{t}(a\mid z_{st})$, we know that $U_{t}(m_{1}\mid z_{st})=L_{t}(m_{K}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})\leq U_{t}(m_{K}\mid z_{st})$, which is described in the following graph:
-$U_{t}(m_{1}\mid z_{st})$$=L_{t}(m_{K}\mid z_{st})$$U_{t}(m_{K-1}\mid z_{st})$$\tilde{L}(m_{K}\mid z_{st})$$U_{t}(m_{K}\mid z_{st})$
-Given the $K$ points $U_{t}(m_{1}\mid z_{st})\leq U_{t}(m_{2}\mid z_{st})\leq...\leq U_{t}(m_{K}\mid z_{st})$, we can find $2\leq j_{1}\leq K$ such that $U_{t}(m_{j_{1}-1}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})\leq U_{t}(m_{j_{1}}\mid z_{st})$.
-Then we can express $\tilde{L}(m_{K}\mid z_{st})$ as follows: for some $\alpha_{1}\in[0,1]$,
-$\tilde{L}(m_{K}\mid z_{st})=\alpha_{1}U_{t}(m_{j_{1}-1}\mid z_{st})+(1-\alpha_{1})U_{t}(m_{j_{1}}\mid z_{st}).$
-Accordingly, we construct the distribution $F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$ as follows:
-$F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})=\alpha_{1}F^{U}_{t}(m_{j_{1}-1}\mid w_{st})+(1-\alpha_{1})F^{U}_{t}(m_{j_{1}}\mid w_{st}).$
-The above construction generates the marginal distribution $\tilde{L}(m_{K}\mid z_{st})$.
-Moreover, it matches the marginal choice probability when $x_{t}=a_{K}$ in condition (16) since $F^{U}_{t}(m_{j}\mid w_{st})=P_{t}(w_{st})$ if $x_{t}=a_{K}$ for any $1\leq j\leq K$ from the definition of $F^{U}_{t}(m_{j}\mid w_{st})$.
-(ii) Now consider that $a=m_{K-1}$.
-Given that $L_{t}(m_{K-1}\mid z_{st})\leq\tilde{L}(m_{K-1}\mid z_{st})\leq U_{t}(m_{K-1}\mid z_{st})$, there either exists $11\leq j_{2}\leq K-$ 1 such that $\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})]$ or $\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})]$, given in the following graph:
-$L_{t}(m_{K-1}\mid z_{st})$$U_{t}(m_{1}\mid z_{st})$$=L_{t}(m_{K}\mid z_{st})$$U_{t}(m_{K-1}\mid z_{st})$$\tilde{L}(m_{K-1}\mid z_{st})$$U_{t}(m_{K}\mid z_{st})$
-Then $\tilde{L}(m_{K-1}\mid z_{st})$ can be expressed as follows, for $\alpha_{2}\in[0,1]$ and $\tilde{\alpha}_{2}\in[0,1]$,
-$\tilde{L}(m_{K-1}\mid z_{st})$
-$=$ $&\alpha_{2}U_{t}(m_{j_{2}-1}\mid z_{st})+(1-\alpha_{2})U_{t}(m_{j_{2}}\mid z_{st})\ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})],\\ &\tilde{\alpha}_{2}L_{t}(m_{K-1}\mid z_{st})+(1-\tilde{\alpha}_{2})U_{t}(m_{1}\mid z_{st})\ \ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})].\end{aligned}\right.$
-Accordingly, the distribution $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$ at $m_{K-1}$ is constructed as
-$F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$
-$=$ $&\alpha_{2}F^{U}_{t}(m_{j_{2}-1}\mid w_{st})+(1-\alpha_{2})F^{U}_{t}(m_{j_{2}}\mid w_{st})\ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[U_{t}(m_{j_{2}-1}\mid z_{st}),U_{t}(m_{j_{2}}\mid z_{st})],\\ &\tilde{\alpha}_{2}F^{L}_{t}(m_{K-1}\mid z_{st})+(1-\tilde{\alpha}_{2})F^{U}_{t}(m_{1}\mid w_{st})\ \ \text{ if }\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})].\end{aligned}\right.$
-Condition (16) is satisfied since $U_{t}(m_{j}\mid w_{st})=P_{t}(w_{st})$ when $x_{t}=a_{K-1}$ for any $11\leq j\leq K-$ 1.
-Now we still need to show that the constructed distribution satisfies monotonicity.
-If $\tilde{L}(m_{K-1}\mid z_{st})\in[L_{t}(m_{K-1}\mid z_{st}),U_{t}(m_{1}\mid z_{st})]$, then $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{U}_{t}(m_{1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$.
-Otherwise, by the monotonicity of $\tilde{L}(a\mid z_{st})$ in $a$, we know that $j_{2}\leq j_{1}$.
-If $j_{2}\leq j_{1}-1$, then monotonicity is satisfied since $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{U}_{t}(m_{j_{2}}\mid w_{st})\leq F^{U}_{t}(m_{j_{1}-1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$.
-If $j_{1}-1<j_{2}=j_{1}$, then it implies that $\alpha_{2}\leq\alpha_{1}$ since $\tilde{L}(m_{K-1}\mid z_{st})\leq\tilde{L}(m_{K}\mid z_{st})$.
-Therefore, we have $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})\leq F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})$.
-The distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ for the remaining points $a\in\{m_{1},...,m_{K-2}\}$ can be constructed in the same way as $F^{*}_{v_{t}\mid W_{st}}(m_{K-1}\mid w_{st})$, so it is omitted here.
-Step 3: we show the construction at the points $a=n_{k}$ for $1\leq k\leq K$.
-For points $a=n_{k}$, we only need to construct a proper distribution to be consistent with Equation (17).
-We still start with the largest point $n_{K}$.
-The distribution $\tilde{L}(n_{K}\mid z_{st})$ can be expressed as follows in different scenarios: for some $2\leq j_{3}\leq K$
-$\tilde{L}(n_{K}\mid z_{st})=\left\{\begin{aligned} &\alpha_{3}\tilde{L}(m_{K}\mid z_{st})+(1-\alpha_{3})*1&\text{ if }n_{K}>m_{K},\\ &\alpha_{4}\tilde{L}(m_{j_{3}-1}\mid z_{st})+(1-\alpha_{4})\tilde{L}(m_{j_{3}}\mid z_{st})&\text{ if }m_{j_{3}-1}<n_{K}<m_{j_{3}},\\ &\alpha_{5}*0+(1-\alpha_{5})\tilde{L}(m_{1}\mid z_{st})&\text{ if }n_{K}<m_{1},\end{aligned}\right.$
-where $\alpha_{3},\alpha_{4},\alpha_{5}\in[0,1]$.
-Accordingly, the distribution $F^{*}_{v_{t}\mid W_{st}}(n_{k}\mid w_{st})$ is constructed as follows,
-$F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})=\left\{\begin{aligned} &\alpha_{3}F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})+(1-\alpha_{3})*1&\text{ if }n_{K}>m_{K},\\ &\alpha_{4}F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}-1}\mid w_{st})+(1-\alpha_{4})F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}}\mid w_{st})&\text{ if }m_{j_{3}-1}<n_{K}<m_{j_{3}},\\ &\alpha_{5}*0+(1-\alpha_{5})F^{*}_{v_{t}\mid W_{st}}(m_{j_{1}}\mid w_{st})&\text{ if }n_{K}<m_{1}.\end{aligned}\right.$
-When $a=n_{K-1}$, the distribution $\tilde{L}(n_{K-1}\mid z_{st})$ can be expressed as follows in different scenarios: for some $11\leq j_{4}\leq j_{3}-$ 1,
-$\tilde{L}(n_{K-1}\mid z_{st})=\left\{\begin{aligned} &\alpha_{6}\tilde{L}(m_{K}\mid z_{st})+(1-\alpha_{6})\tilde{L}(n_{K}\mid z_{st})&\text{ if }m_{K}<n_{K-1}<n_{K},\\ &\alpha_{7}\tilde{L}(m_{j_{3}-1}\mid z_{st})+(1-\alpha_{7})\tilde{L}(n_{K}\mid z_{st})&\text{ if }m_{j_{3}-1}<n_{K-1}<n_{K}<m_{j_{3}},\\ &\alpha_{8}\tilde{L}(m_{j_{4}-1}\mid z_{st})+(1-\alpha_{8})\tilde{L}(m_{j_{4}}\mid z_{st})&\text{ if }m_{j_{4}-1}<n_{K-1}<m_{j_{4}}\leq m_{j_{3}-1},\\ &\alpha_{9}*0+(1-\alpha_{9})\tilde{L}(n_{K}\mid z_{st})&\text{ if }n_{K-1}<n_{K}<m_{1},\end{aligned}\right.$
-where $\alpha_{6},\alpha_{7},\alpha_{8},\alpha_{9}\in[0,1]$.
-The distribution $F^{*}_{v_{t}\mid W_{st}}(n_{K-1}\mid w_{st})$ is constructed accordingly:
-$F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})$
-$=$ $&\alpha_{6}F^{*}_{v_{t}\mid W_{st}}(m_{K}\mid w_{st})+(1-\alpha_{6})F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})&\text{ if }m_{K}<n_{K-1}<n_{K},\\ &\alpha_{7}F^{*}_{v_{t}\mid W_{st}}(m_{j_{3}-1}\mid w_{st})+(1-\alpha_{7})F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})&\text{ if }m_{j_{3}-1}<n_{K-1}<n_{K}<m_{j_{3}},\\ &\alpha_{8}F^{*}_{v_{t}\mid W_{st}}(m_{j_{4}-1}\mid w_{st})+(1-\alpha_{8})F^{*}_{v_{t}\mid W_{st}}(m_{j_{4}}\mid w_{st})&\text{ if }m_{j_{4}-1}<n_{K-1}<m_{j_{4}}\leq m_{j_{3}-1},\\ &(1-\alpha_{9})F^{*}_{v_{t}\mid W_{st}}(n_{K}\mid w_{st})&\text{ if }n_{K-1}<n_{K}<m_{1}.\end{aligned}\right.$
-The monotonicity of the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ among the $2K$ points $\{m_{1},...,m_{K},n_{1},...,n_{K}\}$ is satisfied by construction.
-Furthermore, the construction for the rest of points $\{n_{1},...,n_{K-2}\}$ are similar so it is omitted here.
-Step 4: For any $a\notin\{m_{1},...,m_{K},n_{1},...,n_{K}\}$, The distribution is constructed as $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})=0$ if $a<m_{1}$ and $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})=1$ if $a\geq\bar{m}$.
-When $a\in[m_{1},\bar{m}]$, there must exist some $q_{j}\in\{m_{1},...,m_{K},n_{1},...,n_{K}\}$ such that $q_{j}<a$ and $\tilde{L}(a\mid z_{st})=\tilde{L}(q_{j}\mid z_{st})$.
-Then we set $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})=F^{*}_{v_{t}\mid W_{st}}(m_{j}\mid w_{st})$.
-This completes the construction for the distribution $F^{*}_{v_{t}\mid W_{st}}(a\mid w_{st})$ for any $a$.
-A.3 Proof of Proposition 2 and 3
+We construct the conditional CDF of $v^{*}|W_{i}=w$ using two auxillary CDFs $F_{t}^{L}$ and $F_{t}^{U}$, defined by
+$F_{t}^{L}\left(\left.c\right|w\right)$ =$=\begin{cases}0,&c<w_{t}^{{}^{\prime}}\theta,\\ p_{t}\left(w\right),&w_{t}^{{}^{\prime}}\theta\leq c<\overline{c}_{t},\\ 1,&c\geq\overline{c}_{t},\end{cases}$
+and
+$F_{t}^{U}\left(\left.c\right|w\right)=\begin{cases}0,&c<\underline{c}_{t},\\ p_{t}\left(w\right),&\underline{c}_{t}\leq c<w_{t}^{{}^{\prime}}\theta+\delta,\\ 1,&c\geq w_{t}^{{}^{\prime}}\theta+\delta.\end{cases}$
+where
+$\overline{c}_{t}:=\max{\cal C}_{t}+\delta,\quad\underline{c}_{t}:=\min{\cal C}_{t},\quad{\cal C}_{t}:=\left\{z_{t}^{{}^{\prime}}\beta+\overline{x}_{k}^{{}^{\prime}}\gamma:k=1,...,K\right\}.$
+Clearly, by construction we have
+$F_{t}^{L}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=F_{t}^{U}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=p_{t}\left(w\right).$ (33)
+Furthermore, for any $c<\overline{c}_{t}$, we have
+$F_{t}^{L}\left(\left.c\right|z\right)$ $=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)F_{t}^{L}\left(\left.c\right|w\right)$
+$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\mathbf{\mathbbm{1}}\left\{w_{t}^{{}^{\prime}}\theta\leq c\right\}p_{t}\left(w\right)$
+$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\mathbb{P}\left(\left.Y_{i}=1\text{ and }w_{t}^{{}^{\prime}}\theta\leq c\right|W_{i}=w\right)$
+$=\mathbb{P}\left(\left.Y_{i}=1\text{ and }z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\leq c\right|Z_{i}=z\right)$
+$=L_{t}\left(\left.c\right|z\right),$
+while for $c\geq\overline{c}_{t}$, we have
+$F_{t}^{L}\left(\left.c\right|z\right)=1=U_{t}\left(\left.c\right|z\right).$
+Similarly, for each $c\geq\underline{c}_{t}$, we have
+$F_{t}^{U}\left(\left.c\right|z\right)$ $=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)F_{t}^{U}\left(\left.c\right|w\right)$
+$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\left[1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }w_{t}^{{}^{\prime}}\theta\geq c-\delta\right|W_{i}=w\right)\right]$
+$=\sum_{x}\mathbb{P}\left(\left.X_{i}=x\right|Z_{i}=z\right)\left[1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }w_{t}^{{}^{\prime}}\theta\geq c-\delta\right|W_{i}=w\right)\right]$
+$=1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c-\delta\right|Z_{i}=z\right)$
+$=1-\mathbb{P}\left(\left.Y_{i}=0\text{ and }z_{t}^{{}^{\prime}}\beta+X_{it}^{{}^{\prime}}\gamma\geq c\right|Z_{i}=z\right)$
+$=U_{t}\left(\left.c\right|z\right),$
+where the second last equality holds for sufficiently small $\delta>0$ due to the discreteness of ${\cal C}$. Lastly, for $c<\underline{c}_{t}$, we have
+$F_{t}^{U}\left(\left.c\right|z\right)=0=L_{t}\left(\left.c\right|z\right).$
+In summary, we have
+$F_{t}^{L}\left(\left.c\right|z\right)$ =$=\begin{cases}L_{t}\left(\left.c\right|z\right),&\forall c<\overline{c}_{t},\\ U_{t}\left(\left.c\right|z\right)=1,&\forall c\geq\overline{c}_{t},\end{cases}$
+$F_{t}^{U}\left(\left.c\right|z\right)$ =$=\begin{cases}L_{t}\left(\left.c\right|z\right)=0,&\forall c<\underline{c}_{t},\\ U_{t}\left(\left.c\right|z\right),&\forall c\geq\underline{c}_{t},\end{cases}$ (34)
+Furthermore, observe that
+$L_{t}\left(\left.\cdot\right|z\right)\leq F_{t}^{L}\left(\left.\cdot\right|z\right)\leq F_{t}^{U}\left(\left.\cdot\right|z\right)\leq U_{t}\left(\left.\cdot\right|z\right).$
+Step 2:
+Now, we construct $F_{t}^{*}\left(\left.c\right|w\right)$ for $c\in{\cal C}_{t}$ using the two auxillary CDFs $F_{t}^{L}\left(\left.c\right|w\right)$ and $F_{t}^{U}\left(\left.c\right|w\right)$.
+We rank-order elements in ${\cal C}_{t}$ in ascending order
+$\underline{c}_{t}\equiv c_{t1}\leq c_{t2}\leq...\leq c_{tK}<\overline{c}_{t}\equiv c_{tK}+\delta.$
+(i) We start with the largest element $c_{tK}$. By the definition of $L_{t}$, $U_{t}$, $\overline{L}$ and (10), we know that
+$U_{t}\left(\left.c_{t1}\right|z\right)=L_{t}\left(\left.c_{tK}\right|z\right)\leq\overline{L}\left(\left.c_{tK}\right|z\right)\leq U_{t}\left(\left.c_{tK}\right|z\right).$
+Hence, we can find $2\leq j_{1}\leq K$ such that
+$U_{t}\left(\left.c_{t,j_{1}-1}\right|z\right)\leq\overline{L}\left(\left.c_{tK}\right|z\right)\leq U_{t}\left(\left.c_{t,j_{1}}\right|z\right),$
+so that there exists $\alpha_{1}\in\left[0,1\right]$ such that
+$\overline{L}\left(\left.c_{tK}\right|z\right)=\alpha_{1}U_{t}\left(\left.c_{t,j_{1}-1}\right|z\right)+\left(1-\alpha\right)U_{t}\left(\left.c_{t,j_{1}}\right|z\right).$ (35)
+Then, we set
+$F_{t}^{*}\left(\left.c_{tK}\right|w\right):=\alpha_{1}F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right)+\left(1-\alpha\right)F_{t}^{U}\left(\left.c_{t,j_{1}}\right|w\right),$ (36)
+which ensures that
+$F_{t}^{*}\left(\left.c_{tK}\right|z\right)=\overline{L}\left(\left.c_{tK}\right|z\right).$ (37)
+Furthermore, whenever $w$ is such that $w_{t}^{{}^{\prime}}\theta=c_{tK}$, we have
+$F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=F_{t}^{*}\left(\left.c_{tK}\right|w\right)=p_{t}\left(w\right).$
+(ii) Next, we consider the point $c_{t,K-1}$. Given that
+$L_{t}\left(\left.c_{t,K-1}\right|z\right)\leq\overline{L}\left(\left.c_{t,K-1}\right|z\right)\leq U_{t}\left(\left.c_{t,K-1}\right|z\right),$
+then either there exists some $11\leq j_{2}\leq K-$ 1 such that
+$\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right),U_{t}\left(\left.c_{t,j_{2}}\right|z\right)\right]$
+or
+$\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[L_{t}\left(\left.c_{t,K-1}\right|z\right),L_{t}\left(\left.c_{t,K}\right|z\right)\equiv U_{t}\left(\left.c_{t,1}\right|z\right)\right].$
+Hence, either there exists $\alpha_{2}\in\left[0,1\right]$ s.t.
+$\overline{L}\left(\left.c_{t,K-1}\right|z\right)=\alpha_{2}U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right)+\left(1-\alpha_{2}\right)U_{t}\left(\left.c_{t,j_{2}}\right|z\right)$ (38)
+or there exists $\tilde{\alpha}_{2}\in\left[0,1\right]$ s.t.
+$\overline{L}\left(\left.c_{t,K-1}\right|z\right)=\tilde{\alpha}_{2}L_{t}\left(\left.c_{t,K-1}\right|z\right)+\left(1-\alpha_{2}\right)\tilde{\alpha}_{2}L_{t}\left(\left.c_{t,K}\right|z\right),$ (39)
+so that we can set
+$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)$ :=$:=\begin{cases}\alpha_{2}F_{t}^{U}\left(\left.c_{t,j_{2}-1}\right|w\right)+\left(1-\alpha_{2}\right)F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)\\ \quad\quad\text{if }\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right),U_{t}\left(\left.c_{t,j_{2}}\right|z\right)\right],\\ \tilde{\alpha}_{2}F_{t}^{L}\left(\left.c_{t,K-1}\right|w\right)+\left(1-\tilde{\alpha}_{2}\right)F_{t}^{L}\left(\left.c_{t,K}\right|w\right)\\ \quad\quad\text{if }\left[L_{t}\left(\left.c_{t,K-1}\right|z\right),L_{t}\left(\left.c_{t,K}\right|z\right)\right].\end{cases}$ (40)
+which again ensures
+$F_{t}^{*}\left(\left.c_{t,K-1}\right|z\right)=\overline{L}\left(\left.c_{t,K-1}\right|z\right).$
+We now show that the constructions of $F_{t}^{*}\left(\left.c\right|w\right)$ at $c=c_{t,K}$ and $c_{t,K-1}$ in (36) and (40) satisfy the monotonicity requirement of a CDF, i.e.,
+$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)\leq F_{t}^{*}\left(\left.c_{t,K}\right|w\right).$ (41)
+To see this, first notice from (36) and the monotonicity of $F_{t}^{U}$ that
+$F_{t}^{*}\left(\left.c_{tK}\right|w\right)\geq F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right)$ (42)
+Next, consider the two cases in (40).
+If $\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[L_{t}\left(\left.c_{t,K-1}\right|z\right),L_{t}\left(\left.c_{t,K}\right|z\right)\right],$then by the construction and the monotonicity of $F_{t}^{L},F_{t}^{U}$, we have
+$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)\leq F_{t}^{L}\left(\left.c_{t,K-1}\right|w\right)\leq p_{t}\left(w\right)=F_{t}^{U}\left(\left.c_{t,1}\right|w\right),$
+which together with (42) imply (41).
+If $\overline{L}\left(\left.c_{t,K-1}\right|z\right)\in\left[U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right),U_{t}\left(\left.c_{t,j_{2}}\right|z\right)\right]$, then by the monotonicity of $U_{t}\left(\left.\cdot\right|z\right)$ we know that $j_{2}\leq j_{1}$.
+If in addition $j_{2}\leq j_{1}-1$, then by the monotonicity of $F_{t}^{U}$ we have
+$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)\leq F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)\leq F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right),$
+which together with (42) imply (41). Otherwise, we must have $j_{2}=j_{1}$, then by (35), (38) and the monotonicity of $\overline{L}\left(\left.\cdot\right|z\right)$, we must have
+$\alpha_{2}\geq\alpha_{1}$
+since $\overline{L}\left(\left.c_{t,K-1}\right|z\right)\leq\overline{L}\left(\left.c_{t,K}\right|z\right)$ and $U_{t}\left(\left.c_{t,j_{2}-1}\right|z\right)\leq U_{t}\left(\left.c_{t,j_{2}}\right|z\right)$ Therefore, by (36), (40), and the monotonicity of $F_{t}^{U}$, we have
+$F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)$ $=\alpha_{2}F_{t}^{U}\left(\left.c_{t,j_{2}-1}\right|w\right)+\left(1-\alpha_{2}\right)F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)$
+$\leq\alpha_{1}F_{t}^{U}\left(\left.c_{t,j_{1}-1}\right|w\right)+\left(1-\alpha_{1}\right)F_{t}^{U}\left(\left.c_{t,j_{2}}\right|w\right)=F_{t}^{*}\left(\left.c_{t,K}\right|w\right).$
+(iii) The construction of $F_{t}^{*}\left(\left.c\right|w\right)$ at $c=c_{t,1},...,c_{t,K-2}$ can be carried out in the same way as $F_{t}^{*}\left(\left.c_{t,K-1}\right|w\right)$, and is thus omitted here.
+In summary, $F_{t}^{*}$ is constructed so that
+$F_{t}^{*}\left(\left.c\right|z\right)=\overline{L}\left(\left.c\right|z\right),\quad\forall c\in{\cal C}_{t}$
+Furthermore, since for each $c\in{\cal C}_{t}$, the value $F_{t}^{*}\left(\left.c\right|w\right)$ is always constructed as a weighted average of (two) values in
+$\left\{F^{L}\left(\left.c\right|w\right),F^{U}\left(\left.c\right|w\right):\ c\in{\cal C}\right\},$
+by (32), we have
+$F_{t}^{*}\left(\left.c\right|w\right)=p_{t}\left(w\right),\ \forall c\in{\cal C}_{t}.$
+Hence, for any $w$, we must have $w_{t}^{{}^{\prime}}\theta\in{\cal C}_{t}$ and thus
+$F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)=p_{t}\left(w\right),$
+which ensures (30).
+Step 3:
+We now show how to construct $F_{t}^{*}\left(\left.c\right|w\right)$ at $c\in{\cal C}_{-t}:={\cal C}\backslash{\cal C}_{t}$ to ensure (31), using based on the previously assigned values of $F_{t}^{*}\left(\left.c\right|w\right)$ for $c\in{\cal C}_{t}$ in Step 2.
+We rank-order elements in ${\cal C}_{-t}$ in strict ascending order as follows
+$c_{-t,1}<...<c_{-t,\overline{K}}\text{ for some }\overline{K}\geq 1.$
+(i) We start with the largest element $c_{-t,\overline{K}}$. Then $\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right)$ can be expressed as follows in three different scenarios:
+$\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right):=\begin{cases}\alpha_{3}\overline{L}\left(\left.c_{t,K}\right|z\right)+\left(1-\alpha_{3}\right)\cdot 1,&\text{if }c_{-t,\overline{K}}>c_{t,K},\\ \alpha_{4}\overline{L}\left(\left.c_{t,j-1}\right|z\right)+\left(1-\alpha_{3}\right)\overline{L}\left(\left.c_{t,j}\right|z\right),&\text{if }c_{-t,\overline{K}}\in\left(c_{t,j-1},c_{t,j}\right)\text{ for some }j,\\ \alpha_{5}\cdot 0+\left(1-\alpha_{5}\right)\overline{L}\left(\left.c_{t,1}\right|z\right)&\text{if }c_{-t,\overline{K}}<c_{t,1},\end{cases}$
+for some $\alpha_{3},\alpha_{4},\alpha_{5}\in\left[0,1\right]$. Accordingly, we can set
+$F_{t}^{*}\left(\left.c_{-t,\overline{K}}\right|w\right):=\begin{cases}\alpha_{3}F_{t}^{*}\left(\left.c_{t,K}\right|w\right)+\left(1-\alpha_{3}\right)\cdot 1,&\text{if }c_{-t,\overline{K}}>c_{t,K},\\ \alpha_{4}F_{t}^{*}\left(\left.c_{t,j-1}\right|z\right)+\left(1-\alpha_{3}\right)F_{t}^{*}\left(\left.c_{t,j}\right|z\right),&\text{if }c_{-t,\overline{K}}\in\left(c_{t,j-1},c_{t,j}\right),\\ \alpha_{5}\cdot 0+\left(1-\alpha_{5}\right)F_{t}^{*}\left(\left.c_{t,1}\right|z\right).&\text{if }c_{-t,\overline{K}}<c_{t,1}.\end{cases}$ (43)
+(ii) If $\overline{K}>1$, we now move to the second largest element $c_{-t,\overline{K}-1}$. Then $\overline{L}\left(\left.c_{-t,\overline{K}-1}\right|z\right)$ can be expressed as follows:
+$\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right):=\begin{cases}\alpha_{6}\overline{L}\left(\left.c_{t,K}\right|z\right)+\left(1-\alpha_{6}\right)\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right),&\text{if }c_{t,K}<c_{-t,\overline{K}-1}<c_{-t,\overline{K}}\\ \alpha_{7}\overline{L}\left(\left.c_{t,j_{1}-1}\right|z\right)+\left(1-\alpha_{7}\right)\overline{L}\left(\left.c_{-t,\overline{K}}\right|z\right),&\text{if}c_{t,j_{1}-1}<c_{-t,\overline{K}-1}<c_{-t,\overline{K}}<c_{t,j_{1}}\\ \alpha_{8}\overline{L}\left(\left.c_{t,j_{2}-1}\right|z\right)+\left(1-\alpha_{8}\right)\overline{L}\left(\left.c_{t,j_{2}}\right|z\right),&\text{if }c_{t,j_{2}-1}<c_{-t,\overline{K}-1}<c_{t,j_{2}}<c_{-t,\overline{K}}\\ \alpha_{9}\cdot 0+\left(1-\alpha_{9}\right)\overline{L}\left(\left.c_{t,1}\right|z\right),&\text{if }c_{-t,\overline{K}-1}<c_{-t,\overline{K}}<c_{t,1},\end{cases}$
+Accordingly, we can set
+$F_{t}^{*}\left(\left.c_{-t,\overline{K}-1}\right|w\right):=\begin{cases}\alpha_{6}F_{t}^{*}\left(\left.c_{t,K}\right|w\right)+\left(1-\alpha_{6}\right)F_{t}^{*}\left(\left.c_{-t,\overline{K}}\right|w\right),&\text{if }c_{t,K}<c_{-t,\overline{K}-1}<c_{-t,\overline{K}},\\ \alpha_{7}F_{t}^{*}\left(\left.c_{t,j_{1}-1}\right|w\right)+\left(1-\alpha_{7}\right)F_{t}^{*}\left(\left.c_{-t,\overline{K}}\right|w\right),&\text{if }c_{t,j_{1}-1}<c_{-t,\overline{K}-1}<c_{-t,\overline{K}}<c_{t,j_{1}},\\ \alpha_{8}F_{t}^{*}\left(\left.c_{t,j_{2}-1}\right|w\right)+\left(1-\alpha_{8}\right)F_{t}^{*}\left(\left.c_{t,j_{2}}\right|w\right),&\text{if }c_{t,j_{2}-1}<c_{-t,\overline{K}-1}<c_{t,j_{2}}<c_{-t,\overline{K}},\\ \alpha_{9}\cdot 0+\left(1-\alpha_{9}\right)F_{t}^{*}\left(\left.c_{t,1}\right|w\right),&\text{if }c_{-t,\overline{K}-1}<c_{-t,\overline{K}}<c_{t,1}.\end{cases}$
+(iii) Iteratively, $F_{t}^{*}\left(\left.c\right|w\right)$ can be constructed in the same way at all $c\in{\cal C}_{-t}$.
+By construction, partial stationarity (31) and the monotonicity of $F_{t}^{*}\left(\left.\cdot\right|w\right)$ are both satisfied on ${\cal C}={\cal C}_{t}\cup{\cal C}_{-t}$.
+Step 4:
+Finally, we construct $F_{t}^{*}\left(\left.c\right|w\right)$ for $\𝒞c\in\mathcal{R}\backslash{\cal C}$.
+We set $F_{t}^{*}\left(\left.c\right|w\right)=0$ for $c<c_{t,1}$ and $F_{t}^{*}\left(\left.c\right|w\right)=1$ for $c\geq\overline{c}=\max{\cal C}+\delta$.
+For $c\in\left[c_{t,1},\overline{c}\right]$, there must exists some $\tilde{c}\in{\cal C}$ s.t.
+$c>\tilde{c}\in{\cal C}$ and $\overline{L}\left(\left.c\right|z\right)=\overline{L}\left(\left.\tilde{c}\right|z\right)$, and we then set
+$F_{t}^{*}\left(\left.c\right|w\right):=F^{*}\left(\left.\tilde{c}\right|w\right).$
+This guarantees (31) at any $\𝒞c\in\mathcal{R}\backslash{\cal C}$.
+This completes the construction $F_{t}^{*}\left(\left.c\right|w\right)$ for all $c\in\mathcal{R}$ at each $t=1,...,T$. Together, we have ensured that:
+(a) $F_{t}^{*}\left(\left.\cdot\right|w\right)$ is a proper conditional CDF;
+(b) partial stationarity holds since (31) is satisfied for all $c\in\mathcal{R}$;
+(c) period-$t$ marginal CCPs are matched since (30) holds for all $c\in{\cal C}_{t}$ (in Step 2).
+Observe also that each $F_{t}^{*}\left(\left.\cdot\right|w\right)$ defines a discrete distribution with finite support points. ∎
+Lemma 2 (From Per-Period to All-Period Construction).
+[…]
+$F_{t}^{*}\left(\left.\cdot\right|w\right)$
+as constructed in Lemma 1, such that (28) holds.
+Proof.
+Recall from Lemma 1 that each $F_{t}^{*}\left(\left.\cdot\right|w\right)$ defines a discrete distribution with finite support points.
+Let $\overline{{\cal C}}$ denote the union of support points of $F_{t}^{*}\left(\left.\cdot\right|w\right)$ across all $t=1,...,T$, and let $f_{t}^{*}\left(\left.\cdot\right|w\right)$ denote the corresponding probability mass function for $F_{t}^{*}\left(\left.\cdot\right|w\right)$.
+Then, by definition,
+$F_{t}^{*}\left(\left.c\right|w\right)=\sum_{\tilde{c}\in\overline{{\cal C}}:\tilde{c}\leq c}f_{t}^{*}\left(\left.\tilde{c}\right|w\right),\quad\forall c.$
+We now show how to construct a joint pmf $f^{*}\left(\left.\cdot\right|w\right)$ whose period-$t$ marginals are given by $f_{t}^{*}\left(\left.\cdot\right|w\right)$.
+For each $t$, define
+$c_{t}^{*}:=\max\left\{c\in\overline{{\cal C}}:\ F_{t}^{*}\left(\left.c\right|w\right)=F_{t}^{*}\left(\left.w_{t}^{{}^{\prime}}\theta\right|w\right)\right\},$ (44)
+which exists and is unique by the construction in Lemma 1.
+For each ${\bf c}\equiv\left(c_{1},...,c_{T}\right)\in\overline{{\cal C}}^{T}$, write
+$y_{t}\left(c_{t}\right)$ $:=\mathbf{\mathbbm{1}}\left\{c_{t}\leq c_{t}^{*}\right\},$
+$y\left({\bf c}\right)$ $:=\left(y_{1}\left(c_{1}\right),...,y_{T}\left(c_{T}\right)\right)^{{}^{\prime}}.$
+and define
+$f^{*}\left(\left.{\bf c}\right|w\right):=p\left(\left.y\left({\bf c}\right)\right|w\right)\prod_{t=1}^{T}\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}},$ (45)
+under the convention $10^{0}=$ 1 .
+We show that $f^{*}\left(\left.\cdot\right|w\right)$ is a probability mass function that characterizes a well-defined joint distribution of $\left(v_{i1}^{*},...,v_{iT}^{*}\right)$ and satisfies the requirements in Lemma 2.
+Step 1:
+First, note that the right-hand (45) only involves known (observed or constructed) quantities. In particular:
+  • $p\left(\left.y\right|w\right):=\mathbb{P}\left(\left.Y_{it}=y_{t}\forall t=1,...,T\right|W_{i}=w\right)$ is the (observed) joint CCP of observing a particular path of outcomes $y$ across all periods, given $W_{i}=w$.
+  • $f_{t}^{*}\left(\left.c\right|w\right)$ is the period-$t$ marginal pmf corresponding to $F_{t}^{*}\left(\left.c\right|w\right)$ defined in Lemma 1.
+  • $f_{t}\left(w\right)=\mathbb{P}\left(\left.Y_{it}=1\right|W_{i}=w\right)$ is the observed period-$t$ marginal CCP, with
+    $p_{t}\left(w\right)=F_{t}\left(\left.c_{t}^{*}\right|w\right)=\sum_{\tilde{c}\in\overline{{\cal C}}:\tilde{c}\leq c_{t}^{*}}f_{t}^{*}\left(\left.\tilde{c}\right|w\right).$ (46)
+     Step 2:
+We show that the period-$t$ marginal pmf implied by $f^{*}\left(\left.\cdot\right|w\right)$ coincides with $f_{t}^{*}\left(\left.\cdot\right|w\right)$.
+To see this, observe that, for any $t$ and $y_{t}\in\left\{0,1\right\}$, we have
+$\sum_{c_{t}\in\overline{{\cal C}}:y_{t}\left(c_{t}\right)=y_{t}}\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}}$
+$=$ $y_{t}\sum_{c_{t}\leq c_{t}^{*}}\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)}+\left(1-y_{t}\right)\sum_{c_{t}>c_{t}^{*}}\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{1-p_{t}\left(w\right)}$
+[…]
+$=$ $y_{t}\cdot 1+\left(1-y_{t}\right)\cdot 1$
+$=$ $1,$ (47)
+Hence, for any $c_{t}\in\overline{{\cal C}},$ the period-$t$ marginal implied by $f^{*}\left(\left.\cdot\right|w\right)$ is
+$\sum_{c_{-t}\in\overline{{\cal C}}^{T-1}}f^{*}\left(\left.c_{t},c_{-t}\right|w\right)$
+$=$ $\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}}\sum_{c_{-t}}p\left(\left.y\left(c_{t},c_{-t}\right)\right|w\right)\prod_{s\neq t}\frac{f_{s}^{*}\left(\left.c_{s}\right|w\right)}{p_{s}\left(w\right)^{y_{s}\left(c_{s}\right)}\left(1-p_{s}\left(w\right)\right)^{1-y_{s}\left(c_{s}\right)}}$
+$=$ $\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}}\sum_{y_{-t}}p\left(\left.y_{t}\left(c_{t}\right),y_{-t}\right|w\right)\sum_{c_{-t}:y_{-t}\left(c_{-t}\right)=y_{-t}}\prod_{s\neq t}\frac{f_{s}^{*}\left(\left.c_{s}\right|w\right)}{p_{s}\left(w\right)^{y_{s}\left(c_{s}\right)}\left(1-p_{s}\left(w\right)\right)^{1-y_{s}\left(c_{s}\right)}}$
+$=$ $\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}}\sum_{y_{-t}}p\left(\left.y_{t}\left(c_{t}\right),y_{-t}\right|w\right)\prod_{s\neq t}\sum_{c_{s}:y_{s}\left(c_{s}\right)=y_{s}}\frac{f_{s}^{*}\left(\left.c_{s}\right|w\right)}{p_{s}\left(w\right)^{y_{s}\left(c_{s}\right)}\left(1-p_{s}\left(w\right)\right)^{1-y_{s}\left(c_{s}\right)}}$
+[…]
+$=$ $\frac{f_{t}^{*}\left(\left.c_{t}\right|w\right)}{p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}}p_{t}\left(w\right)^{y_{t}\left(c_{t}\right)}\left(1-p_{t}\left(w\right)\right)^{1-y_{t}\left(c_{t}\right)}$
+$=$ $f_{t}^{*}\left(\left.c_{t}\right|w\right).$
+Step 3:
+We show that $f^{*}\left(\left.\cdot\right|w\right)$ is a valid joint pmf.
+Clearly, $f^{*}\left(\left.{\bf c}\right|w\right)\geq 0$, since all quantities on the right-hand side of (45) are nonnegative.
+In addition, since the period-$t$ marginal of $f^{*}\left(\left.\cdot\right|w\right)$ coincides with $f_{t}^{*}\left(\left.\cdot\right|w\right)$ as established in (2), we must have
+$\sum_{{\bf c}}f^{*}\left(\left.{\bf c}\right|w\right)=\sum_{c_{t}}f_{t}^{*}\left(\left.c_{t}\right|w\right)=1.$
+Hence, $f^{*}\left(\left.c\right|w\right)$ is a valid pmf and thus characterizes a well-defined joint distribution of $\left(v_{i1}^{*},...,v_{iT}^{*}\right)$.
+Step 4:
+Lastly, we show that (28) holds under $f^{*}\left(\left.\cdot\right|w\right)$. For any $y\in\left\{0,1\right\}^{T},$
+[…]
+A.3 Reconciliation with Khan, Ponomareva, and Tamer (2023)
+We show that under Assumption 1 and $X_{it}=Y_{i,t-1}$, our identifying condition (10) implies the following result in Khan, Ponomareva, and Tamer (2023):
+KPT(i): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\left|\gamma_{0}\right|>0.$
+KPT(ii): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{i,s}=0,Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\min\left\{0,\gamma_{0}\right\}>0.$
+KPT(iii): $\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{i,s}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\max\left\{0,\gamma_{0}\right\}>0.$
+KPT(iv): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\max\left\{0,\gamma_{0}\right\}>0.$
+KPT(v): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}>0.$
+KPT(vi): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}+\gamma_{0}>0.$
+KPT(vii): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\min\left\{0,\gamma_{0}\right\}>0.$
+KPT(viii): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0\text{$,$}Y_{i,s-1}=1\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}-\gamma_{0}>0.$
+KPT(ix): $\mathbb{P}\left(\left.Y_{it}=1,Y_{it-1}=0\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,Y_{i,s-1}=0\right|z\right)$Rightarrow $\left(z_{t}-z_{s}\right)^{{}^{\prime}}\beta_{0}>0.$
+Proof.
+With $X_{it}=Y_{i,t-1}$, our inequality restriction (15) can be equivalently rewritten as follows:
+$\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}+\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$
+[…]
+by enumerating the realization of $Y_{i,t-1}$.
+Note that the lower and upper expressions in the inequality (A.3) both have three possible (informative) outcomes depending on the value of $c$, leading to the 9 inequalities in KPT.
+We derive the first two inequalities KPT(i) and KPT(ii), and the rest of inequalities can be derived in the same way.
+KPT(i): consider the event that all indicators in condition (A.3) are equal to one, saying that
+$\max\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0},z_{t}^{{}^{\prime}}\beta_{0}\}\leq c\leq\min\{z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0},z_{s}^{{}^{\prime}}\beta_{0}\},$
+which is equivalent to
+$z_{t}^{{}^{\prime}}\beta_{0}+\max\{0,\gamma_{0}\}-(z_{s}^{{}^{\prime}}\beta_{0}+\min\{0,\gamma_{0}\})=(z_{t}-z_{s})^{{}^{\prime}}\beta_{0}+|\gamma_{0}|\leq 0.$
+Then, when $(z_{t}-z_{s})^{{}^{\prime}}\beta_{0}+|\gamma_{0}|\leq 0$, condition (A.3) becomes
+$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)$ $=\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)+\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)$
+$\leq\ 1-\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=1\right|z\right)-\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=0\right|z\right)$
+$=1-\mathbb{P}(Y_{is}=0\mid z)=\mathbb{P}\left(\left.Y_{is}=1\right|z\right).$
+By contraposition, it implies the same restriction in KPT(i):
+$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>\mathbb{P}\left(\left.Y_{is}=1\right|z\right)\Longrightarrow(z_{t}-z_{s})^{{}^{\prime}}\beta_{0}+|\gamma_{0}|>0.$
+KPT(ii): we first relax condition (A.3) by dropping the last term in the upper expression $\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{s}^{{}^{\prime}}\beta_{0}\geq c\right\}$ and have the following relaxed inequality:
+$\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0}\leq c\right\}+\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)\mathbf{\mathbbm{1}}\left\{z_{t}^{{}^{\prime}}\beta_{0}\leq c\right\}$
+[…]
+Now, consider the event that the indicators in the above restriction are all equal to one, which implies that
+$\max\{z_{t}^{{}^{\prime}}\beta_{0}+\gamma_{0},z_{t}^{{}^{\prime}}\beta_{0}\}\leq c\leq z_{s}^{{}^{\prime}}\beta_{0}+\gamma_{0},$
+and it is equivalent to the following condition:
+$(z_{t}-z_{s})^{{}^{\prime}}\beta_{0}+\max\{0,\gamma_{0}\}-\gamma_{0}=(z_{t}-z_{s})^{{}^{\prime}}\beta_{0}-\min\{0,\gamma_{0}\}\leq 0.$
+Given the above event, condition (A.3) becomes
+$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)$ $=\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=1\right|z\right)+\mathbb{P}\left(\left.Y_{it}=1,\ Y_{i,t-1}=0\right|z\right)$
+$\leq\ 1-\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=1\right|z\right).$
+Similarly, we can derive the same restriction in KPT(ii) by contraposition:
+$\mathbb{P}\left(\left.Y_{it}=1\right|z\right)>1-\mathbb{P}\left(\left.Y_{is}=0,\ Y_{i,s-1}=1\right|z\right)\Longrightarrow(z_{t}-z_{s})^{{}^{\prime}}\beta_{0}-\min\{0,\gamma_{0}\}>0.$
+A.4 Proof of Propositions 3 and 4
 Proof.
@@
 When $X_{it}$ is discrete and there are two periods $T=2$, the identifying condition (10) is given as
-$1-\Pr(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{12})\geq\Pr(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12}),$
-for $a\in\{z_{1}^{\prime}\beta_{0}+a_{1}^{\prime}\gamma_{0},...,z_{1}^{\prime}\beta_{0}+a_{K}^{\prime}\gamma_{0},z_{2}^{\prime}\beta_{0}+a_{1}^{\prime}\gamma_{0},...,z_{2}^{\prime}\beta_{0}+a_{K}^{\prime}\gamma_{0}\}$, and another identifying condition switches the order of period 1 and 2.
-Let $a=z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$, [^11] The value of $a=z_{2}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$ leads to the same identifying condition., then the above upper bound can be further bounded as
-$1-\Pr(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}\mid z_{12})\leq 1-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12}).$
-When $z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0$ which implies $z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}\geq z_{2}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$, then the lower bound can be bounded below as
-$\Pr(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{\mid}z_{12})\leq\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12}).$
+$1-\mathbb{P}(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq c\mid z)\geq\mathbb{P}(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq c\mid z),$
+for $c\in\{z_{t}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0},\ t=1,2,k=1,...,K\}$, and another identifying condition switches the order of period 1 and 2.
+Let $c=z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}$, [^11] The value of $c=z_{2}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}$ leads to the same identifying condition., then the above upper bound can be further bounded as
+$1-\mathbb{P}(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}\mid z)\leq 1-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{k}\mid z).$
+When $z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0$ which implies $z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}\geq z_{2}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}$, then the lower bound can be bounded below as
+$\mathbb{P}(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}\mid z)\leq\mathbb{P}(Y_{i2}=1,X_{i2}=x_{k}\mid z).$
 Combining the above results leads to
-If $\text{If }z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12})\geq\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12}).$
+If $\text{If }z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{k}\mid z)\geq\mathbb{P}(Y_{i2}=1,X_{i2}=x_{k}\mid z).$
 The contraposition of the above inequality yields
-$01-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}>$ 0 .
+$01-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{k}\mid z)<\mathbb{P}(Y_{i2}=1,X_{i2}=x_{k}\mid z)\Longrightarrow\Delta z^{\prime}\beta_{0}>$ 0 .
 Switching the order of the time period leads to another identifying restriction as follows:
-$01-\Pr(Y_{i1}=1,X_{i1}=a_{k}\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=a_{k}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
+$01-\mathbb{P}(Y_{i1}=1,X_{i1}=x_{k}\mid z)<\mathbb{P}(Y_{i2}=0,X_{i2}=x_{k}\mid z)\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
 Therefore, when $\Delta z\in\Delta\mathcal{Z}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified.
@@
 For the identification of $\gamma_{0}$, under the similar analysis for $\beta_{0}$, we have
-$(z_{1},z_{2})\in\mathcal{Z}^{j}_{3}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0},$
-$(z_{1},z_{2})\in\mathcal{Z}^{j}_{4}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}.$
+$(z_{1},z_{2})\in\mathcal{Z}_{3}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}<\Delta z^{\prime}\beta_{0},$
+$(z_{1},z_{2})\in\mathcal{Z}_{4}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}>\Delta z^{\prime}\beta_{0}.$
 As previously shown, when $(z_{1},z_{2})\in\mathcal{Z}_{2}$, it implies that $\Delta z^{\prime}\beta_{0}<0$.
-Therefore, when $(z_{1},z_{2})\in\mathcal{Z}_{2}\cap\mathcal{Z}^{j}_{3}$, we have $(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0}<0$ and the sign of $\gamma^{j}_{0}$ is identified given $x_{1}^{j}\neq x_{2}^{j}$.
-Similarly, when $(z_{1},z_{2})\in\mathcal{Z}_{1}\cap\mathcal{Z}^{j}_{4}$, the sign of $\gamma^{j}_{0}$ is also identified given $(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}>0$.
-Proposition 3 requires that for any $j\leq d_{x}$, either $\mathcal{Z}_{2}\cap\mathcal{Z}^{j}_{3}\neq\emptyset$ or $\mathcal{Z}_{1}\cap\mathcal{Z}^{j}_{4}\neq\emptyset$ so that the sign of $\gamma_{0}^{j}$ is identified for any $j$.
-A.4 Relationship to Khan, Ponomareva, and Tamer (2023)
-Lemma 3.
-Under Assumption 1 and $X_{it}=Y_{i,t-1}$, condition (7) implies the identifying restriction Khan, Ponomareva, and Tamer (2023), presented as follows:
-$(i)$ $P\left(y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+|\gamma_{0}|>0,$
-$(ii)$ $P\left(y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0,$
-$(iii)$ $P\left(y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}>0,$
-$(iv)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}>0,$
-$(v)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}>0,$
-$(vi)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\gamma_{0}>0,$
-$(vii)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0,$
-$(viii)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\gamma_{0}>0,$
-$(ix)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}>0.$
+Therefore, when $(z_{1},z_{2})\in\mathcal{Z}_{2}\cap\mathcal{Z}_{3}^{j}$, we have $(x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}<\Delta z^{\prime}\beta_{0}<0$ and the sign of $\gamma_{0}^{j}$ is identified given $x_{1}^{j}\neq x_{2}^{j}$.
+Similarly, when $(z_{1},z_{2})\in\mathcal{Z}_{1}\cap\mathcal{Z}_{4}^{j}$, the sign of $\gamma_{0}^{j}$ is also identified given $(x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}>\Delta z^{\prime}\beta_{0}>0$.
+Proposition 4 requires that for any $j\leq d_{x}$, either $\mathcal{Z}_{2}\cap\mathcal{Z}_{3}^{j}\neq\emptyset$ or $\mathcal{Z}_{1}\cap\mathcal{Z}_{4}^{j}\neq\emptyset$ so that the sign of $\gamma_{0}^{j}$ is identified for any $j$.
+∎
+A.5 Identification of Counterfactual Parameters
+In the main text, we have focused on the (partial) identification of the index parameters $\theta_{0}$.
+Here we show how our identification results can also be leveraged to (partially) identify counterfactual parameters.
+Write $W_{i}:=\left(Z_{i},X_{i}\right)$ in short, and correspondingly $w:=\left(z,x\right)$, and $w_{t}^{{}^{\prime}}\theta=z_{t}^{{}^{\prime}}\beta+x_{t}^{{}^{\prime}}\gamma$.
+Consider a general counterfactual change in the observable covariates $W_{i}$ from $w$ to $\tilde{w}$, and the consequent counterfactual period-$t$ conditional choice probability of the form
+$\tilde{p}_{t}\left(\tilde{w}\right):=\mathbb{P}\left(\left.v_{it}\leq\tilde{w}^{{}^{\prime}}\theta_{0}\right|W_{i}=w\right).$ (50)
+Importantly, in the definition above, the utility index is changed from $w^{{}^{\prime}}_{t}\theta_{0}$ to the counterfactual $\tilde{w}^{{}^{\prime}}_{t}\theta_{0}$, while the conditional distribution of the latent $v_{it}$ is held unchanged at $\left.v_{it}\right|W_{i}=w$.
+Hence, $\tilde{p}_{t}\left(\tilde{w}\right)$ can be interpreted as a counterfactual CCP induced by an exogenous policy intervention that only changes the characteristics from $w$ to $\tilde{w}$, but leaves all other unobserved individual heterogeneity reflected in the distribution of $v_{it}$ unchanged.
+In particular, note that the (partial) derivative of $\tilde{p}_{t}\left(w\right)$ can be interpreted as average marginal effects.
+Proposition 10 (Bounds on Counterfactual CCP).
+Under model 5 and Assumption 1,
+$\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta\leq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right)\leq\tilde{p}_{t}\left(\tilde{w}\right)\leq 1-\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta\geq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right).$ (51)
 Proof.
-When the endogenous regressor is the lagged dependent variable $X_{it}=Y_{it-1}$, the identifying condition in Proposition 1 is given as
-$U_{s}(a):=1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta_{0}+Y_{i,s-1}\gamma_{0}\geq a\mid z_{st})\\ \geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq a\mid z_{st}):=L_{t}(a).$
-Since $Y_{i,t-1}$ only takes two values $Y_{i,t-1}\in\{0,1\}$, we can express the probability in the above bounds as a mixture of $Y_{i,t-1}=1$ and $Y_{i,t-1}=0$:
-$U_{s}(a)=1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\mathbf{\mathbbm{1}}\{z_{s}^{\prime}\beta_{0}$ $+\gamma_{0}\geq a\}$
-$-\Pr(Y_{is}=0,Y_{i,s-1}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{s}^{\prime}\beta_{0}\geq a\}.$
-Similarly, the lower bound is given as
-$L_{t}(a)=\Pr(Y_{it}=1,Y_{i,t-1}=1\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}$ $+\gamma_{0}\leq a\}$
-$+\Pr(Y_{it}=1,Y_{i,t-1}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}\leq a\}.$
-Now we discuss different scenarios for the value of $(z_{t}^{\prime}\beta_{0},z_{s}^{\prime}\beta_{0},\gamma_{0})$, leading to various identifying restrictions in Khan, Ponomareva, and Tamer (2023) described in Lemma 3.
-Case 1: consider that $z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}\leq z_{s}^{\prime}\beta_{0}+\min\{0,\gamma_{0}\}$.
-Let $a=z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}$, the bounds can be simplified as
-$U_{s}(a)$ $=1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})-\Pr(Y_{is}=0,Y_{i,s-1}=0\mid z_{st})=\Pr(Y_{is}=1\mid z_{st}),$
-$L_{t}(a)$ $=\Pr(Y_{it}=1,Y_{i,t-1}=1\mid z_{st})+\Pr(Y_{it}=1,Y_{i,t-1}=0\mid z_{st})=\Pr(Y_{it}=1\mid z_{st}).$
-By exploiting the condition $\max\{0,\gamma_{0}\}-\min\{0,\gamma_{0}\}=|\gamma_{0}|$, the identifying condition is given as
-$(z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0})+|\gamma_{0}|\leq 0\Longrightarrow U_{s}(a)=\Pr(Y_{is}=1\mid z_{st})\geq L_{t}(a)=\Pr(Y_{it}=1\mid z_{st}).$
-The contraposition of the above condition yields condition (i) in Lemma 3:
-$\Pr(Y_{it}=1\mid z_{st})>\Pr(Y_{is}=1\mid z_{st})\Longrightarrow(z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0})+|\gamma_{0}|>0.$
-Case 2: consider that $z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}\leq z_{s}^{\prime}\beta_{0}+\gamma_{0}$. Let $a=z_{s}^{\prime}\beta_{0}+\gamma_{0}$, the bounds are simplified as
-$U_{s}(a)\leq 1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st}),\quad L_{t}(a)=\Pr(Y_{it}=1\mid z_{st}).$
-Given that $\max\{0,\gamma_{0}\}-\gamma_{0}=-\min\{0,\gamma_{0}\}$, we have the following identifying condition:
-$z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}\leq 0\Longrightarrow 1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\geq\Pr(Y_{it}=1\mid z_{st}).$
-The contraposition of this condition leads to condition (ii) in Lemma 3:
-$\Pr(Y_{it}=1\mid z_{st})>1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\Longrightarrow z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0.$
-The rest of conditions in Lemma 3 can be derived in a similar way so we omit the analysis here.
-A.5 Proof of Proposition 2
+By (52), we have
+$\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta_{0}\leq c\right|W_{i}=w\right)\leq F_{t}\left(\left.c\right|w\right)\leq 1-\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta_{0}\geq c\right|W_{i}=w\right).$
+Since $\tilde{p}_{t}\left(\tilde{w}\right)=F_{t}\left(\left.\tilde{w}_{t}^{{}^{\prime}}\theta_{0}\right|w\right)$, we have
+$\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta_{0}\leq\tilde{w}_{t}^{{}^{\prime}}\theta_{0}\right|W_{i}=w\right)\leq\tilde{p}_{t}\left(\tilde{w}\right)\leq 1-\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta_{0}\geq\tilde{w}_{t}^{{}^{\prime}}\theta_{0}\right|W_{i}=w\right),$
+and hence
+$\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta\leq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right)\leq\tilde{p}_{t}\left(\tilde{w}\right)\leq 1-\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta\geq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right).$
+The lower and upper bounds in Proposition 10 above are identified since the involved conditional probabilities are all about observed data $\left(Y_{i},W_{i}\right)$ for each $\theta\in\Theta_{I}$, while the set $\Theta_{I}$ is identified by Proposition 1.
+Hence, Proposition 10 establishes the partial identification of the counterfactual CCP $\tilde{p}_{t}\left(\tilde{w}\right)$.
+Note that the (partial) identification of counterfactual CCP $\tilde{p}_{t}\left(\tilde{w}\right)$ relies on the identification of the index parameter $\theta_{0}$ as well as the identification of the latent conditional distribution $\left.v_{it}\right|W_{i}=w$, which also involves the endogenous covariates $X_{i}$.
+It turns out that, our key identification strategy in Section 2.2 also provides a straightforward way to derive bounds on $F_{t}\left(\left.c\right|w\right)$, the CDF of $\left.v_{it}\right|W_{i}=w$ at any point $c$, by taking conditional expectations of (7) and (8) given $W_{i}=w$ (instead of $Z_{i}=z$ as in Section 2.2):
+$\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta_{0}\leq c\right|W_{i}=w\right)\leq F_{t}\left(\left.c\right|w\right)\leq 1-\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta_{0}\geq c\right|W_{i}=w\right),$ (52)
+which can then be combined with Proposition 1 to derive the bounds in Proposition 10.
+A.6 Proof of Proposition 8
 Proof.
-The identifying restriction (9) in Proposition 4 is given as follows:
-$U_{t,order}(a\mid z_{st})$ $:=1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})$ (18)
-$\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st}):=L_{t,order}(a\mid z_{st}).$
-The proof idea of Corollary 2 is the same as Appendix 1 for Corollary 1.
-Let $q_{k}\in\mathcal{Q}_{2}$ denote a point in $\mathcal{Q}_{2}$, and we can rank the $2K(J-1)$ points as $q_{1}\leq q_{2}...\leq q_{2K(J-1)}$.
-When $a>q_{2K(J-1)}$, then $\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})=0$ for any $j$ such that the upper bound is one $U_{t,order}(a\mid z_{st})=1$.
-Therefore, condition (18) holds for any $\theta$ and it is not informative for $\theta$.
-When $a<q_{1}$, then $\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st})=0$ for any $2\leq j\leq K$ so that the lower bound is zero $L_{t,order}(a\mid z_{st})=0$ and it is not informative for $\theta$.
-When $a\in(q_{k},q_{k+1})$, we show that the identifying condition (18) at $q_{k+1}$ contains more information than the one at $a$.
-The lower bound $L_{t,order}(a\mid z_{st})$ is increasing in $a$, therefore the lower bound evaluated at $q_{k+1}$ is larger than the one evaluated at $a$.
-The upper bound is the same when evaluated at $q_{k+1}$ and at $a$, as there does not exist any point $b_{j}-z_{s}^{\prime}\beta-a_{k}^{\prime}\gamma$ that lies in $[a,q_{k+1})$ from the definition of $q_{k}$.
-Therefore, if the parameter $\theta$ satisfies condition (18) evaluated at $q_{k+1}$, it must satisfy condition (18) evaluated at $a$.
-The identifying condition at $a$ does not provide additional information for $\theta$.
-A.6 Proof of Lemma 1
+Let $v_{ijt}:=\alpha_{ij}+\epsilon_{ijt}$, for any set $K\subset\mathcal{J}$, the probability of selecting a choice $j\in K$ conditional on $W_{i}=w$ is given as:
+$\mathbb{P}(Y_{it}^{K}\mid w)=\mathbb{P}(Y_{it}\in K\mid w)=\mathbb{P}\left(\exists j\in K\text{ s.t.
+}w_{ijt}^{\prime}\theta_{0}+v_{ijt}\geq w_{ikt}^{\prime}\theta_{0}+v_{ikt}\ \forall k\in K^{c}\mid w\right).$
+The above observed probability restricts the conditional distribution of $v_{ikt}-v_{ijt}\mid w$ and can be exploited to bound this distribution.
+We define $Q_{t}(c_{jk}\mid w)$ as follows: for $c_{jk}\in\mathcal{R}$,
+$Q_{t}(c_{jk}\mid w):=\mathbb{P}\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq c_{jk}\ \forall k\in\mathcal{J}\setminus K\mid w\right).$
+Then, we can derive lower and upper bounds for the above probability using variations in observed choice probabilities.
+When $c_{jk}$ satisfies $c_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ for any $j\in K$ and $k\in\mathcal{J}\setminus K$, then $Q_{t}(c_{jk}\mid w)$ can be bounded below as
+$Q_{t}(c_{jk}\mid w)$ $\geq\mathbb{P}\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall k\in\mathcal{J}\setminus K\mid w\right)$
+$=\mathbb{P}(Y_{it}\in K\mid w).$
+Therefore, the lower bound for $Q_{t}(c_{jk}\mid w)$ is established as
+$Q_{t}(c_{jk}\mid w)\geq\mathbb{P}(Y_{it}\in K,\ c_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in k\in\mathcal{J}\setminus K\mid w).$
+The above inequality holds since either $c_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the lower bound is zero.
+By taking expectation of $X_{i}$ given $z$, we can bound the conditional distribution $Q_{t}(c_{jk}\mid z)$ as
+$Q_{t}(c_{jk}\mid z)$ $\geq\mathbb{P}\left(Y_{it}\in K,c_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right)$
+$=\mathbb{P}\left(Y_{it}^{K}=1,c_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right).$
+Similarly, the conditional probability $Q_{t}(c_{jk}\mid w)$ can be bounded above as
+$Q_{t}(c_{jk}\mid w)\leq\mathbb{P}(Y_{it}^{K}=1\mid w)\mathbf{\mathbbm{1}}\{c_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,\mathcal{J}\setminus K\}+\\ 1-\mathbf{\mathbbm{1}}\{c_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\}.$
+The above inequality holds since either $c_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the upper bound is one with $c_{jk}>(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$.
+After taking expectation of $X_{i}$ given $z$, the upper bound for $Q_{t}(c_{jk}\mid z)$ is obtained as
+$Q_{t}(c_{jk}\mid z)\leq\mathbb{P}\left(Y_{it}^{K}=1,c_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z\right)\\ +1-\mathbb{P}\left(c_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right).$
+Rearranging the above formula yields
+$Q_{t}(c_{jk}\mid z)\leq 1-\mathbb{P}\left(Y_{it}^{K}=0,c_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right).$
+Under Assumption 1, the conditional probability $Q_{t}(c_{jk}\mid z)$ is the same for any $t$.
+Therefore, the smallest upper bound of $Q_{t}(c_{jk}\mid z)$ should be larger than the largest lower bound over all periods, yielding the identifying condition (23) as follows:
+$1-\max_{s=1,...,T}\mathbb{P}(Y_{is}^{K}=0,(z_{js}-z_{ks})^{\prime}\beta_{0}+(X_{ijs}-X_{iks})^{\prime}\gamma_{0}\geq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z)\\ \geq\max_{t=1,...,T}\mathbb{P}(Y_{it}^{K}=1,(z_{jt}-z_{kt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\leq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z).$
+A.7 Proof of Proposition 9
 Proof.
-When all regressors are exogenous in the sense $\epsilon_{is}\mid(W_{ist},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{ist},\alpha_{i})$, the identifying restriction in Proposition 4 is given as
-$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{s}^{\prime}\theta_{0}\geq a\}\\ \geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$ (19)
-For any $a\in\mathcal{R}$, the above identifying restriction is only informative when there exists $2\leq j_{1}\leq J$ and $11\leq j_{2}\leq J-$ 1 such that
-$b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq a,\quad b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\leq a.$ (20)
-It is because if the above condition does not hold, then either the upper bound in condition (19) becomes one or the lower bound becomes zero.
-Then condition (19) holds for any $\theta$ and there is no identifying power.
-Condition (20) can be satisfied for some $a$ if and only if there exists $2\leq j_{1}\leq J$ and $11\leq j_{2}\leq J-$ 1 such that $b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}$.
-In this case, the identifying condition (19) becomes
-if $\text{if }b_{j_{1}}-w_{s}^{\prime}\theta_{0}\geq b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\Longrightarrow\sum_{j=1}^{j_{1}-1}\Pr(Y_{is}=y_{j}\mid w_{st})=1-\sum_{j=j_{1}}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\\ \geq\sum_{j=1}^{j_{2}}\Pr(Y_{it}=y_{j}\mid w_{st}).$
-Replacing $j_{1}$ with $\tilde{j}_{1}=j_{1}-1$ yielding the results in Proposition 4.
-A.7 Proof of Proposition 5 and Proposition 6
-Proof.
-We first prove that $\beta_{0}$ is point identified up to scale by showing that for any $\Delta z\in\Delta\mathcal{Z}_{order}$, the sign of $\Delta z^{\prime}\beta_{0}$ is identified.
-The identifying condition for $\theta_{0}$ in Proposition 4 is given as
-$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{12})\geq\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12}),$
-and the same condition also holds when changing the order of period 1 and 2.
-Let $a=b_{k}-z_{1}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}$ for $2\leq k\leq J$, then we know that $b_{j}-z_{1}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}\geq a$ if $j\geq k$.
-Therefore, the upper bound becomes
-$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{st})\geq 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x).$
-Moreover, if $-z_{2}^{\prime}\beta_{0}\leq-z_{1}^{\prime}\beta_{0}$, then $b_{j+1}-z_{2}^{\prime}\beta_{0}-x^{\prime}\gamma_{0}\leq a$ for $j+1\leq k$. The lower bound can be bounded below by
-$\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
-Combining the results yields the following conditions:
-$\Delta z^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x)\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
-The contraposition of the above condition generates
-$01-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x)\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
-Therefore, when $z\in\mathcal{Z}_{1,order}$, the sign of $\Delta z^{\prime}\beta_{0}<0$ is identified.
-Similarly, the sign of $\Delta z^{\prime}\beta_{0}>0$ is also identified when $z\in\mathcal{Z}_{2,order}$.
-Given that the sign of $\Delta z^{\prime}\beta_{0}$ is identified, the parameter $\beta_{0}$ is point identified up to scale under the large support condition $\Delta z$.
-The analysis is the same as proof A.3 for Proposition 2, so it is omitted here.
-Now we look at the result for $\gamma_{0}$ in Proposition 6. Let $a=b_{k}-z_{1}^{\prime}\beta_{0}-x_{1}^{\prime}\gamma_{0}$, then the upper bound becomes
-$1-\sum_{j=1}^{J}\Pr(Y_{i1}=y_{j},b_{j}-z_{1}^{\prime}\beta_{0}-X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{st})\geq 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}).$
-If $b_{k}-z_{2}^{\prime}\beta_{0}-x_{2}^{\prime}\gamma_{0}\leq a$, the lower bound becomes
-$\sum_{j=1}^{J}\Pr(Y_{i2}=y_{j},b_{j+1}-z_{2}^{\prime}\beta_{0}-X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12}).$
-Combining the results leads to
-if $\text{if }-z_{2}^{\prime}\beta_{0}-x_{2}^{\prime}\gamma_{0}\leq-z_{1}^{\prime}\beta_{0}-x_{1}^{\prime}\gamma_{0}\Longrightarrow\\ 1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1})\geq\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12}).$
-Similarly, the contraposition of the above condition gives
-$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<(x_{1}-x_{2})^{\prime}\gamma_{0}.$
-Therefore, when $\Delta z\in\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}$, it implies that
-$0<\Delta z^{\prime}\beta_{0}<(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}.$
-Given that $x_{1}^{j}-x_{2}^{j}\neq 0$, it implies that the sign of $\gamma^{j}_{0}$ is identified.
-Similarly when $\Delta z\in\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}$, we have $0<\Delta z^{\prime}\beta_{0}<(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}$.
-The sign of $\gamma^{j}_{0}$ is also identified.
-For any $j\leq d_{x}$, Proposition 6 requires that either $\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}\neq\emptyset$ or $\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}\neq\emptyset$, so the sign of $\gamma^{j}_{0}$ is identified for any $j$.
-A.8 Proof of Proposition 7
-Proof.
-For any set $K\subset\mathcal{J}$, the conditional probability of selecting a choice $j\in K$ given $W_{ist}=w_{st}$ is:
-$\Pr(Y_{it}\in K\mid w_{st})=\Pr\left(\exists j\in K\text{ s.t. }w_{ijt}^{\prime}\theta_{0}+v_{ijt}\geq w_{ikt}^{\prime}\theta_{0}+v_{ikt}\ \forall k\in K^{c}\mid w_{st}\right).$
-The above observed probability provides information regarding the distribution of $v_{ikt}-v_{ijt}$.
-Therefore, we can use observed data to bound the following distribution: for $a_{jk}\in\mathcal{R}$,
-$Q_{t}(a_{jk}\mid w_{st}):=\Pr\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq a_{jk}\ \forall k\in K^{c}\mid w_{st}\right).$
-When $a_{jk}$ satisfies $a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ for any $(j,k)$, then the above probability can be bounded below as
-$Q_{t}(a_{jk}\mid w_{st})$ $\geq\Pr\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall k\in K^{c}\mid w_{st}\right)$
-$=\Pr(Y_{it}\in K\mid w_{st}).$
-Therefore, the lower bound for $Q_{t}(a_{jk}\mid w_{st})$ is established as
-$Q_{t}(a_{jk}\mid w_{st})\geq\Pr(Y_{it}\in K\mid w_{st})\mathbf{\mathbbm{1}}\{a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}.$
-The above inequality holds since either $a_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the lower bound is zero.
-Since Assumption 1 is only conditioning on the exogenous covariate $z_{st}$, we can bound the conditional distribution $Q_{t}(a_{jk}\mid z_{st})$ by taking expectation of $X_{it}$ given $z_{st}$:
-$Q_{t}(a_{jk}\mid z_{st})\geq\Pr\left(Y_{it}\in K,a_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
-Similarly, an upper bound for the conditional probability $Q_{t}(a_{jk}\mid w_{st})$ is derived as follows:
-$Q_{t}(a_{jk}\mid w_{st})\leq\Pr(Y_{it}\in K\mid w_{st})\mathbf{\mathbbm{1}}\{a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}+\\ 1-\mathbf{\mathbbm{1}}\{a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in K^{c}\}.$
-The above inequality holds since either $a_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ and the probability $Q_{t}(a_{jk}\mid w_{st})$ is weakly increasing or the upper bound is one.
-After taking expectation of $X_{ist}$ given $z_{st}$, the upper bound for $Q_{t}(a_{jk}\mid z_{st})$ is obtained as
-$Q_{t}(a_{jk}\mid z_{st})\leq\Pr\left(Y_{it}\in K,a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right)\\ +1-\Pr\left(a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
-Rearranging the above formula yields
-$Q_{t}(a_{jk}\mid z_{st})\leq 1-\Pr\left(Y_{it}\in K^{c},a_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z_{st}\right).$
-Under Assumption 1, the conditional probability $Q_{t}(a_{jk}\mid z_{st})$ is the same over the two periods $(s,t)$.
-Therefore, the upper bound at period $s$ of $Q_{t}(a_{jk}\mid z_{st})$ should be larger than the lower bound at period $t$, yielding the identifying condition (10) as follows:
-$1-\Pr(Y_{is}\in K^{c},(z_{js}-z_{ks})^{\prime}\beta_{0}+(X_{ijs}-X_{iks})^{\prime}\gamma_{0}\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st})\\ \geq\Pr(Y_{it}\in K,(z_{jt}-z_{kt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st}).$
-A.9 Proof of Proposition 8
-Proof.
-Since the observed outcome $Y_{it}$ is censored, we either observe $Y_{it}=y>0$ or $Y_{it}=0$. Let $v_{it}=-(\alpha_{i}+\epsilon_{it})$, the conditional probability of $Y_{it}=0$ is given as,
-$\Pr(Y_{it}=0\mid w_{st})=\Pr(Y^{*}_{it}\leq 0\mid w_{st})=\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w_{st}).$
+Since the observed outcome $Y_{it}$ is censored at 0, we either observe $Y_{it}=y>0$ or $Y_{it}=0$. Let $v_{it}:=-(\alpha_{i}+\epsilon_{it})$, the conditional probability of $Y_{it}=0$ is given as,
+$\mathbb{P}(Y_{it}=0\mid w)=\mathbb{P}(Y_{it}^{*}\leq 0\mid w)=\mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w).$
 When $y>0$, the conditional distribution is given as
-$\Pr(Y_{it}\leq y\mid w_{st})$ $=\Pr(Y^{*}_{it}\leq 0,Y_{it}\leq y\mid w_{st})+\Pr(0<Y^{*}_{it},Y_{it}\leq y\mid w_{st})$
-$=\Pr(Y^{*}_{it}\leq 0\mid w_{st})+\Pr(0<Y^{*}_{it}\leq y\mid w_{st})$
-$=\Pr(Y^{*}_{it}\leq y\mid w_{st})$
-$=\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-y\mid w_{st}).$
-Combining the two scenarios, the full conditional distributional of $Y_{it}\mid W_{ist}$ is characterized as follows:
-$\Pr(Y_{it}\leq y\mid w_{st})=\left\{\begin{aligned} \Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+&x_{t}^{\prime}\gamma_{0}-y\mid w_{st})\quad&\text{ if }y\geq 0,\\ &0\quad&\text{ if }y<0.\end{aligned}\right.$
-Given observed distribution of $Y_{it}\mid W_{ist}$, the identification approach is to bound the distribution $\Pr(v_{it}\leq a\mid z_{st})$.
-We first look at the upper bound for the conditional distribution $v_{it}\mid w_{st}$ given all covariates $W_{ist}=w_{st}$.
-$\Pr(v_{it}\geq a\mid w_{st})\leq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\geq a\}+\\ \Pr(Y_{it}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a\},$
-where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ such that $\Pr(Y_{it}\leq y\mid w_{st})=\Pr(v_{it}\geq a\mid w_{st})$, or $\Pr(v_{it}\geq a\mid w_{st})\leq\Pr(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w_{st})=\Pr(Y_{it}=0\mid w_{st})$ when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a$.
-Taking expectation over the endogenous covariate $X_{ist}$ yields the upper bound for the distribution $v_{it}\mid z_{st}$:
-$\Pr(v_{it}\geq a\mid z_{st})\leq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\\ \Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st}).$
-The above upper bound can be also also expressed as
-$\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq a\mid z_{st})+\Pr(Y_{it}=0,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<a\mid z_{st})$
-$=$ $\Pr(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a,z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\geq a\mid z_{st})+\Pr(Y_{it}=0\mid z_{st})$
-$=$ $\Pr(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{it}=0\mid z_{st}).$
-Similarly, the conditional distribution $v_{it}\mid w_{st}$ can be bounded below
-$\Pr(v_{it}\geq a\mid w_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a),$
-where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-a\geq 0$ such that $\Pr(Y_{it}\leq y\mid w_{st})=\Pr(v_{it}\geq a\mid w_{st})$, or the lower bound is zero when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<a$.
-Taking expectation over $X_{ist}$ leads to the following lower bound:
-$\Pr(v_{it}\geq a\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a\mid z_{st}).$
-The conditional stationarity condition requires $v_{is}\mid Z_{ist}\stackrel{{\scriptstyle d}}{{\sim}}v_{it}\mid Z_{ist}$, which implies that the bounds for $v_{it}\mid Z_{ist}$ must have intersections over any pair of periods $(s,t)$.
-This restriction generates the following identifying condition for $\theta_{0}$:
-$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-a\mid z_{st})$
-for any $a\in\mathcal{R}$, any $s,t$, and any $z_{st}$.
-A.10 Dynamic Censored Models with Latent Dependent Variables
-Consider the following dynamic censored models with the latent lagged outcome $Y^{*}_{i,t-1}$:
-$Y^{*}_{it}$ $=Z_{it}^{\prime}\beta_{0}+Y^{*}_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
-$Y_{it}$ $=\max\{Y^{*}_{it},0\},$
-In this model, the endogenous variable $X_{it}$ is the lagged outcome: $X_{it}=Y^{*}_{i,t-1}$.
-However, the variable $Y^{*}_{i,t-1}$ is not observed in data, so the results in Proposition 8 cannot be directly applied here.
+$\mathbb{P}(Y_{it}\leq y\mid w)$ $=\mathbb{P}(Y_{it}^{*}\leq 0,\ Y_{it}\leq y\mid w)+\mathbb{P}(0<Y_{it}^{*},\ Y_{it}\leq y\mid w)$
+$=\mathbb{P}(Y_{it}^{*}\leq 0\mid w)+\mathbb{P}(0<Y_{it}^{*}\leq y\mid w)$
+$=\mathbb{P}(Y_{it}^{*}\leq y\mid w)$
+$=\mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-y\mid w).$
+Combining the two scenarios, the conditional distributional of $Y_{it}\mid W_{i}$ is characterized as follows:
+$\mathbb{P}(Y_{it}\leq y\mid w)=\left\{\begin{aligned} \mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+&x_{t}^{\prime}\gamma_{0}-y\mid w)\quad&\text{ if }y\geq 0,\\ &0\quad&\text{ if }y<0.\end{aligned}\right.$
+Given observed distribution of $Y_{it}\mid W_{i}$, we can bound the distribution $\mathbb{P}(v_{it}\leq c\mid w)$ above as
+$\mathbb{P}(v_{it}\geq c\mid w)\leq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\mid w)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\geq c\}+\\ \mathbb{P}(Y_{it}=0\mid w)\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<c\},$
+where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ such that $\mathbb{P}(Y_{it}\leq y\mid w)=\mathbb{P}(v_{it}\geq c\mid w)$, or $\mathbb{P}(v_{it}\geq c\mid w)\leq\mathbb{P}(v_{it}\geq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}\mid w)=\mathbb{P}(Y_{it}=0\mid w)$ when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<c$.
+Taking expectation over the endogenous covariate $X_{i}$ yields the upper bound for the distribution $v_{it}\mid Z_{i}=z$:
+$\mathbb{P}(v_{it}\geq c\mid z)\leq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq c\mid z)+\\ \mathbb{P}(Y_{it}=0,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<c\mid z).$
+Rearranging the formula, the above upper bound can be equivalently written as
+$\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\geq c\mid z)+\mathbb{P}(Y_{it}=0,\ z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}<c\mid z)$
+$=$ $\mathbb{P}(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-c,\ z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma\geq c\mid z)+\mathbb{P}(Y_{it}=0\mid z)$
+$=$ $\mathbb{P}(0<Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-c\mid z)+\mathbb{P}(Y_{it}=0\mid z).$
+Similarly, the conditional distribution $v_{it}\mid w$ can be bounded below
+$\mathbb{P}(v_{it}\geq c\mid w)\geq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\mid w),$
+where the above condition holds since either $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ so that there exists $y=z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}-c\geq 0$ such that $\mathbb{P}(Y_{it}\leq y\mid w)=\mathbb{P}(v_{it}\geq c\mid w)$, or the lower bound is zero when $z_{t}^{\prime}\beta_{0}+x_{t}^{\prime}\gamma_{0}<c$.
+Taking expectation over $X_{i}$ leads to the following lower bound:
+$\mathbb{P}(v_{it}\geq c\mid z)\geq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c\mid z).$
+Given the bounds on the distribution $\mathbb{P}(v_{it}\geq c\mid z)$, the partial stationarity assumption implies the following identifying restriction for $\theta_{0}$:
+$\max_{t}\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}-c\mid z)\leq\max_{s}\{\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-c\mid z)+\mathbb{P}(Y_{is}=0\mid z)\},$
+for any $c\in\mathcal{R}$ and any $z$. ∎
+A.8 Dynamic Censored Models with Latent Dependent Variables
+Consider the following dynamic censored models with the latent lagged outcome $Y_{i,t-1}^{*}$:
+$Y_{it}^{*}$ $=Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}^{*}\gamma_{0}+\alpha_{i}+\epsilon_{it},$
+$Y_{it}$ $=\max\{Y_{it}^{*},0\},$
+In this model, the endogenous variable $X_{it}$ is the lagged outcome: $X_{it}=Y_{i,t-1}^{*}$.
+However, the variable $Y_{i,t-1}^{*}$ is not observed in data, so the results in Proposition 9 cannot be directly applied here.
 Due to this feature in the dynamic model, we need to adjust the results in Proposition 9.
-Given that $Y^{*}_{i,t-1}=Y_{i,t-1}$ when $Y_{i,t-1}>0$, we can further relax the lower and upper bounds in (8) to identify $\theta_{0}$.
+Given that $Y_{i,t-1}^{*}=Y_{i,t-1}$ when $Y_{i,t-1}>0$, we can further relax the lower and upper bounds in (9) to identify $\theta_{0}$.
 The lower bound in condition (9) can be bounded below as follows:
-$\Pr(Y_{it}\leq z_{t}^{\prime}\beta+Y^{*}_{i,t-1}\gamma-a\mid z_{st})\\ \geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}\gamma-a,Y_{i,t-1}>0\mid z_{st}):=L_{t,cen}(a\mid z_{st};\theta).$
+$\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}^{*}\gamma-c\mid z)\\ \geq\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta+Y_{i,t-1}\gamma-c,Y_{i,t-1}>0\mid z):=L_{t,cen}(c\mid z;\theta).$
 Similarly, the upper bound in condition (9) can be further bounded above
-$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+Y^{*}_{i,s-1}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\leq U_{s,cen}(a\mid z_{st};\theta),$
-where $U_{s,cen}(a\mid z_{st};\theta)$ is defined as
-$U_{s,cen}(a\mid z_{st};\theta):=\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}\gamma-a,Y_{i,s-1}>0\mid z_{st})\\ +\Pr(Y_{is}>0,Y_{i,s-1}=0\mid z_{st})+\Pr(Y_{is}=0\mid z_{st}).$
+$\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}^{*}\gamma-c\mid z)+\mathbb{P}(Y_{is}=0\mid z)\leq U_{s,cen}(c\mid z;\theta),$
+where $U_{s,cen}(c\mid z;\theta)$ is defined as
+$U_{s,cen}(c\mid z;\theta):=\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+Y_{i,s-1}\gamma-c,\ Y_{i,s-1}>0\mid z)\\ +\mathbb{P}(Y_{is}>0,\ Y_{i,s-1}=0\mid z)+\mathbb{P}(Y_{is}=0\mid z).$
 For the dynamic model, an identified set for $\theta_{0}$ is characterized by the following lemma:
 Lemma 3.
-Under Assumption 1 and $X_{it}=Y_{i,t-1}^{*}$, the identified set for $\theta_{0}$ is characterized by the set of parameters $\theta$ that satisfy the following conditions:
-$U_{s,cen}(a\mid z_{st};\theta)\geq L_{t,cen}(a\mid z_{st};\theta).$
-for any $a\in\mathcal{R}$, $s,t\leq T$, and $z_{st}$.
+Under Assumption 1 and $X_{it}=Y_{i,t-1}^{*}$, $\theta_{0}\in\tilde{\Theta}_{I,cen}$, where the identified set $\tilde{\Theta}_{I,cen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
+$\max_{t=1,...,T}L_{t,cen}(c\mid z;\theta)\leq\min_{s=1,...,T}U_{s,cen}(c\mid z;\theta).$
+for any $c\in\mathcal{R}$ and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.